### v1

# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

Abstract
The convergence rate of a Markov chain to its stationary distribution is typically assessed using the concept of total variation mixing time. However, this worst-case measure often yields pessimistic estimates and is challenging to infer from observations. In this paper, we advocate for the use of the average-mixing time as a more optimistic and demonstrably easier-to-estimate alternative. We further illustrate its applicability across a range of settings, from two-point to countable spaces, and discuss some practical implications.
Keywords— Ergodic Markov chain; average mixing time; countable state space; $\beta$-mixing.
1   Introduction
The total variation distance variant of the mixing time $t_{\mathsf{mix}}$ of a Markov chain $X_{1},\dots,X_{n}$ on a discrete state space $\mathcal{X}$, governed by a transition operator $P$, quantifies the time to convergence to its stationary distribution $\pi$. Two observations on the chain separated in time by about $t_{\mathsf{mix}}$ will be approximately independent and sampled from $\pi$. Estimates on the mixing time yield exponential concentration inequalities (Paulin, 2015), and in turn, generalization bounds for stable machine learning algorithms trained with Markovian data (Mohri and Rostamizadeh, 2010; Agarwal and Duchi, 2012; Oliveira et al., 2022), regret bounds for reinforcement learning algorithms (Ortner, 2020), and convergence diagnostics for Markov Chain Monte Carlo methods, which are widely used in mathematical physics (Metropolis et al., 1953; Hastings, 1970), computer science (Sinclair, 2012) and statistics (Chib and Greenberg, 1995).
However, $t_{\mathsf{mix}}$ is an intrinsically pessimistic notion, defined in terms of the worst-case initial distribution. Furthermore, $t_{\mathsf{mix}}$ is generally unknown a priori, and theoretical upper bounds, when available, are typically conservative. Finally, estimation of $t_{\mathsf{mix}}$ from a single trajectory of observations of length $n$ is known to be a statistically hard problem which requires at least
$n\in\Omega\left(\frac{1}{\pi_{\star}}+t_{\mathsf{mix}}\left|\mathcal{X}\right|\right),$ (1)
for estimation down to constant multiplicative error (Hsu et al., 2019, Theorem 3.1, Theorem 3.2), where $\left|\mathcal{X}\right|$ is the size of the space and $\pi_{\star}$ is the minimum stationary probability (refer to Section 1.4). The sample complexity lower bound in (1) is prohibitively large for vast state spaces, and even vacuous in the infinite case.
In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in (Espuny Díaz et al., 2022; Münch and Salez, 2022). Specifically, we will confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space. Furthermore, we initiate the program of estimating $t_{\mathsf{mix}}^{\sharp}$ from a single trajectory and we demonstrate that it is statistically less costly than the problem pertaining to $t_{\mathsf{mix}}$. In fact, contrary to the worst-case problem, our analysis carries to a large subset of countable-state Markov chains. One should rightfully expect $t_{\mathsf{mix}}^{\sharp}$ to be less informative than $t_{\mathsf{mix}}$ since it quantifies a weaker notion of mixing. However, we confirm that knowledge of $t_{\mathsf{mix}}^{\sharp}$ remains of practical value in the sense where it can be used in known powerful decoupling techniques (Yu, 1994) and appear as a natural parameter controlling the large deviation rate for additive functionals. Along the way, we explore the connections between the average-mixing time and $\beta$-mixing, argued by Vidyasagar (2013); Mohri and Rostamizadeh (2010) to be “just the right” assumption for the analysis of machine learning algorithms with weakly-dependent data.
1.1   Related work
Average variants of the mixing time have recently made their apparition in the literature. For instance, Berestycki et al. (2018) analyzed random walks on the giant component of the Erdös–Rényi random graph $\mathcal{G}(n,p)$ with $p=\Theta(1/n)$, and showed that the mixing time when starting randomly from any vertex is the order $\mathcal{O}(\log n)$ while it was known to be of the order $\mathcal{O}(\log^{2}n)$ for the mixing time from a worst starting point. Following up, Espuny Díaz et al. (2022) extended the approach on a larger collection of graphs with small bottlenecks, and show applicability of their approach for the Newman–Watts small-world model and supercritically percolated expander graphs. In (Lovász and Kannan, 1999), the authors introduce the notion of a conductance averaged over subsets of states with various sizes, and demonstrate that it can lead to tighter convergence bounds for counting and volume computation problems. Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered in (Münch and Salez, 2022) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
In particular, we stress that the notion of average-mixing treated in this paper is different from the one discussed in Peres and Sousi (2015); Anderson et al. (2021), defined by taking an average over two successive steps on the chain.
1.2   Main contributions
We direct the reader to Section 1.4 for the notation.
1.2.1 Highlight the significance of the average-mixing time
Observing that the average-mixing time $t^{\sharp}_{\mathsf{mix}}$ recently analyzed by Münch and Salez (2022) corresponds to the stationary $\beta$-mixing time of the Markov chain,
$t^{\sharp}_{\mathsf{mix}}(\xi)=\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\beta(s)\leq\xi\right\},$
we briefly demonstrate how the average-mixing time (together with an a priori known order of convergence) can be used to quantitatively capture the convergence rate of a Markov chain in lieu of the worst-case mixing time. Specifically, we show (Lemma 2.3) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ naturally appears in the sample complexity for controlling the deviation probability from 00 of a centered function $f\colon\mathcal{X}\to[-1,1]$. Namely, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$, it is sufficient to have a trajectory of
$n\geq\frac{\left(e/b\right)^{1/b}}{b\varepsilon^{2}}t^{\sharp}_{\mathsf{mix}}\left(\delta/e\right)\operatorname{polylog}\left(\frac{1}{\delta},\frac{1}{\varepsilon}\right),$
where $\operatorname{polylog}$ is a poly-logarithmic function of its parameters and $e$ is the Euler number. Similarly, in the polynomial $\beta$-mixing setting of order $b$, we show that for some natural constant $C\in\mathbb{R}_{+}$,
$n\geq C\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(2t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\right)^{\frac{1}{b}},$
is sufficient.
1.2.2 Estimation of $\beta$-mixing coefficients
For a general time-discrete, time-homogeneous Markov chain over a countable state space, we establish a parametric rate of convergence for its individual $\beta$-coefficients from a single stationary trajectory of observations $X_{1},\dots,X_{n}$. Namely, for $s\in\mathbb{N}$, we construct an estimator $\widehat{\beta}(s)$ for $\beta(s)$, and we show (Theorem 3.1) that
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
with
$\mathcal{B}_{p}^{(s)}\leq\left\|\beta^{(s)}\right\|_{1/p}^{1/p},\qquad\mathcal{J}_{p}^{(s)}=\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
and where for $x,x^{\prime}\in\mathcal{X}$, $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$. Here, $\mathcal{B}_{p}^{(s)}$ is a mixing parameter roughly controlled from above by the $\ell_{1/p}$ quasi-norm of the sequence of $\beta$-mixing coefficients of the $s$-skipped Markov chain, and $\mathcal{J}_{p}^{(s)}$ is an entropic quantity pertaining to the $s$-skipped Markov chain. Crucially, the parameter $p\in[1,\infty)$ is free and will allow us to balance the two factors in the upper bound.
1.2.3 Estimation of the average-mixing time
In the uniformly ergodic setting, for $\xi\in(0,1)$ and $\delta\in(0,1)$, we show (Theorem 3.4) that a single path length of
$n\geq\frac{t_{\mathsf{mix}}}{\xi^{2}}\inf_{p\geq 1}\left\{p\mathcal{J}_{p,3\xi/4}\right\}\operatorname{polylog}\left(\frac{1}{\xi},\frac{1}{\delta},t_{\mathsf{mix}}\right),$
where $t_{\mathsf{mix}}$ is the mixing time of $P$ and
$\mathcal{J}_{p,\xi}\doteq\sup\left\{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$
is sufficient in order to obtain an estimate $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ of the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\xi)$ which satisfies
$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(3\xi/2),t_{\mathsf{mix}}^{\sharp}(\xi/2)\right]\end{split}$
with probability at least $1-\delta$. Furthermore, for finite spaces, setting $\xi=1/4$ for simplicity, the parameter $p$ can be chosen so as to obtain an upper bound (Corollary 4.2) of
$n\geq Ct_{\mathsf{mix}}\log\max\left\{\left|\mathcal{X}\right|,\frac{t_{\mathsf{mix}}}{\delta}\right\},$
where $C\in\mathbb{R}_{+}$ is a universal constant. One can thus demonstrably estimate with high probability the average-mixing time sub-linearly in the number of states, answering a question posed by Luca Zanetti and John Sylvester (Zanetti and John, 2023). This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024).
1.3   Outline
Section 2 introduces the average-mixing time, illustrates its significance and analyzes some of its properties. In Section 2.2, under geometric ergodicity and time-reversibility assumptions, we establish an upper bound of $t^{\sharp}_{\mathsf{mix}}$ in terms of the relaxation time of the chain, which is non-vacuous over countable state spaces and under suitable conditions on the stationary distribution. In Section 2.3 we recall some important connection between the average-mixing time and $\beta$-mixing properties of the chain.
Section 3 is dedicated to the problem of inferring the $\beta$-mixing coefficients and average-mixing time from a single trajectory of observations. In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain; first individually and then as a sequence, with respect to the uniform norm. In Section 3.3, we specialize our analysis to the uniformly ergodic setting, and in Section 3.4 we deduce an upper bound for the problem of estimating the average-mixing time from a single trajectory under uniform ergodicity.
In Section 4, we demonstrate that our contribution holds implications at various scales of state spaces and provide more explicit rates under additional structural assumptions on the underlying chain.
Technical proofs have been deferred to Section 5 for readability.
Finally, Section 6 introduces a class of Markov chain which illustrates some of our results.
1.4   Notation and setting
We let $\mathcal{X}$ be a countably infinite space of symbols equipped (for simplicity) with a total order. Real functions on $\mathcal{X}$ are written as infinite row vectors and for $x\in\mathcal{X}$, $e_{x}$ is the function defined by $x^{\prime}\in\mathcal{X}$, $e_{x}(x^{\prime})=\delta\left[x=x^{\prime}\right]$. For $p\in(0,\infty)$, and a function $v\in\mathbb{R}^{\mathcal{X}}$,
$\left\|v\right\|_{p}\doteq\left(\sum_{x\in\mathcal{X}}\left|v(x)\right|^{p}\right)^{1/p},$
for convenience, we define
$\left\|v\right\|_{0}\doteq\left|\left\{x\in\mathcal{X}\colon\left|v(x)\right|>0\right\}\right|,$
and for any $p\in[0,\infty)$,
$\ell_{p}\doteq\left\{s\in\mathbb{R}^{\mathcal{X}}\colon\left\|s\right\|_{p}<\infty\right\}.$
The set of all probability distributions over $\mathcal{X}$, written $\mathcal{P}(\mathcal{X})$, corresponds to non-negative functions summing to unity. A time-discrete, time-homogeneous Markov chain
$X_{1},X_{2},X_{3}\dots$
on $\mathcal{X}$ is defined by a transition operator represented by an infinite stochastic matrix $P$
$P(x,x^{\prime})=\mathbb{P}\left(X_{t+1}=x^{\prime}|X_{t}=x\right).$
We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$. For a finite trajectory of length $n\in\mathbb{N}$, started from an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
$\mathbb{P}_{\mu}\left(X_{1}=x_{1},\dots,X_{n}=x_{n}\right)=\mu(x_{1})\prod_{t=1}^{n-1}P(x_{t},x_{t+1}).$
By abuse of notation, matrix multiplication will retain its meaning in the countable setting. Namely, for $\mu\in\mathcal{P}(\mathcal{X})$ and $P\in\mathcal{W}(\mathcal{X})$, we will write,
$\mu P=\sum_{x,x^{\prime}\in\mathcal{X}}\mu(x)P(x,x^{\prime})e_{x^{\prime}}.$
Extensions of notation to $Pf^{\intercal}$ with $f\in\ell_{1}$, or for $P^{t}$ are similarly defined. We say that $P$ is ergodic when $P$ is irreducible, aperiodic and positive-recurrent (Levin et al., 2009). In that case, there exists a unique stationary distribution $\pi$ verifying $\pi P=\pi$, and we can define the distance to stationarity
$d(t)\doteq\sup_{x\in\mathcal{X}}\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}},$
where $\left\|\cdot\right\|_{\mathsf{TV}}$ is the total variation distance defined for $\mu,\nu\in\mathcal{P}(\mathcal{X})$ by $\left\|\mu-\nu\right\|_{\mathsf{TV}}\doteq\frac{1}{2}\left\|\mu-\nu\right\|_{1}$. We say that $P$ is uniformly ergodic whenever, there exist $M>0$ and $\rho\in(0,1)$ such that for any $t\in\mathbb{N}$,
$d(t)\leq M\rho^{t}.$
We write $\mathcal{W}^{\star}(\mathcal{X})$ the set of uniformly ergodic Markov chains. We recall that the (worst-case) mixing time is defined as
$\begin{split}t_{\mathsf{mix}}(\xi)\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d(t)\leq\xi\right\}.\end{split}$
It is furthermore customary to define
$\qquad t_{\mathsf{mix}}\doteq t_{\mathsf{mix}}(1/4),$ (2)
as a consequence of the following property (Levin et al., 2009, Lemma 4.11)
$t_{\mathsf{mix}}(\xi)\leq\left\lceil\log_{2}\frac{1}{\xi}\right\rceil t_{\mathsf{mix}}.$ (3)
Elements related directly to the notion of average-mixing time will be introduced in the next section for clarity of the exposition.
2   The average-mixing time
In this section, we first recall the definition of the average-mixing time, and subsequently explore some of its basic properties. For $P\in\mathcal{W}(\mathcal{X})$, the average distance to stationarity of the chain induced from $P$ at time $t\in\mathbb{N}$ is expressed as
$d^{\sharp}(t)\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}},$
where $\pi$ is the stationary distribution of $P$. The above should be understood as the total variance distance of the chain to stationarity, averaged with respect to $\pi$ over the starting state. The limit $\lim_{t\to\infty}d^{\sharp}(t)=0$ always exists [^1] This follows from a connection between average-mixing time and $\beta$-mixing discussed at Section 2.3, thus for $\xi\in(0,1)$, the average-mixing time of $P$, defined (Münch and Salez, 2022) as
$\begin{split}t^{\sharp}_{\mathsf{mix}}(\xi)&\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d^{\sharp}(t)\leq\xi\right\},\end{split}$
always verifies $t^{\sharp}_{\mathsf{mix}}(\xi)<\infty$.
Remark 2.1.
Similar to $t_{\mathsf{mix}}$ in (2), it is tempting to introduce $t_{\mathsf{mix}}^{\sharp}\doteq t_{\mathsf{mix}}^{\sharp}(1/4)$. However, since $d^{\sharp}$ does not enjoy the same sub-multiplicativity properties as $d$ (3) (refer to (Levin et al., 2009, Lemma 4.11)), it is generally not enough to consider a fixed value of $\xi\in(1/2)$, and in this manuscript we will keep $\xi$ to be a user-fixed value.
2.1   Average-mixing versus worst-case mixing
Following their respective definitions, it holds that $d^{\sharp}(t)\leq d(t)$. Thus, for any fixed $\xi\in(0,1)$, average-mixing is always faster than mixing. The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2022, Remark 1), which in particular include random walks on groups [^2] A chain is called transitive when for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$.. In general however, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as formalized below.
Proposition 2.1.
Let $M\in\mathbb{R}_{+}$ be arbitrarily large. There exists a transition operator $P$ such that for any $\xi\in(0,1)$, it holds that
$t_{\mathsf{mix}}(\xi)>Mt_{\mathsf{mix}}^{\sharp}(\xi).$
Proof.
Lemma 4.6 establishes this fact in the case where $\mathcal{X}=\left\{0,1\right\}$. ∎
2.2   Spectral methods under reversibility and geometric ergodicity
In this section, we endow $\mathbb{R}^{\mathcal{X}}$ with the inner product $\langle f,g\rangle_{\pi}\doteq\sum_{x\in\mathcal{X}}f(x)g(x)\pi(x)$, where $\pi$ is a positive distribution. We regard $P$ as a linear operator on the resulting Hilbert space $\ell_{2}(\pi)\doteq(\mathbb{R}^{\mathcal{X}},\langle\cdot,\cdot\rangle_{\pi})$ and further assume $P$ to be geometrically ergodic. Recall that $P\in\mathcal{W}^{\star}(\mathcal{X})$ is called $\pi$-reversible when the following detailed-balance equation holds,
$\pi(x)P(x,x^{\prime})=\pi(x^{\prime})P(x^{\prime},x),\qquad\forall x,x^{\prime}\in\mathcal{X},$
which is equivalent to stating that $P$ is self-adjoint in $\ell_{2}(\pi)$. Under reversibility, the spectrum of $P$, denoted $\operatorname{Spec}(P)$, is contained in the real line. The absolute spectral gap of $P$ is then defined as
$\gamma_{\star}\doteq 1-\sup\left\{\left|\lambda\right|\colon\lambda\in\operatorname{Spec}(P),\lambda\neq 1\right\}.$
Geometric ergodicity and existence of a positive spectral gap are known to be equivalent (Kontoyiannis and Meyn, 2012). We first establish that, at least under mild assumptions on $P$, the average-mixing time can be controlled by the relaxation time $t_{\mathsf{rel}}\doteq 1/\gamma_{\star}$, even over countably infinite spaces.
Lemma 2.1.
Let $P\in\mathcal{W}^{\star}(\mathcal{X})$ be $\pi$-stationary, geometrically ergodic, reversible and let $\xi\in(0,1)$. When $P$ regarded as the linear operator $P\colon\ell_{2}(\pi)\to\ell_{2}(\pi)$ is compact [^3] For instance, being trace-class is a sufficient condition for $P$ to be compact., it holds that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{2\xi},$
where $t_{\mathsf{rel}}$ is the relaxation time of $P$.
The half quasi-norm of the stationary distribution $\|\pi\|_{1/2}$, making its first appearance in our study, is a natural measure of complexity of the stationary distribution $\pi$, directly related to its Rényi entropy of order $1/2$. Note that $\|\pi\|_{1/2}$ could be infinite, in which case Lemma 2.1 becomes vacuous. However, Lemma 2.1 compares favorably with the bound of Münch and Salez (2022, Theorem 5, Corollary 2), which is inversely proportional to the minimum (non-zero) transition probability $P_{\star}\doteq\min_{x,x^{\prime}\in\mathcal{X}}P(x,x^{\prime})$. Additionally, Lemma 2.1 can be instructively compared with Levin et al. (2009, Theorem 12.4) for the mixing time $t_{\mathsf{mix}}(\xi)$,
$t_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{1}{\xi\pi_{\star}},$
where $\pi_{\star}\doteq\min_{x\in\mathcal{X}}\pi(x)$, which is only informative [^4] We note in passing that the bound at Levin et al. (2009, Theorem 12.4) could be improved as $t_{\mathsf{mix}}(\xi)\leq\frac{1}{2}t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{\xi\pi_{\star}}$. when $\left|\mathcal{X}\right|<\infty$.
2.3   Connection with $\beta$-mixing
Recall that for a process $\{X_{t}\}_{t\in\mathbb{Z}}$ on $\mathcal{X}$, for any $s\in\mathbb{Z}$, the $\beta$-mixing coefficient (Bradley, 2005; Doukhan, 2012) is defined by
$\beta(s)\doteq\sup_{r\in\mathbb{N}}\beta\left(\sigma\left(\left\{X_{t}\colon t\leq r\right\}\right),\sigma\left(\left\{X_{t}\colon t\geq r+s\right\}\right)\right),$
where $\sigma(\{X_{t}\}_{t})$ is the $\sigma$-field generated by the random variables $\{X_{t}\}_{t}$, and for two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$,
$\beta(\mathcal{A},\mathcal{B})\doteq\sup\frac{1}{2}\sum_{i=1}^{I}\sum_{j=1}^{J}\left|\mathbb{P}(A_{i}\cap B_{i})-\mathbb{P}(A_{i})\mathbb{P}(B_{j})\right|,$
where the supremum is taken over all pairs of finite partitions $\{A_{1},\dots,A_{I}\}$ and $\{B_{1},\dots,B_{J}\}$ of $\mathcal{X}$ such that for any $1\leq i\leq I$ and any $1\leq j\leq J$, $A_{i}\in\mathcal{A}$ and $B_{j}\in\mathcal{B}$ (Bradley, 2005, (7)). The random process $\{X_{t}\}_{t\in\mathbb{Z}}$ is called $\beta$-mixing, or absolutely regular, when
$\lim_{s\to\infty}\beta(s)=0.$
Recall that a stationary, countable-state Markov chain $X$ is $\beta$-mixing if and only if it is irreducible and aperiodic (Bradley, 2005, Theorem 3.2).
Lemma 2.2 (Davydov (1974, Proposition 1)).
We let $\{X_{t}\}_{t\in\mathbb{N}}$ be a countable-state Markov chain with transition operator $P$. Writing $\mu^{(0)}$ for the initial distribution, and for $t>1$, $\mu^{(t)}=\mu^{(0)}P^{t}$ for the marginal distribution after time step $t$, it holds that
$\beta(s)=\sup_{t\in\mathbb{N}}\sum_{x\in\mathcal{X}}\mu^{(t)}(x)\left\|e_{x}P^{s}-\mu^{(s+t)}\right\|_{\mathsf{TV}}.\$
It follows from the above lemma that when the chain is stationary, $\mu^{(0)}=\pi=\mu^{(t)}$ for any $t\in\mathbb{N}$, and
$\begin{split}\beta(s)=\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{s}-\pi\right\|_{\mathsf{TV}}=d^{\sharp}(s),\\ \end{split}$
thus the average-mixing time is essentially the “stationary $\beta$-mixing time” of the chain555 Similarly, the worst-case mixing time can be connected to the notion of $\phi$-mixing (Davydov, 1974). . Note that with the convention $P^{0}=I$, we obtain
$\beta(0)=1-\left\|\pi\right\|_{2}^{2},$
which depends only on the stationary properties of the chain. When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially. When there exists $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially fast when $b\neq 1$, and exponentially fast when $b=1$. In Section 6, we will introduce a class of Markov transition operators, whose flavor of mixing (algebraic, sub-exponential, exponential) can be fully controlled. Notably (refer to the proof of Theorem 3.2), when $P$ is uniformly ergodic with mixing time $t_{\mathsf{mix}}$, it holds that
$\beta(s)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$ (4)
hence the process is exponentially $\beta$-mixing with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. However, nothing a priori precludes the $\beta$-mixing rate from being superior to (4). The connection between the average-mixing time and $\beta$-mixing provides additional motivation for the problem of estimating the $\beta$-mixing coefficients, whose knowledge is known to lead to powerful decoupling techniques (Yu, 1994). The problem of inferring mixing coefficients from a single trajectory of observations has recently received attention from the community. In the most recent work (Khaleghi and Lugosi, 2021), consistent estimators for $\beta$- and $\alpha$-mixing coefficients were proposed. In Section 3, we will provide more explicit and finer finite-sample bounds for the estimation problem in the specialized Markovian setting. A regime of particular interest to us is when the chain $\beta$-mixes much faster than $\Theta(t_{\mathsf{mix}})$. Henceforth in this manuscript, when they coincide under stationarity, we may use the notations $\beta$ and $d^{\sharp}$ interchangeably.
2.4   Bounding the deviation of additive functionals evaluated on Markov chains
We now illustrate how the average-mixing time, through its connection to $\beta$-mixing naturally appears as a quantity of interest when bounding deviations of functions evaluated on a trajectory sampled from a Markov chain. More formally, let $X_{1},\dots,X_{n}$ be a Markov chain over a countable space $\mathcal{X}$, and whose dynamics are governed by $P\in\mathcal{W}(\mathcal{X})$ with stationary distribution $\pi$. For simplicity, we consider a real bounded function $f$ on $\mathcal{X}$, which is centered in the sense where its stationary expected value vanishes $\mathbb{E}_{\pi}f=0$. The ergodic theorem states that the empirical mean taken on an infinite trajectory will similarly vanish,
$\lim_{n\to\infty}\frac{1}{n}\sum_{t=1}^{n}f(X_{t})=0.$
In practice however, we are more likely to be interested in controlling the deviation of the sample mean from its expected value in a non-asymptotic setting.
Lemma 2.3 (Bounding deviation of additive functionals evaluated on Markov chains).
Let $f\colon\mathcal{X}\to[-1,1]$ with $\mathbb{E}_{\pi}f=0$, let $\varepsilon,\delta\in(0,1)$, and let $t^{\sharp}_{\mathsf{mix}}\colon(0,1)\to\mathbb{N}$. For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, and average-mixing time bounded from above by $t^{\sharp}_{\mathsf{mix}}$, the following statements hold.
  1. 1.
    (Sub-exponential) If there exist $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ with $\beta_{0}\geq 1$, $b\in(0,1]$, such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$, then there exists a universal constant $C\in\mathbb{R}_{+}$ such that for
    $n\geq\frac{C}{\varepsilon^{2}}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right)\log\left(\frac{1}{\delta}\right),$
    with
    $\xi(\varepsilon,\delta)\doteq\frac{\delta\varepsilon^{2}}{16\log\left(\frac{4}{\delta}\right)},$
    it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$. The above upper on the path length can be further streamlined as
    $n\geq\frac{C}{\varepsilon^{2}}\frac{1}{b}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\delta/e\right)\log\left(\frac{1}{\delta}\right)\log\left(\frac{\log(1/\delta)}{\varepsilon}\right),$
    for some universal constant $C\in\mathbb{R}_{+}$.
  2. 2.
    (Polynomial) If there exist $\beta_{1},b\in\mathbb{R}_{+}$ with $\beta_{1}\geq 1$, such that for any $s\in\mathbb{N}$, $\beta(s)\leq\frac{\beta_{1}}{s^{b}}$, then there exists a universal constant $C\in\mathbb{R}_{+}$ such that for
    $n\geq C\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(2t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\right)^{\frac{1}{b}},$
    it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$.
Proof sketch.
We adopt the traditional method as described in Yu (1994). This approach combines a blocking technique, credited to Bernstein (1927) in the context of $m$-dependent processes, together with an application of Hoeffding’s inequality to control the deviation probability of a sum of independent observations. We further optimize for the block size and reformulate the results to obtain an expression involving the average-mixing time under various rates of $\beta$-mixing. Refer to Section 5.1 for the proof. ∎
As a consequence of Lemma 2.3 , in both the sub-exponential and polynomial settings, provided an assumption on the convergence order $b$, estimates on the average-mixing time to proximity $\Theta(\delta)$ yield corresponding deviation bounds. In the uniformly ergodic setting, corresponding to the assumption of the exponential setting ($b=1$), we obtain a deviation inequality involving the average-mixing time in lieu of the mixing time (Chung et al., 2012, Theorem 3.1). The lemma generalizes to non-centered functions, lower deviations and arbitrarily bounded functions. It is also noteworthy that in the sub-exponential setting, the convergence order $b$ solely affects the constant in the sample complexity. Our formulation above serves as additional motivation for analyzing and inferring the average-mixing time.
3   Estimation from a single trajectory
In this section, we analyze the estimation of $\beta$-mixing coefficients and the average-mixing time from a single trajectory of observations.
3.1   Estimation of $\beta$-mixing coefficients under general ergodicity condition
In this section we analyse plug-in estimators for $\beta$-mixing coefficients. We recall that in this general setting, the mixing time can be infinite, and the $\beta$-mixing coefficients can decay at arbitrary slow rate. We will focus on exponential, sub-exponential and polynomial rates of convergence. Our analysis relies on stationary skipped trajectories, where for $s\in\mathbb{N},s<n$, we denote
$\begin{split}X^{(s)}\doteq X_{1},X_{1+s},X_{1+2s},\dots,X_{1+\lfloor(n-1)/s\rfloor s},\end{split}$
with transition operator $P^{s}$ and initial distribution $\pi$. It will be convenient to define the following counting random variables,
$\begin{split}N_{x}^{(s)}&\doteq\sum_{t=1}^{\lfloor(n-1)/s\rfloor}\delta\left[X_{1+s(t-1)}=x\right],\\ N_{xx^{\prime}}^{(s)}&\doteq\sum_{t=1}^{\lfloor(n-1)/s\rfloor}\delta\left[X_{1+s(t-1)}=x,X_{1+st}=x^{\prime}\right].\end{split}$
We use the shortcut notation $N_{x}=N_{x}^{(1)}$ and $N_{xx^{\prime}}=N_{xx^{\prime}}^{(1)}$ to respectively denote number of visits to state $x$ and transitions from $x$ to $x^{\prime}$ on the original chain. We define the following estimator $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ for $\beta(s)$,
$\widehat{\beta}(s)=\frac{1}{2\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|,$ (5)
and by convention, we set $\widehat{\beta}(s)=0$ for $s\geq n$.
Theorem 3.1.
Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, it holds that
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{\mathcal{B}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}}\left(\sqrt{\frac{3}{2}\mathcal{J}_{p}^{(s)}}+2\sqrt{\left\|\pi\right\|_{1-1/p}^{1-1/p}}\right)\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
where for $x,x^{\prime}\in\mathcal{X}$, we wrote $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$, and
$\mathcal{B}_{p}^{(s)}\doteq\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p},\qquad\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}.$
We observe that the two key quantities $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$ control the estimation rate. The former essentially captures $\beta$-mixing properties of the process, while the latter is tightly connected to entropic properties of the distribution of (skipped) stationary pairs. In the following Corollary 3.1, we readily provide more explicit upper bounds on $\mathcal{B}_{p}^{(s)}$ under additional assumptions on the mixing rate. The analysis of $\mathcal{J}_{p}^{(s)}$ is deferred to Section 4.
Corollary 3.1.
Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$ with $\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}<\infty$, the following statements hold.
  1. 1.
    (Exponential) If $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s)$, with $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$ and $t\in\mathbb{N}\cup\left\{0\right\}$,
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}\left(\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}\right)\mathcal{J}_{p}^{(s)}}.\\ \end{split}$
  2. 2.
    (Sub-exponential) If $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$, with $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$, $b\in(0,1)$, and $s\in\mathbb{N}\cup\left\{0\right\}$.
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{\beta_{0}^{1/p}}{\lfloor(n-1)/s\rfloor}\left(1+\exp(-\beta_{1}s^{b}/p)+\frac{\Gamma\left(\frac{1}{b},s^{b}\beta_{1}/p\right)}{bs(\beta_{1}/p)^{1/b}}\right)\mathcal{J}_{p}^{(s)}},\end{split}$
[…]
  3. 3.
    (Polynomial) If $\beta(s)\leq\beta_{1}/s^{b}$, with $\beta_{1}\in\mathbb{R}_{+}$, $b\in(1,\infty)$, $b/p\in(1,\infty)$, $s\in\mathbb{N}$ and $\beta(0)=\beta_{0}\in\mathbb{R}_{+}$,
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{\beta_{0}^{1/p}+\zeta(b/p)\frac{\beta_{1}^{1/p}}{s^{b/p}}}{\lfloor(n-1)/s\rfloor}\mathcal{J}_{p}^{(s)}},\end{split}$
    where $\zeta$ is the Riemann zeta function defined by $\zeta(r)=\sum_{t=1}^{\infty}t^{-r}$.
3.2   Comparison with prior results
In McDonald et al. (2015, Theorem 4), the authors establish the nearly parametric convergence rate for continuously valued Markov processes
$\mathbb{E}{\left|\widehat{\beta}(s)-\beta(s)\right|}\leq\mathcal{O}\left(\sqrt{\frac{W_{0}(n)}{n}}\right),$
where $W_{0}$ is the Lambert function (Corless et al., 1996). They also posit that their findings may be considered conservative when applied to discrete-valued processes. Theorem 3.1 validates their remark by recovering the superior parametric rate $\mathcal{O}(\sqrt{1/n})$ under convergence of $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$. Our method shares similarities with Khaleghi and Lugosi (2021), as both employ Rio’s covariance inequality Rio (1999, Corollary 1.1). However, compared to both Khaleghi and Lugosi (2021) and McDonald et al. (2015, Theorem 4), we distinguish ourselves by focusing on the discrete space setting and deriving explicit rates in terms of fine-grained properties of the transition operator.
3.3   Estimation of $\beta$-mixing coefficients under uniform ergodicity
Under uniform ergodicity, the mixing time is finite, and the $\beta$-mixing coefficients decays exponentially. In this setting, we show how to obtain a sample complexity upper bound for high-probability estimates on the invividual $\beta$-mixing coefficients.
Theorem 3.2.
Let $s,n\in\mathbb{N}$ with $n>2s+1$, let $p\in\mathbb{R}_{+},p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any uniformly ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$,
  1. $(i)$
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 8\sqrt{\frac{1}{n-1}\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}},\\ \end{split}$
    where $t_{\mathsf{mix}}$ is the mixing time of $P$.
  2. $(ii)$
    Let $\varepsilon,\delta\in(0,1)$. For
    $n\geq 1+\frac{2^{8}}{\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{2}{\delta}\right\},$
    with probability at least $1-\delta$, it holds that
    $\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon.$
Using a truncated estimator with a cut-off determined from a uniform bound on the complexity terms $\|Q^{(s)}\|_{1-1/p}^{1-1/p}$, we can control the sequence of $\beta$-mixing coefficients with respect to the uniform norm modulo an additional logarithmic factor in the sample complexity.
Theorem 3.3.
Let $\varepsilon\in(0,1/4],\delta\in(0,1),p\in\mathbb{R}_{+},p\geq 1$ and $\mathcal{J}_{p}\in\mathbb{R}_{+}$. Let $n,S\in\mathbb{N}$, and for $s\in\mathbb{N}$, let $\widehat{\beta}_{[S]}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined as
$\widehat{\beta}_{[S]}(s)=\begin{cases}\widehat{\beta}(s)&\text{ for }s\leq S\\ 0&\text{ otherwise,}\end{cases}$
where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$ such that,
$\sup_{s\in\mathbb{N}}\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\leq\mathcal{J}_{p},$
when
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\varepsilon}\right\rceil\max\left\{\mathcal{J}_{p},18\log\left(\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\varepsilon}\right\rceil\right)\right\},$
with probability at least $1-\delta$, it holds that,
$\left\|\widehat{\beta}_{[S(n,\varepsilon,\delta,p,\mathcal{J}_{p})]}-\beta\right\|_{\infty}\leq\varepsilon,$
where the cut-off is set as
$S(n,\varepsilon,\delta,p,\mathcal{J}_{p})\doteq\frac{(n-1)\varepsilon^{2}}{2^{9}\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}}.$
Remark 3.1.
We note but do not pursue the fact that a more delicate bound on the sample complexity in Theorem 3.3 could be obtained, replacing a logarithm with a Lambert $W$ function.
In Section 4, we show how to obtain bounds on the entropy term in several natural settings.
3.4   Estimation of the average-mixing time
We move on to the estimation of the average-mixing time. Since the average distance to stationarity $d^{\sharp}=\beta$ does not generally enjoy sub-multiplicativity (it does not hold that $t^{\sharp}_{\mathsf{mix}}(\xi)(1/4)$ can be bounded by $t^{\sharp}_{\mathsf{mix}}$ and a polynomial of $\log\frac{1}{\xi}$), we formalize desired guarantees somewhat differently from Hsu et al. (2019); Wolfer and Kontorovich (2024). Indeed, for the relaxation time $t_{\mathsf{rel}}$, we were seeking an estimator $\widehat{t_{\mathsf{rel}}}$ to multiplicative error $\varepsilon$. Namely, the estimators in Hsu et al. (2019); Wolfer and Kontorovich (2024) guarantee that with high probability,
$\left|\frac{\widehat{t_{\mathsf{rel}}}}{t_{\mathsf{rel}}}-1\right|\leq\varepsilon.$
Here, we keep the distance to stationarity as a free parameter, and construct an estimator for the average-mixing time for proximity parameter $\xi$ up to multiplicative accuracy $\varepsilon$. Define for $\xi\in(0,1)$,
$\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\doteq\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\widehat{\beta}(s)\leq\xi\right\},$ (6)
where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (5).
Theorem 3.4.
Let $\varepsilon\in(0,1/4),\delta,\xi\in(0,1),p\in\mathbb{R}_{+}$ and $\mathcal{J}_{p,\xi}\in\mathbb{R}_{+}$. Let $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (6). For any uniformly ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$ such that,
$\sup\left\{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\}\leq\mathcal{J}_{p,\xi},$
when
$n\geq 1+\frac{2^{8}t_{\mathsf{mix}}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+\frac{p}{\log 2}\right)\mathcal{J}_{p,\frac{3\xi}{4}},9\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+1\right)\log\left(\frac{4t_{\mathsf{mix}}}{\delta}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil\right)\right\},$
with probability at least $1-\delta$, it holds that
$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
3.5   Extension to non-stationary Markov chains
Although results in this section are stated assuming a stationary start, it is possible to generalize them to non-stationary chains as follows. Let $\mu\in\mathcal{P}(\mathcal{X})$ absolutely continuous with respect to $\pi$. Inspired by the proof of Paulin (2015, Proposition 3.10), an application of Hölder’s inequality with $q,r\geq 1$ with $11/q+1/r=$ 1 yields,
$\mathbb{E}_{\mu}f(X_{1},\dots,X_{n})\leq\left\|\mu(X)/\pi(X)\right\|_{\pi,q}\left\|f(X_{1},\dots,X_{n})\right\|_{\pi,r},$
and in particular, for any $\varepsilon>0$, we can bound the probability of an $\varepsilon$-deviation assuming a non-stationary start as
$\mathbb{P}_{\mu}\left(f(X_{1},\dots,X_{n})>\varepsilon\right)\leq\left\|\mu(X)/\pi(X)\right\|_{\pi,q}\mathbb{P}_{\pi}\left(f(X_{1},\dots,X_{n})>\varepsilon\right)^{1/r}.$
Setting $q=r=2$ recovers Paulin (2015, Proposition 3.10). The reader is invited to consult (Paulin, 2015, Section 3.3) for additional methods involving “burning” a prefix of the observed trajectory.
3.6   Comparison with an alternative method
We point out the fact that for the special case $p=2$, by using the techniques in Paulin (2015), one can obtain an upper bound of the order (ignoring logarithmic factors) of
$\frac{t_{\mathsf{mix}}}{\varepsilon^{2}\xi^{2}}\mathcal{J}_{2,\xi}$
for the problem of estimating the average-mixing time in uniformly ergodic Markov chains. Indeed, the mixing time of the $s$-skipped pairs of observations is upper bounded by $t_{\mathsf{mix}}^{(s)}+s\leq\lceil t_{\mathsf{mix}}/s\rceil+s$. From Paulin (2015, Theorem 3.2, Proposition  3.4), it follows that
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]&\leq 2\lfloor(n-1)/s\rfloor\lceil t_{\mathsf{mix}}/s\rceil\pi(x),\\ \operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]&\leq 2\lfloor(n-1)/s\rfloor(\lceil t_{\mathsf{mix}}/s\rceil+s)\pi(x)P^{s}(x,x^{\prime}),\\ \end{split}$
which yields
$\begin{split}\mathbb{E}\left|\widehat{\beta}(s)-\beta(s)\right|\leq&\sqrt{2\frac{\lceil t_{\mathsf{mix}}/s\rceil+s}{\lfloor(n-1)/s\rfloor}}\left(\left\|Q^{(s)}\right\|_{1/2}^{1/2}+2\left\|\pi\right\|_{1/2}^{1/2}\right).\\ \end{split}$
Our method is superior in that it introduces a trade-off parameter $p$ to obtain an upper bound of order
$\frac{t_{\mathsf{mix}}}{\varepsilon^{2}\xi^{2}}\inf_{p\geq 1}\left\{p\mathcal{J}_{p,\xi}\right\}.$
This enables the use of more versatile bounds on the entropic term, which we explore in Section 4.
4   Implications at various scales of state space
In infinitely denumerable state spaces, recall that for any $s\in\mathbb{N}$, the well-definition of the estimation rate for the $\beta(s)$-mixing coefficient (Theorem 3.1) hinges on the finiteness of the following quantity
$\mathcal{J}^{(s)}_{p}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
where $p\in[1,\infty)$ is a user-fixed parameter. It is instructive to observe that on one hand
$\lim_{p\to\infty}\mathcal{J}_{p}^{(s)}=1,$
while on the other hand, $\mathcal{B}_{p}^{(s)}$ increases with $p$,
$\lim_{p\to\infty}\mathcal{B}_{p}^{(s)}\asymp\frac{n}{s},$
leading to a trade-off in the choice of the parameter in Theorem 3.1. In particular, provided a control which holds uniformly over $s$,
$\mathcal{J}_{p,\xi}\doteq\sup\left\{\mathcal{J}^{(s)}_{p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$ (7)
the free parameter $p$ can be appropriately selected in Theorem 3.4, in order to optimize the upper bound on the sample complexity of estimating the average-mixing time. It is also noteworthy that since the $\ell_{1-1/p}$ quasi-norm and the $(1-1/p)$th order Rényi entropy $H_{1-1/p}$ of a distribution $\mu\in\mathcal{P}(\mathcal{X})$ are related by the expression
$p\log\left\|\mu\right\|_{1-1/p}^{1-1/p}=H_{1-1/p}(\mu),$
the quantity $\mathcal{J}_{p,\xi}$ pertains to a specific entropic property inherent in the Markov process. Unfortunately, this term is generally not easily amenable to analysis; hence, our focus lies in establishing bounds for its evaluation. Under ergodicity of $P$, the pair-wise limit
$\lim_{s\to\infty}Q^{(s)}(x,x^{\prime})=\pi(x)\pi(x^{\prime}),\qquad\forall x,x^{\prime}\in\mathcal{X},$
hints at a heuristic lower bound in $\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}$ on $\mathcal{J}_{p,\xi}$. Using the stationary regime as a baseline, it seems natural to examine chains where the entropic term does not undergo relative explosive growth in the transient regime. For instance, for some multiplicative factor $K\geq 1$, an upper bound
$\mathcal{J}_{p}^{(s)}\leq K\left\|\pi\right\|^{2(1-1/p)}_{1-1/p},$ (8)
with a controlled $K$ readily simplifies the sample complexities appearing in Section 3. In the present section, we proceed to discuss bounds on the entropic quantities $\mathcal{J}_{p}^{(s)}$ and $\mathcal{J}_{p,\xi}$ in terms of state-wise conditional entropy, structural properties of the connection graph of the Markov chain, or the alphabet size. We conclude with a discussion on the binary state space, where the dominant complexity stems from the mixing bottlenecks, revealing that our results hold implications at every scale.
4.1   Ergodic properties
In this subsection, we derive various controls of $\mathcal{J}_{p}^{(s)}$ based on ergodic properties of the chain. Specifically, we introduce three conditions. Condition 1 is the classical notion of $V$-geometric ergodicity, which is known to be weaker than uniform ergodicity. Condition 2 imposes a stronger criterion, under which an ergodic chain is necessarily uniformly ergodic. Both Conditions 1 and 2 allow to derive bounds on $\mathcal{J}_{p}^{(s)}$ that are uniform over $s$. However, these conditions are quite stringent. We introduce a natural extension of these two conditions, Condition 3, which also allows to control $\mathcal{J}_{p}^{(s)}$. However, under Condition 3, this bound might be growing (sub-linearly) with $s$. We summarize the relations between these conditions in Figure 1.
{tikzcd}
Figure 1: Relation between the various ergodic conditions in Subsection 4.1. In this diagram, $\Rightarrow$ is to be read as ”implies under the assumptions of the result referred to”.
Condition 1 ($V$-geometric ergodicity (Douc et al., 2018, Section 18)).
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. We say that $P$ is $V$-geometrically ergodic for some function $V\colon\mathcal{X}\to[1,\infty)$, when there exist constants $\theta\in(0,1)$ and $M\in\mathbb{R}_{+}$ such that for any $x\in\mathcal{X}$,
$\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s},$
where for any $v\in\mathbb{R}^{\mathcal{X}}$, we wrote
$\left\|\mu\right\|_{V}\doteq\sup_{x\in\mathcal{X}}V(x)v(x).$
Condition 2 (Existence of a dominating distribution).
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Suppose that there exists a universal constant $C\in\mathbb{R}_{+}$ and a probability measure $\nu\in\mathcal{P}(\mathcal{X})$, such that for all $x^{\prime}\in\mathcal{X}$,
$\sup_{x\in\mathcal{X}}P(x,x^{\prime})\leq C\nu(x^{\prime}).$
Condition 3 (Weak $V$-geometric ergodicity).
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. We say that $P$ is weakly $V$-geometrically ergodic when there exists functions
$V\colon\mathcal{X}\to[1,\infty),\qquad D\colon\mathcal{X}\to[1,\infty),$
and a constant $b\geq 0$ such that for any $s\in\mathbb{N}$,
$\sup_{x^{\prime}\in\mathcal{X}}V(x^{\prime})(P^{s}(x,x^{\prime})-\pi(x^{\prime}))\leq D(x)+b(s-1).$
Lemma 4.1.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary weakly $V$-geometrically ergodic transition operator, with $V,D,b$ defined as in Condition 3, and let $p\geq 1$ such that
$\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty.$
Then,
$\mathcal{J}_{p}^{(s)}\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right).$
Proof.
Directly from the assumptions,
$\mathcal{J}_{p}^{(s)}$ $\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left(\pi(x)\left[\pi(x^{\prime})+\frac{1}{V(x^{\prime})}\left(D(x)+b(s-1)\right)\right]\right)^{1-1/p}$
$\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\sum_{x,x^{\prime}\in\mathcal{X}}\left[\pi(x)\frac{1}{V(x^{\prime})}\left(D(x)+b(s-1)\right)\right]^{1-1/p}$
$\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right),$
where the two inequalities stem from the sub-additivity of the function $x\mapsto x^{1-1/p}$. ∎
Corollary 4.1 (Under $V$-geometric ergodicity).
Let $P$ a $\pi$-stationary $V$-geometrically ergodic transition operator, with $V,M,\theta$ defined as in Condition 1. Then Condition 3 is satisfied with
$D=M\theta V\text{ and }b=0.$
If the following summability conditions are met for some $p\geq 1$,
$\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\pi\circ V\right\|_{1-1/p}^{1-1/p}<\infty,$
we obtain a control of $\mathcal{J}_{p}^{(s)}$ uniformly over $s$:
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\\ \leq\|\pi\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left\|\pi\circ D\right\|^{1-1/p}_{1-1/p}.$
Proof.
We first check Condition 3. For any $x,x^{\prime}\in\mathcal{X}$,
$V(x)(P^{s}(x,x^{\prime})-\pi(x^{\prime}))$ $\leq\sup_{x^{\prime\prime}\in\mathcal{X}}V(x^{\prime\prime})|P^{s}(x,x^{\prime\prime})-\pi(x^{\prime\prime})|$
$=\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s}\leq MV(x)\theta,$
where the second inequality stems from the definition of $V$-geometric ergodicity. Thus, Condition 3 is satisfied with $D=M\theta V$ and $b=0$. An application of Lemma 4.1 gives the the second claim. ∎
Remark 4.1.
Under the assumption $D\propto V$, we can clarify the summability assumptions in Corollary 4.1.
  • Let us first assume that $\pi(x)\propto x^{-\alpha}$ for $\alpha\in(1,\infty)$. Then, for $V(x)\propto D(x)\propto x^{\beta}$, with $\beta\in\mathbb{R}_{+}$, $\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty$ holds only when $\beta>p/(p-1)$, and $\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty$ holds only when $\alpha-\beta>p/(p-1)$, thus we must have $p/(p-1)<\beta<\alpha-p/(p-1)$, which will only be possible for $\alpha>2p/(p-1)$.
  • Let us now assume now that $\pi(x)\propto a^{-x}$ for $a\in(0,1)$. Then, for $V(x)\propto D(x)\propto x^{\beta}$, $\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty$ always holds but we still require $\beta>p/(p-1)$ in order for $\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty$ to hold.
Lemma 4.2 (Under the existence of a dominating distribution).
Assume Condition 2 is satisfied. If moreover, for $p\geq 1$,
$\left\|\pi\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\nu\right\|_{1-1/p}^{1-1/p}<\infty,$
then,
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq C^{1-1/p}\left\|\nu\right\|_{1-1/p}^{1-1/p}\left\|\pi\right\|_{1-1/p}^{1-1/p}.$
Proof.
It directly follows from the assumptions that for any $x,x^{\prime}\in\mathcal{X}$, and for any $s\in\mathbb{N}$,
$\begin{split}Q^{(s)}(x,x^{\prime})&=\pi(x)\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s-1}(x,x^{\prime\prime})P(x^{\prime\prime},x^{\prime})\\ &\leq C\pi(x)\nu(x^{\prime})\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s-1}(x,x^{\prime\prime})=C\pi(x)\nu(x^{\prime}).\end{split}$
As a result,
$\sum_{x,x^{\prime}\in\mathcal{X}}Q^{(s)}(x,x^{\prime})^{1-1/p}\leq C^{1-1/p}\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}\nu(x^{\prime})^{1-1/p}.$
In practice, Condition 3 can be readily verified through the following lemma.
Lemma 4.3.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Assume there exists functions $V\colon\mathcal{X}\to[1,\infty)$, $D\colon\mathcal{X}\to[1,\infty)$ and a constant $b\geq 0$ such that
$\sup_{x^{\prime}\in\mathcal{X}}V(x^{\prime})(P(x,x^{\prime})-\pi(x^{\prime}))\leq D(x)$
and, for all $x\in\mathcal{X}$, $PD(x)\leq D(x)+b$. Then, Condition 3 holds.
In Section 6, we use Lemma 4.3 to check Condition 3 on a whole family of transition operators $(P_{T})_{T\in\mathcal{T}}$. Note that Lemma 4.3 also shows that Condition 2 immediately implies Condition 3 with $D=C$, $V=1$ and $b=0$, which completes Figure 1.
Proof.
By recursion, for any $t\in\mathbb{N}$,
$\forall x\in\mathcal{X},\,P^{t}D(x)\leq D(x)+bt.$ (9)
Indeed, (9) holds for $t=1$ by assumption. Assume it holds for $t-1$, then
$P^{t}D(x)$ $=P[P^{t-1}D](x)$
$=\sum_{x^{\prime}\in\mathcal{X}}P(x,x^{\prime})[P^{t-1}D](x^{\prime})$
$\leq\sum_{x^{\prime}\in\mathcal{X}}P(x,x^{\prime})[D(x^{\prime})+b(t-1)]$
$=PD(x)+b(t-1)$
$\leq D(x)+bt.$
Thus, for any $x^{\prime\prime}\in\mathcal{X}$,
$V(x^{\prime\prime})[P^{t}(x,x^{\prime\prime})-\pi(x^{\prime\prime})]$ $=\sum_{x^{\prime}\in\mathcal{X}}P^{t-1}(x,x^{\prime})[P(x^{\prime},x^{\prime\prime})-\pi(x^{\prime\prime})]V(x^{\prime\prime})$
$\leq\sum_{x^{\prime}\in\mathcal{X}}P^{t-1}(x,x^{\prime})D(x^{\prime})$
$=P^{t-1}D(x)$
$\leq D(x)+b(t-1),$
where the last inequality follows from (9). ∎
4.2   Uniformly bounded conditional entropy
When for each state $x\in\mathcal{X}$, the conditional distribution $e_{x}P$ has bounded entropy, it is possible to guarantee finiteness of $\mathcal{J}_{p,\xi}$.
Lemma 4.4.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Suppose that for $p\geq 1$,
$\mathcal{I}_{p}\doteq\sup_{x\in\mathcal{X}}\left\{\left\|e_{x}P\right\|^{1-1/p}_{1-1/p}\right\}<\infty.$
Then, for any $s\in\mathbb{N}$, it holds that
$\mathcal{J}_{p}^{(s)}\leq\mathcal{I}_{p}^{s}\left\|\pi\right\|_{1-1/p}^{1-1/p}.$
Proof.
Suppose that the claim holds for $s\geq 1$. Then,
$\begin{split}\mathcal{J}_{p}^{(s+1)}&=\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)^{1-1/p}\left(\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s}(x,x^{\prime\prime})P(x^{\prime\prime},x^{\prime})\right)^{1-1/p}\\ &\leq\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s}(x,x^{\prime\prime})^{1-1/p}P(x^{\prime\prime},x^{\prime})^{1-1/p}\\ &=\sum_{x,x^{\prime\prime}\in\mathcal{X}}\pi(x)^{1-1/p}P^{s}(x,x^{\prime\prime})^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}P(x^{\prime\prime},x^{\prime})^{1-1/p}\\ &\leq\sum_{x,x^{\prime\prime}\in\mathcal{X}}\pi(x)^{1-1/p}P^{s}(x,x^{\prime\prime})^{1-1/p}\sup_{x\in\mathcal{X}}\left\|e_{x}P\right\|_{1-1/p}^{1-1/p}\\ &\leq\mathcal{J}_{p}^{(s)}\mathcal{I}_{p}.\end{split}$
The case $s=1$ is easy to verify, thus the lemma holds by induction. ∎
While a direct consequence of Lemma 4.4 is that
$\mathcal{J}_{p,\xi}\leq\mathcal{I}_{p}^{t^{\sharp}_{\mathsf{mix}}(\xi)}\left\|\pi\right\|_{1-1/p}^{1-1/p},$
guaranteeing finiteness of $\mathcal{J}_{p,\xi}$ for any $\xi\in(0,1)$, the above approach is only practical when bounding $\mathcal{J}_{p}^{(s)}$ for restrictively small values of $s$.
4.3   Infinite graphs with controlled growth
Natural structural assumptions on the underlying connection graph of a Markov chain also enable us to control the entropic term. Recall that an ergodic transition operator $P\in\mathcal{W}(\mathcal{X})$ induces a fully connected digraph $(\mathcal{X},\mathcal{E})$, where
$\mathcal{E}\doteq\left\{(x,x^{\prime})\in\mathcal{X}^{2}\colon P(x,x^{\prime})>0\right\}$
is the set of directed edges. The growth of a graph (Trofimov, 1985; Campbell et al., 2022) is defined as a function $g\colon\mathbb{N}\to\mathbb{N}\cup\left\{\infty\right\}$, where for $s\in\mathbb{N}$,
$g(s)\doteq\sup_{\mathcal{S}\subset\mathcal{X},\rho\left((\mathcal{X},\mathcal{E})|_{\mathcal{S}}\right)\leq s}\left|\mathcal{S}\right|,$
where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the of a directed graph, which is the maximum eccentricity of any vertex in the graph [^6] With the convention that a disconnected graph has infinite diameter., $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.
Condition 4 (Connection graph with polynomial growth).
A graph is said to have at most polynomial growth when there exists $G\in\mathbb{R}_{+}$ and $q\geq 1$ such that for any $s\in\mathbb{N}$, it holds that $g(s)\leq Gs^{q}$. Then $q$ and $G$ are respectively called the degree and factor of growth.
Lemma 4.5.
Let $p,q\in[1,\infty),\xi\in(0,1)$ and let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. If the connection graph of $P$ has polynomial growth of degree $q$ with factor $G$, it holds that
$\mathcal{J}_{p,\xi}\leq G^{1/p}t^{\sharp}_{\mathsf{mix}}(\xi)^{q/p}\left\|\pi\right\|^{1-1/p}_{1-1/p}.$
Proof.
For $x\in\mathcal{X}$, let us denote $\mathcal{S}_{x}^{(s)}\subset\mathcal{X}$ the set of states accessible from $x$ in $s$ steps. We have that
$\begin{split}\mathcal{J}_{p}^{(s)}&=\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}P^{s}(x,x^{\prime})^{1-1/p}\leq\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{S}_{x}^{(s)}}\left(\frac{1}{\mathcal{S}_{x}^{(s)}}\right)^{1-1/p}\\ &=\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}{\mathcal{S}_{x}^{(s)}}^{1/p}\leq\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}(s^{q}G)^{1/p},\\ \end{split}$
where the first inequality stems from the uniform distribution maximizing the $\ell_{1-1/p}$ quasi-norm, and the second inequality follows from the polynomial growth assumption. ∎
Observe that
$\lim_{p\to\infty}\left\|\pi\right\|_{1-1/p}^{1-1/p}=\lim_{p\to\infty}\left[t^{\sharp}_{\mathsf{mix}}(\xi)\right]^{q/p}=1,$
that is, for large values of $p$, the upper bound of Lemma 4.5 still approaches unity.
Example 4.1 (Linear growth and stationary distribution with polynomially decaying tail).
We let $\mathcal{X}=\mathbb{N}$, and suppose that there exists $C\in\mathbb{R}_{+}$, and $\eta\in(1/8,1/4)$ such that $\pi(x)\leq\frac{C}{x^{1+\eta}}$ for any $x\in\mathcal{X}$. Taking $p=2(1+1/\eta)$, it holds that
$\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\leq C^{1-1/p}\sum_{x\in\mathcal{X}}\frac{1}{x^{(1+\eta)(1-1/p)}}=C^{1-1/p}\sum_{x\in\mathcal{X}}\frac{1}{x^{1+\eta/2}}=C^{2\frac{\eta+2}{\eta+1}}\zeta(1+\eta/2),$
where $\zeta$ is the Riemann Zeta function. When the connection graph of $P$ enjoys linear growth, that is $q=1$, an application of Theorem 3.4, yields that for $\delta\in(0,1/2)$, a trajectory length of
$n\geq\widetilde{\mathcal{O}}\left(\frac{t_{\mathsf{mix}}^{11/10}}{\delta^{2}}\right),$
is sufficient to obtain $\widehat{t_{\mathsf{mix}}}^{\sharp}(2\delta/5)\geq t^{\sharp}_{\mathsf{mix}}(\delta/2)$ with probability at least $1-\delta$.
4.4   General finite space
When the state space $\mathcal{X}$ is finite, the graph-related methods discussed in the previous section remain applicable. Alternatively, a more direct approach yields the following bound,
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left(\frac{1}{\left|\mathcal{X}\right|^{2}}\right)^{1-1/p}=\left|\mathcal{X}\right|^{2/p},$ (10)
agnostic of the mixing properties of $P$. Note that (10) matches the heuristic lower bound $\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}$ in the worst-case where $P$ is bistochastic.
Corollary 4.2 (of Theorem 3.4 for $\left|\mathcal{X}\right|<\infty$).
Let $\mathcal{X}$ be a finite space. Let $\varepsilon\in(0,1/4)$, and $\delta,\xi\in(0,1)$. Let $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (6). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$. When
$n\geq 1+\frac{2^{8}et_{\mathsf{mix}}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+2\log_{2}\left|\mathcal{X}\right|\right),9\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+1\right)\log\left(\frac{4t_{\mathsf{mix}}}{\delta}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil\right)\right\},$
with probability at least $1-\delta$, it holds that
$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
Proof.
We plug (10) in Theorem 3.4 and optimize for $p$. The function $p\mapsto p\left|\mathcal{X}\right|^{2/p}$ is minimized for $p=2\log\left|\mathcal{X}\right|$, whence the corollary. ∎
4.5   The two-point space
In this section, we analyze the case where $\left|\mathcal{X}\right|=2$, illustrating the possibly stark difference between the worst-case mixing time and the average-mixing time. For $p,q\in(0,1)$, we let
$P_{p,q}=\begin{pmatrix}1-p&p\\ q&1-q\end{pmatrix}.$
Quantities such as $d$ or $d^{\sharp}$ subscripted with $p,q$ will be understood as pertaining to $P_{p,q}$. The matrix $P_{p,q}$ is irreducible, and its stationary distribution is readily given by
$\pi_{p,q}^{\intercal}=\frac{1}{p+q}\begin{pmatrix}q\\ p\end{pmatrix}.$
It follows from a direct computation (see e.g. the proof of Wolfer (2023+, Lemma 8.1)) that
$\begin{split}P_{p,q}^{t}&=1^{\intercal}\pi_{p,q}+\frac{(1-p-q)^{t}}{p+q}\begin{pmatrix}p&-p\\ -q&q\end{pmatrix},\\ \end{split}$
and as a result,
$\begin{split}d_{p,q}^{\sharp}(t)&=2\frac{pq}{(p+q)^{2}}\left|1-p-q\right|^{t},\qquad d_{p,q}(t)=\frac{p\lor q}{p+q}\left|1-p-q\right|^{t}.\end{split}$
When $p=q$, the stationary distribution is uniform, $d_{p,q}^{\sharp}(t)=d_{p,q}(t)$, and
$t_{\mathsf{mix}}(\xi)=t_{\mathsf{mix}}^{\sharp}(\xi).$
However, for $\eta\doteq p/q<1$, the ratio of the distances is a function of the parameter $\eta$,
$\frac{d_{p,q}(t)}{d^{\sharp}_{p,q}(t)}=\frac{(1+\eta)^{2}}{2\eta}\doteq r_{\eta}(t),$
whence the following lemma holds.
Lemma 4.6.
There exists a family of Markov chains
$\left\{P_{p,q}\colon(p,q)\in(0,1)^{2}\right\}\in\mathcal{W}^{\star}(\left\{0,1\right\})$
such that for any $\xi\in(0,1)$ and any $M>0$, there exist $p,q\in(0,1)$ such that
$t_{\mathsf{mix}}(\xi)>Mt_{\mathsf{mix}}^{\sharp}(\xi).$
5   Proofs
In this section, we compile technical proofs that were either outlined briefly or omitted in the manuscript. The subsequent technical tool will prove to be convenient.
Lemma 5.1.
Let $A,B\in\mathbb{R}_{+}$, with $A,B\geq e$.
$B\geq 2A\log A\implies B\geq A\log B.$
Proof.
The function $x\mapsto x/\log x$ is increasing on $[e,\infty)$ and in this range, it holds that $x/\log x\geq e>2$. Hence, for $B\geq 2A\log A$, it holds that
$\frac{B}{\log B}\geq\frac{2A\log A}{\log A+\log(2\log A)}\geq\frac{2A\log A}{\log A+\log A}=A,$
whence the lemma holds. ∎
5.1   Proof of Lemma 2.3
The proof is standard, and based on the blocking technique credited to Bernstein and further developed in the series of papers (Yu, 1993, 1994; Eberlein, 1984; Volkonskii and Rozanov, 1959). For simplicity, we assume that $n=2Bs$, where $B,s\in\mathbb{N}$. For $b\in[B]$, we write
$\begin{split}X^{[2b]}&=X^{[2b]}_{1},\dots,X^{[2b]}_{s}=X_{(2b-1)s+1},\dots,X_{2bs},\\ X^{[2b-1]}&=X^{[2b-1]}_{1},\dots,X^{[2b-1]}_{s}=X_{(2b-2)s+1},\dots,X_{(2b-1)s},\\ \end{split}$
and consider the partitioned process
$X=X^{[1]},X^{[2]},\dots,X^{[2B]}.$
In other words, we decompose the observed trajectory into $B$ even and $B$ odd blocks of size $s$, as illustrated in Figure 2.
{tikzpicture}
Figure 2: Decomposing the process into blocks.
Invoking Hoeffding’s inequality (Hoeffding, 1963) and Yu (1994, Corollary 2.7), it holds that
$\begin{split}\mathbb{P}\left(\frac{1}{n}\sum_{t=1}^{n}f(X_{t})>\varepsilon\right)&\leq\sum_{\sigma\in\left\{0,1\right\}}\mathbb{P}\left(\sum_{b=1}^{B}\sum_{t=1}^{s}f\left(X_{t}^{[2b-\sigma]}\right)>n\varepsilon/2\right)\\ &\leq 2\exp\left(-\frac{n\varepsilon^{2}}{4s}\right)+2\left(B-1\right)\beta(s).\end{split}$ (11)
We first prove the claim in the sub-exponential $\beta$-mixing setting, that is for $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ with parameters $b,\beta_{0},\beta_{1}$ in their specified range. For any $\xi\in(0,1)$, it holds that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\left\lceil\left(\frac{1}{\beta_{1}}\log\frac{\beta_{0}}{\xi}\right)^{1/b}\right\rceil.$
Observing that
$2(B-1)\beta(s)\leq\frac{n\beta_{0}}{s}\exp(-\beta_{1}s^{b}),$
and denoting $W_{0}$ the principal branch of the product logarithm Lambert function (Corless et al., 1996), for
$s=\left\lceil\left(\frac{1}{b\beta_{1}}W_{0}\left(b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b}\right)\right)^{1/b}\right\rceil,$
it holds that $22(B-1)\beta(s)\leq\delta/$ 2. As a consequence, the deviation probability at (11) is bounded by $\delta$ when the trajectory length $n$ satisfies the inequality,
$n\geq\frac{4}{\varepsilon^{2}}\left\lceil\left(\frac{1}{b\beta_{1}}W_{0}\left(b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b}\right)\right)^{1/b}\right\rceil\log\left(\frac{4}{\delta}\right).$
We now give an approximate albeit more tractable sufficient condition on $n$. Since $b\beta_{1}\leq 1$, for the above argument of $W_{0}$ to be no smaller than $e$, it is enough to have,
$n\geq\frac{\delta}{2\beta_{0}}\left(\frac{e}{b\beta_{1}}\right)^{1/b}.$
Since $\beta_{0}\geq 1,b\leq 1$ and $x\leq\frac{1}{e}\implies x\leq\log\frac{1}{x}$,
$\frac{\delta}{2\beta_{0}}\left(\frac{e}{b\beta_{1}}\right)^{1/b}\leq\frac{e}{2}\log\left(\frac{e\beta_{0}}{\delta}\right)\left(\frac{e}{b\beta_{1}}\right)^{1/b}\leq\frac{e}{2}\left(\frac{e}{b\beta_{1}}\log\left(\frac{e\beta_{0}}{\delta}\right)\right)^{1/b}\leq\frac{e}{2}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}(\delta/e).$
Therefore,
$n\geq\frac{e}{2}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}(\delta/e),$
is sufficient. Further relying on the upper bounds
$x\geq e\implies W_{0}(x)\leq\log(x)\;\text{and}\;x\geq 1\implies\lceil x\rceil\leq 2x,$
it suffices for $n$ to satisfy
$n^{b}\geq\left(\frac{8}{\varepsilon^{2}}\log\left(\frac{4}{\delta}\right)\right)^{b}\frac{1}{b\beta_{1}}\log\left(b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b}\right).$
Setting
$A=\left[\frac{16\beta_{0}}{\delta\varepsilon^{2}}\log\left(\frac{4}{\delta}\right)\right]^{b},\qquad B=b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b},$
and invoking Lemma 5.1, it is enough that
$n\geq\frac{8}{\varepsilon^{2}}\log\left(\frac{4}{\delta}\right)2^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right),$
with $\xi(\varepsilon,\delta)=\frac{\delta\varepsilon^{2}}{16\log\left(\frac{4}{\delta}\right)}$. In order to streamline this upper bound, observe that
$\frac{1}{\beta_{1}}\log\left(\frac{\beta_{0}}{\xi(\varepsilon,\delta)}\right)\leq t^{\sharp}_{\mathsf{mix}}(\delta/e)+t^{\sharp}_{\mathsf{mix}}(\delta/e)\log\left(\frac{16\log(4/\delta)}{e\varepsilon^{2}}\right),$
and recall that for $b\in(0,1]$
$x,y\in\mathbb{R}_{+}\implies(x+y)^{1/b}\leq\frac{1}{b}\left(x^{1/b}+y^{1/b}\right).$
Finally, let us now move on to the geometric case, that is,
$\beta(s)\leq\frac{\beta_{1}}{s^{b}},$
for some $\beta_{1},b>0$. For any $\xi\in(0,1)$, it holds that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\left\lceil\left(\frac{\beta_{1}}{\xi}\right)^{1/b}\right\rceil.$
Having $\beta_{1}\geq 1$ ensures that $\frac{2n\beta_{1}}{\delta}\geq 1$. Setting
$s=\left\lceil\left(\frac{2n\beta_{1}}{\delta}\right)^{\frac{1}{b+1}}\right\rceil,$
yields
$\frac{n}{s}\beta(s)\leq\frac{n\beta_{1}}{s^{b+1}}\leq n\beta_{1}\left(\frac{\delta}{2n\beta_{1}}\right)=\frac{\delta}{2}.$
Finally, setting
$n\geq\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(\frac{2\beta_{1}}{\delta}\right)^{\frac{1}{b}}$
leads to
$2\exp\left(-\frac{n\varepsilon^{2}}{4s}\right)\leq 2\exp\left(-\frac{n\varepsilon^{2}}{8\left(\frac{2n\beta_{1}}{\delta}\right)^{\frac{1}{b+1}}}\right)\leq\frac{\delta}{2}.$
For both the sub-exponential and the geometric case, an adaption of the proof handles the general case where $n$ is not divisible by the block size $s$, modulo paying a universal multiplicative constant. ∎
5.2   Proof of Lemma 2.1
The proof closely follows the techniques in Levin et al. (2009, Lemma 12.2, Theorem 12.4), albeit in a countably infinite alphabet setting. The linear operator defined by $P$ is self-adjoint in $\ell_{2}(\pi)$, hence normal. The assumption on compactness of $P$ thus enables diagonalization methods. Suppose that $P$ is reversible. We can write,
$\begin{split}d^{\sharp}(t)&\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|\delta_{x}P^{t}-\pi\right\|_{\mathsf{TV}}\\ &=\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\left|P^{t}(x,x^{\prime})-\pi(x^{\prime})\right|\\ &=\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\pi(x^{\prime})\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\\ &\stackrel{{\scriptstyle(\spadesuit)}}{{\leq}}\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\pi(x^{\prime})\frac{(1-\gamma_{\star})^{t}}{\sqrt{\pi(x)\pi(x^{\prime})}}\\ &=\frac{1}{2}(1-\gamma_{\star})^{t}\sum_{x,x^{\prime}\in\mathcal{X}}{\sqrt{\pi(x)\pi(x^{\prime})}}\\ &=\frac{1}{2}(1-\gamma_{\star})^{t}\left\|\pi\right\|_{1/2}.\\ \end{split}$
For the inequality ($\spadesuit$), we used compactness of $P$ and the fact Levin et al. (2009, Lemma 12.2) that for the $\langle\cdot,\cdot\rangle_{\pi}$-orthonormal basis of real-valued eigenfunctions $(f_{i})_{x\in\mathcal{X}}$,
$\begin{split}P^{t}(x,x^{\prime})&=\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x^{\prime})\pi(x^{\prime})\lambda^{t}_{i}f_{i}(x)=\pi(x^{\prime})+\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}(x^{\prime})\pi(x^{\prime})\lambda^{t}_{i}f_{i}(x)\\ \delta_{x}&=\sum_{i=1}^{\left|\mathcal{X}\right|}\langle\delta_{x},f_{i}\rangle_{\pi}f_{i}=\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x)\pi(x)f_{i}\end{split}$
$\pi(x)=\langle\delta_{x},\delta_{x}\rangle_{\pi}=\Biggl{\langle}\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x)\pi(x)f_{i},\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x)\pi(x)f_{i}\Biggr{\rangle}_{\pi}=\pi(x)^{2}\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x)^{2},$
thus as in Levin et al. (2009, Theorem 12.4),
$\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\leq\left|\lambda_{\star}\right|^{t}\sqrt{\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x)\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x^{\prime})}.$
Solving for a proximity parameter $\xi\in(0,1)$, we obtain that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{2\xi}.$
5.3   Proof of Theorem 3.1
For $s\in\mathbb{N}$,
$\begin{split}2\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|=&\mathbb{E}_{\pi}\left|\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|-\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{s}-\pi\right\|_{1}\right|\\ \leq&\mathbb{E}_{\pi}\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N_{xx^{\prime}}^{(s)}}{\lfloor(n-1)/s\rfloor}-\pi(x)P^{s}(x,x^{\prime})\right|\\ &+\mathbb{E}_{\pi}\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|.\end{split}$
We further decompose the argument in the second expectation as
$\begin{split}&\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|\\ &\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\frac{N^{(s)}_{x}\pi(x^{\prime})}{\lfloor(n-1)/s\rfloor}\right|+\left|\frac{N^{(s)}_{x}\pi(x^{\prime})}{\lfloor(n-1)/s\rfloor}-\pi(x)\pi(x^{\prime})\right|\\ &=2\sum_{x\in\mathcal{X}}\left|\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}-\pi(x)\right|.\\ \end{split}$
Noting that by stationarity,
$\mathbb{E}_{\pi}\left[\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}\right]=\pi(x),\qquad\mathbb{E}_{\pi}\left[\frac{N_{xx^{\prime}}^{(s)}}{\lfloor(n-1)/s\rfloor}\right]=\pi(x)P^{s}(x,x^{\prime}),$
it follows from Jensen’s inequality that
$\begin{split}2\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq\frac{1}{\lfloor(n-1)/s\rfloor}\left(\sum_{x,x^{\prime}\in\mathcal{X}}\sqrt{\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]}+2\sum_{x\in\mathcal{X}}\sqrt{\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]}\right).\end{split}$ (12)
We proceed to bound the two variance terms. We first treat $\operatorname{Var}_{\pi}\left[N_{x}\right]$ for a fixed $x\in\mathcal{X}$. This will in turn yield bounds for general $\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]$ and $\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]$. The sequence $(\delta\left[X_{t}=x\right])_{t\in\mathbb{N}}$ being stationary at second order, we expand the variance of the sum (see e.g. Rio (1999, (1.6))) as,
$\operatorname{Var}_{\pi}\left[N_{x}\right]=(n-1)\operatorname{Var}_{\pi}\left[\delta\left[X_{1}=x\right]\right]+2\sum_{t=1}^{n-2}(n-1-t)\operatorname{Cov}_{\pi}(\delta\left[X_{1}=x\right],\delta\left[X_{1+t}=x\right]).$ (13)
From Rio (1993, Theorem 1) [^7] This generalization of Ibragimov’s covariance inequality (Ibragimov, 1962) also appears in Davydov (1968) but with a weaker constant. (see also Rio (1999, (1.12b)) and Doukhan (2012, Theorem 3)), for a Hölder triplet $p,r,q\geq 1,\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$ and any $t\in\mathbb{N}\cup\left\{0\right\}$,
$\begin{split}\operatorname{Cov}_{\pi}(\delta\left[X_{1}=x\right],\delta\left[X_{1+t}=x\right])&\leq 2(2\alpha(t))^{1/p}\left\|\delta\left[X_{1}=x\right]\right\|_{\pi,r}\left\|\delta\left[X_{t}=x\right]\right\|_{\pi,q}\\ &\leq 2\beta(t)^{1/p}\pi(x)^{1-1/p},\end{split}$
where $\alpha$ is Rosenblatt’s strong mixing coefficient [^8] We note that $\alpha$ is defined as twice Rosenblatt’s coefficient in Rio (1999). (Rosenblatt, 1956), we wrote $\left\|X\right\|_{\pi,r}=\left(\mathbb{E}_{\pi}\left|X\right|^{r}\right)^{1/r}$ for a real random variable $X$, and for the second inequality, we relied on the inequality $2\alpha(t)\leq\beta(t)$ (see e.g. (Bradley, 2005, (1.11))). Plugging in (13), it follows that
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}\right]&\leq 2(n-1)\beta(0)^{1/p}\pi(x)^{1-1/p}+4\sum_{t=1}^{n-2}(n-1-t)\beta(t)^{1/p}\pi(x)^{1-1/p}\\ &\leq 4(n-1)\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{n-2}\beta(t)^{1/p}\right),\end{split}$
and for a general $s<n$, noting that the mixing coefficient $\beta^{(s)}(t)$ pertaining to the $s$-skipped chain satisfies $\beta^{(s)}(t)=\beta(st)$, we obtain
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]&\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p}\right)\\ &\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor\pi(x)^{1-1/p}\mathcal{B}^{(s)}_{p}.\\ \end{split}$
We verify (see e.g. Wolfer and Kontorovich (2021, Lemma 6.1)) that the process consisting of a sliding window over pairs,
$\{E_{t}\}_{t\in\mathbb{N}}\doteq\left\{\left(X_{t},X_{t+1}\right)\right\}_{t\in\mathbb{N}}$
forms a Markov chain over the state space =$\mathcal{E}=\left\{x,x^{\prime}\in\mathcal{X}\colon P(x,x^{\prime})>0\right\}$, called the Hudson expansion (Kemeny and Snell, 1983; Wolfer and Watanabe, 2022) of the original chain, with transition operator
$\overline{P}(e=(x_{1},x_{2}),e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime}))=\delta\left[x_{2}=x_{1}^{\prime}\right]P(x_{2},x_{2}^{\prime}),$
and with stationary distribution $\overline{\pi}((x,x^{\prime}))=\pi(x)P(x,x^{\prime})$. Writing $\overline{\beta}(s)$ for the $\beta$-mixing coefficient of $\overline{P}$, we compute
$\begin{split}\overline{\beta}(0)&=1-\left\|\overline{\pi}\right\|_{2}^{2}\\ \overline{\beta}(1)&=\beta(0)=1-\left\|\pi\right\|_{2}^{2},\end{split}$
and it follows from the Markov property that for $t\in\mathbb{N},t<n-1$,
$\overline{P}^{t+1}((x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime}))=P(x_{1}^{\prime},x_{2}^{\prime})P^{t}(x_{2},x_{1}^{\prime}),$
and thus, for $t\geq 1$,
$\begin{split}\overline{\beta}(t+1)&=\frac{1}{2}\sum_{e=(x_{1},x_{2})\in\mathcal{E}}\overline{\pi}(e)\sum_{e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime})\in\mathcal{E}}\left|\overline{P}^{t+1}(e,e^{\prime})-\overline{\pi}(e^{\prime})\right|\\ &=\frac{1}{2}\sum_{(x_{1},x_{2})\in\mathcal{E}}\pi(x_{1})P(x_{1},x_{2})\sum_{(x^{\prime}_{1},x^{\prime}_{2})\in\mathcal{E}}\left|P^{t}(x_{2},x_{1}^{\prime})-\pi(x_{1}^{\prime})\right|P(x_{1}^{\prime},x_{2}^{\prime})\\ &=\frac{1}{2}\sum_{x_{2}\in\mathcal{X}}\pi(x_{2})\sum_{x^{\prime}_{1}\in\mathcal{X}}\left|P^{t}(x_{2},x_{1}^{\prime})-\pi(x_{1}^{\prime})\right|=\beta(t).\end{split}$
More generally, for $s\in\mathbb{N},t\in\mathbb{N}\cup\left\{0\right\}$, $1\leq s,\leq n-1$, and $0\leq t<\lfloor(n-1)/s\rfloor$,
$\overline{\beta}^{(s)}(t+1)=\beta^{(s)}(t)=\beta(st).$
As a consequence
$\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor Q^{(s)}(x,x^{\prime})^{1-1/p}\left(\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\overline{\beta}^{(s)}(t)^{1/p}\right),$
where we recall that $Q^{(s)}(x,x^{\prime})=\pi(x)^{s}P(x,x^{\prime})$. We further bound
$\begin{split}\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\overline{\beta}^{(s)}(t)^{1/p}&=\frac{1}{2}\left(1-\left\|Q^{(s)}\right\|_{2}^{2}\right)+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\beta(s(t-1))^{1/p}\\ &\leq\frac{1}{2}\left(1-\left\|\pi\right\|_{2}^{2}\right)+\sum_{t=0}^{\lfloor(n-1)/s\rfloor-3}\beta(st))^{1/p}\\ &\leq\frac{3}{2}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-3}\beta(st)^{1/p}\leq\frac{3}{2}\mathcal{B}_{p}^{(s)},\end{split}$
and plugging in (12),
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}}\left(\sqrt{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\frac{3}{2}\mathcal{B}_{p}^{(s)}}+2\sqrt{\left\|\pi\right\|_{1-1/p}^{1-1/p}\mathcal{B}_{p}^{(s)}}\right),\end{split}$
which yields the first inequality in the theorem. The second inequality can be obtained from observing that $\left\|Q^{(s)}\right\|_{1-1/p}\geq\left\|\pi\right\|_{1-1/p}$. ∎
5.4   Proof of Corollary 3.1
Let us first analyze the exponential case, by assuming that $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s)$.
$\begin{split}\mathcal{B}_{p}^{(s)}=\beta_{0}^{1/p}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\exp\left(-\beta_{1}st/p\right)&\leq\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}.\end{split}$
We proceed to address the sub-exponential case. Let us assume that $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$ for $\beta_{0}>0,\beta_{1}>0$ and $b\in(0,1)$. Then,
[…]
Finally, in the polynomial case, where for $s\geq 1$, $\beta(s)\leq\beta_{1}/s^{b}$,
$\begin{split}\mathcal{B}_{p}^{(s)}&\leq\beta_{0}^{1/p}+\beta(s)^{1/p}\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\frac{1}{t^{b/p}}\leq\beta_{0}^{1/p}+\zeta(b/p)\beta(s)^{1/p}.\\ \end{split}$
5.5   Proof of Theorem 3.2
Recall that for a uniformly ergodic Markov chain,
$\beta(s)\leq d(s)\leq d(\lfloor s/t_{\mathsf{mix}}\rfloor t_{\mathsf{mix}})\leq\exp(-\lfloor s/t_{\mathsf{mix}}\rfloor\log 2)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$
where the third inequality is adapted from Levin et al. (2009, (4.33)), i.e. $\beta$ is decays exponentially with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. Thus, from the fact that $x>0\implies 1/(1-\exp(-x))\leq 1+1/x,$ and Corollary 3.1, it holds that
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{2^{1/p}}{\lfloor(n-1)/s\rfloor}\left(1+\frac{pt_{\mathsf{mix}}}{s\log 2}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}},\\ \end{split}$
and $(i)$ follows from $n>s+1$ and $x>1\implies\lfloor x\rfloor\geq x/2$ and $2^{1/p}\leq 2$. From $(i)$, when
$n\geq 1+\frac{2^{8}}{\varepsilon^{2}}\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
it holds that, $\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon/2$ and in this case, the probability of the above deviation can be upper bounded by fluctuations around the mean as follows,
$\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)\leq\mathbb{P}_{\pi}\left(\left|\left|\widehat{\beta}(s)-\beta(s)\right|-\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\right|>\varepsilon/2\right).$
Let
$x_{[s]}^{n}=x_{1},x_{1+s},x_{1+2s},\dots,x_{1+\lfloor(n-1)/s\rfloor s}\in\mathcal{X}$
be some $s$-skipped trajectory, let $t\in\left\{0,\dots,\lfloor(n-1)/s\rfloor\right\}$, $x^{\prime}_{1+ts}\in\mathcal{X}$, and denote
$x^{n,(t)}=x_{1},x_{1+s},x_{1+2s},\dots,x_{1+(t-1))s},x^{\prime}_{1+ts},x_{1+(t+1)s}\dots,x_{1+\lfloor(n-1)/s\rfloor s}$
the trajectory where the observation $x_{1+ts}$ at time $1+ts$ was replaced with observation $x^{\prime}_{1+ts}$. We show that the estimator satisfies the bounded differences condition. We decompose,
$\begin{split}&\left|\left|\widehat{\beta}(s)(X)-\beta(s)\right|-\left|\widehat{\beta}(s)(X^{(t)})-\beta(s)\right|\right|\\ &\leq\left|\widehat{\beta}(s)(X)-\widehat{\beta}(s)(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}(X)-N_{xx^{\prime}}^{(s)}(X^{(t)})\right|\\ &+\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|.\\ \end{split}$
We further upper bound the second term as follows,
$\begin{split}&\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &+\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X^{(t)})-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x^{\prime}\in\mathcal{X}}\left|N_{x^{\prime}}^{(s)}(X)-N_{x^{\prime}}^{(s)}(X^{(t)})\right|+\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x\in\mathcal{X}}\left|N_{x}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})\right|.\\ \end{split}$
We obtain
$\left|\left|\widehat{\beta}(s)(X)-\beta(s)\right|-\left|\widehat{\beta}(s)(X^{(t)})-\beta(s)\right|\right|\leq\frac{4}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}.$
From a version of McDiarmid’s inequality for uniformly ergodic Markov chains (Paulin, 2015, Corollary 2.10), and the fact that the mixing time of the $s$-skipped chain $t_{\mathsf{mix}}^{(s)}$ verifies $t_{\mathsf{mix}}^{(s)}\leq\lceil t_{\mathsf{mix}}/s\rceil$,
$\begin{split}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)&\leq 2\exp\left(-\frac{2(\varepsilon/2)^{2}}{9t_{\mathsf{mix}}^{(s)}\lfloor(n-1)/s\rfloor\left(\frac{8}{\lfloor(n-1)/s\rfloor}\right)^{2}}\right)\\ &\leq 2\exp\left(-\frac{(n-1)\varepsilon^{2}}{9\cdot 2^{8}(t_{\mathsf{mix}}+s)}\right),\end{split}$
which completes the proof of claim $(ii)$. ∎
5.6   Proof of Theorem 3.3
Suppose that
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},18\log\left(\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\right)\right\}.$
Invoking Lemma 5.1 with
$a\doteq\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon),\qquad b\doteq\frac{2}{\delta_{n-1}},$
and where $\delta_{n-1}\doteq\frac{\delta}{n-1}$, it holds that
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}.$ (14)
Writing $S=S(n,\varepsilon,\delta,p,\mathcal{J}_{p})$, we thus have
$\begin{split}n-1&\geq\frac{n-1}{2}+\frac{2^{8}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}\\ &\geq\frac{2^{8}}{\varepsilon^{2}}\left[S+\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\right]\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}\\ &\geq\frac{2^{8}}{\varepsilon^{2}}\max\left\{\left(S+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\mathcal{J}_{p},9(S+t_{\mathsf{mix}})\log\frac{2}{\delta_{n-1}}\right\}.\end{split}$
Furthermore, observing from (14) that
$S\geq\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\geq t^{\sharp}_{\mathsf{mix}}(\varepsilon),$
it follows that $\beta(s)<\varepsilon$ for $s>S$. An application of the union bound and Theorem 3.2$(ii)$ yield that
$\mathbb{P}_{\pi}\left(\left\|\widehat{\beta}-\beta\right\|_{\infty}>\varepsilon\right)\leq\sum_{s=1}^{S}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)\leq S\delta_{n-1}\leq\delta.$
The claim then follows from $t_{\mathsf{mix}}(\varepsilon)\leq\lceil\log_{2}\varepsilon^{-1}\rceil t_{\mathsf{mix}}$ for any $\varepsilon\in(0,1/4)$ (Levin et al., 2009, (4.34)). ∎
5.7   Proof of Theorem 3.4
We first bound the probability that the estimator overshoots, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)>t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$. On this event, for any $s\leq t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, and in particular for $s=t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, it must be that $\widehat{\beta}(s)>\xi$. From Theorem 3.2$(ii)$, for
$n\geq 1+\frac{2^{8}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{4}{\delta}\right\},$
it holds that
$\begin{split}\mathbb{P}_{\pi}\left(\widehat{\beta}(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))>\xi\right)&\leq\mathbb{P}_{\pi}\left(\widehat{\beta}(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))-\beta(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))>\xi-\xi(1-\varepsilon)\right)\\ &\leq\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))-\beta(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))\right|>\xi\varepsilon\right)\leq\frac{\delta}{2}.\end{split}$
As a second step, we bound the probability that $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ underestimates the true parameter, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)<t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon))$. From Theorem 3.2$(ii)$ and an application of the union bound, when for any $s\leq t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))$ the trajectory length satisfies
$n\geq 1+\frac{2^{8}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{4t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))}{\delta}\right\},$
it holds that
$\begin{split}\mathbb{P}_{\pi}\left(\exists s<t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon)),\widehat{\beta}(s)\leq\xi\right)&\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\widehat{\beta}(s)\leq\xi\right)\\ &\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\beta(s)-\widehat{\beta}(s)\geq-\xi+\xi(1+\varepsilon)\right)\\ &\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|\geq\xi\varepsilon\right)\leq\frac{\delta}{2}.\end{split}$
The theorem follows from $t^{\sharp}_{\mathsf{mix}}(\xi(1-\varepsilon))\leq t_{\mathsf{mix}}(\xi(1-\varepsilon))\leq t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\xi(1-\varepsilon)}\right\rceil\leq t_{\mathsf{mix}}\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil$. ∎
6   Example
In this section, we identify $\mathcal{X}=\mathbb{N}$, and we study the properties of a family of transition operators $(P_{T})_{T\in\mathcal{T}}$.
Definition 6.1.
Let $q\in(0,1)$, a sequence $\nu=(\nu_{x})_{x=1,\dots}$ taking values in $[0,1]$, and a probability distribution $\mu=(\mu_{x})_{x=2,\dots}$. We define:
$P_{T}=\left(\begin{array}[]{c c c c c}1-q&q\mu_{2}&q\mu_{3}&q\mu_{4}&\dots\\ \nu_{2}&1-\nu_{2}&0&0&\dots\\ \nu_{3}&0&1-\nu_{3}&0&\dots\\ \nu_{4}&0&0&1-\nu_{4}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{array}\right).$
We define $\mathcal{T}$ as the set of triples $T=(q,\mu,\nu)$ such that $q>0$, $\nu_{x}>0$ and $\mu_{x}>0$ for all $x$ and $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<+\infty$.
A first observation is that when $q>0$ and $\nu_{x}>0$ and $\mu_{x}>0$ for all $x$, then the chain is aperiodic and irreducible. Thus it has an invariant positive measure (not necessarily finite). Solving $\pi P_{T}=\pi$, we find that the invariant measure satisfies, for $x\geq 2$,
$\pi(x)=\frac{q\mu_{x}}{\nu_{x}}\pi(1).$
So, if $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}=+\infty$, the chain is recurrent null. But if $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<\infty$ then it is recurrent positive, and the unique invariant probability measure, that will be denoted by $\pi_{T}$, is given by
$\pi_{T}(1)=\frac{1}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\text{ and }\pi_{T}(x)=\frac{q\mu_{x}}{\nu_{x}\left(1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}\right)}\text{ for }x>1.$ (15)
Thus, when $T=(q,\mu,\nu)\in\mathcal{T}$ the chain is recurrent positive. At first, we will study the ergodic properties of $P_{T}$ and evaluate its $\beta$-mixing coefficients. Secondly, we will apply the results of Section 4 to control $\mathcal{J}_{p}^{(s)}$ for $P_{T}$.
6.1   Ergodic and mixing properties of $P_{T}$
We remind that the Dobrushin coefficient of $P_{T}$ is defined by
$\kappa:=\sup_{x,y\in\mathbb{N}}\|P_{T}(x,\cdot)-P_{T}(y,\cdot)\|_{{\rm TV}}.$ (16)
An application of Theorem 18.2.4 in Douc et al. (2018) gives, for any $t\in\mathbb{N}$,
$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\leq\kappa^{t}.$ (17)
Thus, when $\kappa<1$, the chain is uniformly ergodic.
Note for example that if $q<1$ and $\nu_{x}\geq 1-q$, then for any $x$ we have $P_{T}(x,1)\geq 1-q$. Thus, for any $x$ and $y$, $\|P_{T}(x,\cdot)-P_{T}(y,\cdot)\|_{{\rm TV}}\leq q<1$, and thus $\kappa<q<1$. So the chain defined by $P_{T}$ is uniformly ergodic. In this case, $t_{\mathsf{mix}}(\xi)\leq\lceil\log(\xi)/\log q\rceil$ and the chain is exponentially $\beta$-mixing.
A first interesting observation is that when $\nu_{x}\rightarrow 0$, the chain cannot be uniformly ergodic.
Lemma 6.1.
Assume $T=(q,\mu,\nu)\in\mathcal{T}$. If $\nu_{x}\rightarrow 0$ for $x\rightarrow\infty$, we have
$d(t)=\sup_{x\geq 1}\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\geq\frac{1}{2}$
and thus $t_{{\rm mix}}=+\infty$.
Proof: we have, for a fixed $t$,
$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\sum_{y=1}^{\infty}|P_{T}^{t}(x,y)-\pi_{T}(y)|$
$\geq\frac{1}{2}|P_{T}^{t}(x,x)-\pi_{T}(x)|$
$\geq\frac{1}{2}|(1-\nu_{x})^{t}-\pi_{T}(x)|.$
Note that $\pi_{T}(x)\rightarrow 0$ for $x\rightarrow\infty$, while $(1-\nu_{x})^{t}\rightarrow 1$. $\square$
Of course, even though it is not uniformly ergodic, as the chain remains ergodic for $T\in\mathcal{T}$. Thus, for $t\rightarrow+\infty$ we have $\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\rightarrow 0$. Thus, by the dominated convergence theorem, in the stationary case, we have $\beta(t)=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\rightarrow 0.$ We can actually provide more accurate information on $\beta(t)$. This requires a careful analysis that will be simplified by additional technical assumptions on $T$.
Definition 6.2.
We define $\mathcal{S}\subset\mathcal{T}$ as the set of triples $T=(q,\mu,\nu)\in\mathcal{T}$ such that $q\in(0,1/4]$, $(\nu_{x})_{x=1,\dots}$ is nonincreasing and takes values in $(0,1/2]$ and $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<+\infty$.
Lemma 6.2.
Assume $T=(q,\mu,\nu)\in\mathcal{S}$. For any $K$ large enough to ensure $\sum_{x>K}\pi_{T}(x)\leq 1/2$,
$\beta(t)\leq 2(1-\nu_{K})^{t}+7\sum_{x>K}\pi_{T}(x)+2tq\sum_{x>K}\mu_{x}.$
We can simplify a little the above by noting that
$q\mu_{x}=\pi_{T}(x)\nu_{x}\left(1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}\right)\leq\frac{\pi_{T}(x)\nu_{x}}{\pi_{T}(1)}\leq\frac{\pi_{T}(x)}{2\pi_{T}(1)}$
and thus
$\beta(t)\leq 2(1-\nu_{K})^{t}+\left(7+\frac{t}{\pi_{T}(1)}\right)\sum_{x>K}\pi_{T}(x).$
Before we provide the proof, let us discuss various possible regimes:
  • if $\nu_{i}x\rightarrow 0$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ for some $C^{\prime},b>0$ then there is a $C>0$ such that for $x>1$, $\pi(x)=C\exp(-bx)$ and thus
    $\beta(t)$ $\leq 2(1-\nu_{K})^{t}+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}$
    $\leq 2\exp(-t\nu_{K})+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}.$
    For example if $\nu_{x}=\min(1/2,1/\log\log(x))$ we can take $K=t$ to obtain:
    $\beta(t)\leq 2\exp\left(-\frac{t}{\log\log(t)}\right)(1+o(1)).$
  • if $\nu_{x}=\min(1/2,1/x^{a})$ for some $a>0$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ for some $C^{\prime}>0$ and $b>1$ then there is a $C>0$ such that for $x>1$, $\pi_{T}(x)=C\exp(-bx)$ and thus
    $\beta(t)\leq 2\exp\left(-\frac{t}{K^{a}}\right)+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}$
    which, if we choose $K=\lfloor t^{\frac{1}{a+1}}\rfloor$ we obtain
    $\beta(t)\leq\frac{\frac{Ct}{\pi_{T}(1)}\exp(-bt^{\frac{1}{a+1}})}{1-\exp(-b)}(1+o(1)).$
  • if $\nu_{x}=\min(1/2,1/x)$ and $\mu_{x}=C^{\prime}\nu_{x}/x^{b}$ for some $C^{\prime}>0$ and $b>1$ then there is a $C>0$ such that for $x>1$, $\pi_{T}(x)=C/x^{b+1}$ and thus there is a $C^{\prime\prime}$ with
    $\beta(t)\leq 2\exp(-t/K)+C^{\prime\prime}\left(7+\frac{t}{\pi_{T}(1)}\right)\frac{1}{(K+1)^{b}}$
    and the choice $K=\lfloor t/((b-1)\log t)\rfloor$ gives
    $\beta(t)$ $\leq\left(\frac{2}{t^{d}}+\frac{C^{\prime\prime}(b-1)^{b}}{\pi_{T}(1)}\frac{\log^{b}(t)}{t^{b-1}}\right)(1+o(1))$
    $=\frac{C^{\prime\prime}(b-1)^{b}}{\pi(1)}\frac{\log^{b}(t)}{t^{b-1}}(1+o(1)).$
Of course, these estimates could be seen as pessimistic. However, we also have lower bounds for $\beta(t)$.
Lemma 6.3.
Assume $T=(q,\mu,\nu)\in\mathcal{S}$. Assume $\nu_{x}=\min(1/2,1/x^{a})$ for some $a>0$. Then, we have, for $t\geq 2$,
$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[\frac{1}{4}-\pi_{T}(\lceil t^{1/a}\rceil)\right].$
That is,
$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{8}(1+o(1)).$
This is important with respect to the previous examples:
  • by taking $a=1$ thus $\nu_{x}=\min(1/2,1/x)$ and $\mu(x)=C^{\prime}\nu_{x}/x^{b}=C^{\prime}/x^{b+1}$ for some $b>1$ we have $\pi_{T}(x)=C/x^{b+1}$ and thus
    $\frac{C}{8(t+1)^{b+1}}\left[1+o(1)\right]\leq\beta(t)\leq\frac{C^{\prime\prime}(b-1)^{b}}{\pi_{T}(1)}\frac{\log^{b}(t)}{t^{b-1}}(1+o(1)).$
  • in the case $\nu_{x}=\min(1/2,1/x^{a})$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ and thus $\pi_{T}(x)=C\exp(-bx)$, this gives
    $\frac{C\exp[-b(t^{\frac{1}{a}}+1)]}{8}\left[1+o(1)\right]\leq\beta(t)\leq\frac{\frac{Ct}{\pi_{T}(1)}\exp(-bt^{\frac{1}{a+1}})}{1-\exp(-b)}(1+o(1)).$
We conclude this subsection by the proofs of Lemmas 6.2 and 6.3.
Proof of Lemma 6.2.
The proof relies on the following approximation $P_{T,K}$ of $P_{T}$:
$P_{T,K}=\left(\begin{array}[]{c c c c c c}1-q-q\sum_{x>K+1}\mu_{x}&q\mu_{2}&\dots&q\mu_{K}&0&\dots\\ \nu_{2}&1-\nu_{2}&0&0&0&\dots\\ \vdots&\vdots&\ddots&\vdots&\vdots&\dots\\ \nu_{K}&0&&1-\nu_{K}&0&\dots\\ \nu_{K+1}&0&\dots&0&1-\nu_{K+1}&\dots\\ \vdots&\vdots&&\vdots&\vdots&\ddots\end{array}\right).$
Let $\pi_{T,K}$ denote its invariant distribution:
$\pi_{T,K}(1)=\frac{1}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\text{ and }\pi_{T,K}(x)=\frac{q\mu_{x}}{\nu_{x}\left(1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}\right)}\text{ for }1<x\leq K$
and $\pi_{T,K}(x)=0$ if $x>K$. Then, note that
$\beta(t)$ $=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
$\leq\sum_{x=1}^{K}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}+\sum_{x>K}\pi_{T}(x).$
Then, for $x\leq K$, we have:
$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\\ \leq\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}+\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}+\|\pi_{T,K}(\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}.$ (18)
Let us control each term in (18) separately. First,
$2\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}$ $=\left|\frac{1}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}-\frac{1}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\right|$
$\quad\quad\quad+\sum_{x=2}^{K}\left|\frac{\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}-\frac{\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\right|+\sum_{x>K}\pi_{T}(x)$
$=\left|1-\frac{1+q\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\right|\sum_{x=1}^{K}\pi_{T}(x)+\sum_{x>K}\pi_{T}(x)$
$\leq\left|1-\frac{1+q\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\right|+\sum_{x>K}\pi_{T}(x)$
$=\frac{\sum_{x>K}\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}+\sum_{x>K}\pi_{T}(x)$
$=\frac{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\frac{\sum_{x>K}\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}+\sum_{x>K}\pi_{T}(x)$
$=\left(\frac{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}+1\right)\sum_{x>K}\pi_{T}(x)$
$=\left(\frac{1}{1-\sum_{x>K}\pi_{T}(x)}+1\right)\sum_{x>K}\pi_{T}(x)$
$\leq 3\sum_{x>K}\pi_{T}(x)$
as soon as $K$ large enough to ensure $\sum_{x>K}\pi_{T}(x)\leq 1/2$.
We then tackle the term $\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}$ in (18). In order to do so, we study the Dobrushin coefficient of $P_{T,K}$,
$\kappa:=\sup_{x,y\leq K}\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$
and show it satisfies $\kappa<1$. Exactly as we derived (17) from (16) above, an application of Theorem 18.2.4 in Douc et al. (2018) will allow to conclude. When $1<x<y$,
$\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\left(|\nu_{x}-\nu_{y}|+|1-\nu_{x}|+|1-\nu_{y}|\right)$
$=1-\nu_{y}.$
In the case $1=x<y$, and there we need $q\leq 1/4$ and $\nu_{y}\leq 1/2$,
$\|P_{T,K}(1,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\left(\left|1-q-q\sum_{x>K}\mu_{x}-\nu_{y}\right|+|1-\nu_{y}-q\mu_{y}|+\sum_{x\leq K,x\neq y}q\mu_{x}\right)$
$=\frac{1}{2}\left[2-2\nu_{z}-q\left(\sum_{x=2}^{\infty}\mu_{x}+\sum_{x>K}\mu_{x}+\mu_{y}-\sum_{x\leq K,x\neq y}\mu_{x}\right)\right]$
$\leq 1-\nu_{y}.$
So, $\kappa=\sup_{x,y\leq K}\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}=1-\nu_{K}$ which leads to the bound
$\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}\leq(1-\nu_{K})^{t}.$
Finally, we control the term $\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}$ in (18). This can be done thanks to the theory of the stability of Markov chains, that studies how fast two Markov chains with very similar transition operators will diverge, see for example Mitrophanov (2003). More precisely, (3.17) in the proof of Corollary 3.2 in Mitrophanov (2003) shows that
$\sup_{1\leq x\leq K}\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}\leq t\sup_{1\leq x\leq K}\|P_{T}(x,\cdot)-P_{T,K}(x,\cdot)\|_{{\rm TV}}.$
But it is direct to check that
$\sup_{1\leq x\leq K}\|P_{T}(x,\cdot)-P_{T,K}(x,\cdot)\|_{{\rm TV}}$ $=\|P_{T}(1,\cdot)-P_{T,K}(1,\cdot)\|_{{\rm TV}}$
$=q\sum_{x>K}\mu_{x}.$
Thus, we obtain
$\sup_{1\leq x\leq K}\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}\leq tq\sum_{x>K}\mu_{x}.$
Putting everything together in (18),
$\beta(t)\leq\sum_{x=1}^{K}2\pi_{T}(x)\left[(1-\nu_{K})^{t}+3\sum_{x>K}\pi_{T}(x)+qt\sum_{x>K}\mu_{x}\right]+\sum_{x>K}\pi_{T}(x)$
which ends the proof.
Proof of Lemma 6.3.
We have, for a fixed $t$ and $y$,
$\beta(t)$ $=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
$\geq\pi_{T}(y)\|P_{T}^{t}(y,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
$=\frac{\pi_{T}(y)}{2}\sum_{y=1}^{\infty}|P_{T}^{t}(k,y)-\pi_{T}(y)|$
$\geq\frac{\pi_{T}(y)}{2}|P_{T}^{t}(y,y)-\pi_{T}(y)|$
$\geq\frac{\pi_{T}(y)}{2}\left[(1-1/(2y^{a}))^{t}-\pi_{T}(y)\right].$
Then, take $y=\lceil t^{1/a}\rceil$ to get
$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[(1-1/t)^{t}-\pi_{T}(\lceil t^{1/a}\rceil)\right]\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[\frac{1}{4}-\pi_{T}(\lceil t^{1/a}\rceil)\right]$
when $t>2$. ∎
6.2   Control of $\mathcal{J}_{p}^{(s)}$ for $P_{T}$
Proposition 6.1.
Assume $T=(q,\mu,\nu)\in\mathcal{T}$ and
$\sum_{x\geq 2}\frac{\mu_{x}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{x}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
Then Condition 3 is satisfied for $P_{T}$ with $D(1)=V(1)=1$, $D(k)=V(k)=1/\sqrt{\mu_{k}}$ for $k>1$ and $b=q\sum_{k\geq 2}\sqrt{\mu_{k}}$. Moreover, the sumability conditions of Lemma 4.1 are also satisfied, and
$\mathcal{J}_{p}^{(s)}\leq\left\|\pi_{T}\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi_{T}\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi_{T}\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right)<\infty.$
This gives a nontrivial example of chain that satisfies Condition 3.
Proof of Proposition 6.1.
First, let us evaluate $P_{T}D(x)$. We start with $x>1$,
$P_{T}D(x)=\sum_{y}P_{T}(x,y)D(y)=\nu_{x}D(1)+(1-\nu_{x})D(x)\leq D(x),$
and then when $x=1$,
$P_{T}D(1)=\sum_{y}P_{T}(1,y)D(y)=(1-q)D(1)+q\sum_{y\geq 2}\mu_{y}D(y)\leq D(1)+q\sum_{y\geq 2}\sqrt{\mu_{y}}.$
Thus $P_{T}D(x)\leq D(x)+b$ with $b=q\sum_{y\geq 2}\sqrt{\mu_{y}}$.
We can now check the assumption in Lemma 4.3:
$\sup_{y\in\mathcal{X}}V(y)(P_{T}(x,y)-\pi_{T}(y))\leq D(x).$
We actually prove the stronger:
$\forall(x,y)\text{, }P_{T}(x,y)\leq\frac{D(x)}{V(y)}.$
Indeed, for $x=y=1$,
$P_{T}(1,1)=1-q\leq 1=\frac{D(1)}{V(1)}.$
For $x>1$,
$P_{T}(x,x)=1-\nu_{x}\leq 1=\frac{D(x)}{V(x)}.$
Still for $x>1$,
$P_{T}(x,1)=\nu_{x}\leq 1\leq\frac{1}{\sqrt{\mu_{x}}}=\frac{D(x)}{V(1)}.$
Finally, for $y>1$,
$P_{T}(1,y)=q\mu_{y}\leq\mu_{y}\leq\sqrt{\mu_{y}}=\frac{D(1)}{V(y)}.$
Finally, let us check the summability conditions:
$\sum_{y}\left[\frac{1}{V(y)}\right]^{\frac{1}{2}(1-\frac{1}{p})}=1+\sum_{y>1}\mu_{y}^{\frac{1}{4}(1-\frac{1}{p})}\leq 1+\sum_{y>1}\frac{\mu_{y}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{y}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
by assumption. Then
$\sum_{x}\left[\pi_{T}(x)D_{x}\right]^{\frac{1}{2}(1-\frac{1}{p})}\leq 1+\sum_{x>1}\left[\frac{\mu_{x}}{\nu_{x}}\frac{1}{\sqrt{\mu_{x}}}\right]^{\frac{1}{2}(1-\frac{1}{p})}\leq 1+\sum_{x>1}\frac{\mu_{x}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{x}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
still by assumption. This ends the proof. ∎
Acknowledgments
GW would like to thank Luca Zanetti and John Sylvester for posing the original question of estimating the average-mixing time, and Aryeh Kontorovich for the interesting discussions.
Funding
GW is supported by the Special Postdoctoral Researcher Program (SPDR) of RIKEN. The visit of Pierre Alquier to RIKEN in October 2023, during which part of this work was conducted was sponsored by the SPDR program.

### v2

@@
 The sample complexity lower bound in (1) is prohibitively large for vast state spaces, and even vacuous in the infinite case.
-In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in (Espuny Díaz et al., 2022; Münch and Salez, 2022).
+In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in (Díaz et al., 2024; Münch and Salez, 2023).
 Specifically, we will confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space.
@@
 (2018) analyzed random walks on the giant component of the Erdös–Rényi random graph $\mathcal{G}(n,p)$ with $p=\Theta(1/n)$, and showed that the mixing time when starting randomly from any vertex is the order $\mathcal{O}(\log n)$ while it was known to be of the order $\mathcal{O}(\log^{2}n)$ for the mixing time from a worst starting point.
-Following up, Espuny Díaz et al.
+Following up, Díaz et al.
 (2024) extended the approach on a larger collection of graphs with small bottlenecks, and show applicability of their approach for the Newman–Watts small-world model and supercritically percolated expander graphs.
 In (Lovász and Kannan, 1999), the authors introduce the notion of a conductance averaged over subsets of states with various sizes, and demonstrate that it can lead to tighter convergence bounds for counting and volume computation problems.
-Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered in (Münch and Salez, 2022) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
+Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered in (Münch and Salez, 2023) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
 In particular, we stress that the notion of average-mixing treated in this paper is different from the one discussed in Peres and Sousi (2015); Anderson et al.
@@
 $t^{\sharp}_{\mathsf{mix}}(\xi)=\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\beta(s)\leq\xi\right\},$
-we briefly demonstrate how the average-mixing time (together with an a priori known order of convergence) can be used to quantitatively capture the convergence rate of a Markov chain in lieu of the worst-case mixing time.
-Specifically, we show (Lemma 2.3) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ naturally appears in the sample complexity for controlling the deviation probability from 00 of a centered function $f\colon\mathcal{X}\to[-1,1]$.
+we briefly demonstrate how the average-mixing time, together with an a priori known order of convergence, can be used to quantitatively capture the convergence rate of a Markov chain in lieu of the worst-case mixing time.
+Specifically, we show (Lemma 2.3) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ naturally appears in the sample complexity for controlling the deviation probability from zero of a centered function $f\colon\mathcal{X}\to[-1,1]$.
 Namely, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$, it is sufficient to have a trajectory of
@@
 For a general time-discrete, time-homogeneous Markov chain over a countable state space, we establish a parametric rate of convergence for its individual $\beta$-coefficients from a single stationary trajectory of observations $X_{1},\dots,X_{n}$.
-Namely, for $s\in\mathbb{N}$, we construct an estimator $\widehat{\beta}(s)$ for $\beta(s)$, and we show (Theorem 3.1) that
+Namely, for $s\in\mathbb{N}$, we construct an estimator $\widehat{\beta}(s)$ for $\beta(s)$, and we show (refer to Theorem 3.1) that the absolute mean deviation can be controlled as follows,
 $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
@@
 $\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(3\xi/2),t_{\mathsf{mix}}^{\sharp}(\xi/2)\right]\end{split}$
-with probability at least $1-\delta$. Furthermore, for finite spaces, setting $\xi=1/4$ for simplicity, the parameter $p$ can be chosen so as to obtain an upper bound (Corollary 4.2) of
+with probability at least $1-\delta$. Furthermore, for finite spaces, setting $\xi=1/4$ for simplicity, the parameter $p$ can be chosen so as to obtain an upper bound (refer to Corollary 4.2) of
 $n\geq Ct_{\mathsf{mix}}\log\max\left\{\left|\mathcal{X}\right|,\frac{t_{\mathsf{mix}}}{\delta}\right\},$
 where $C\in\mathbb{R}_{+}$ is a universal constant.
-One can thus demonstrably estimate with high probability the average-mixing time sub-linearly in the number of states, answering a question posed by Luca Zanetti and John Sylvester (Zanetti and John, 2023).
+One can thus demonstrably estimate with high probability the average-mixing time sub-linearly in the number of states, answering a question posed by Luca Zanetti and John Sylvester.
 This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024).
@@
 Section 2 introduces the average-mixing time, illustrates its significance and analyzes some of its properties.
-In Section 2.2, under geometric ergodicity and time-reversibility assumptions, we establish an upper bound of $t^{\sharp}_{\mathsf{mix}}$ in terms of the relaxation time of the chain, which is non-vacuous over countable state spaces and under suitable conditions on the stationary distribution.
+In Section 2.2, under geometric ergodicity and time-reversibility assumptions, we establish an upper bound of $t^{\sharp}_{\mathsf{mix}}$ in terms of the relaxation time of the chain, which remains non-vacuous over countable state spaces and under suitable conditions on the stationary distribution.
 In Section 2.3 we recall some important connection between the average-mixing time and $\beta$-mixing properties of the chain.
@@
 In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain; first individually and then as a sequence, with respect to the uniform norm.
-In Section 3.3, we specialize our analysis to the uniformly ergodic setting, and in Section 3.4 we deduce an upper bound for the problem of estimating the average-mixing time from a single trajectory under uniform ergodicity.
+From Section 3.3, we specialize our analysis to the uniformly ergodic setting, and in Section 3.4 we deduce an upper bound for the problem of estimating the average-mixing time from a single trajectory.
 In Section 4, we demonstrate that our contribution holds implications at various scales of state spaces and provide more explicit rates under additional structural assumptions on the underlying chain.
@@
 We let $\mathcal{X}$ be a countably infinite space of symbols equipped (for simplicity) with a total order.
-Real functions on $\mathcal{X}$ are written as infinite row vectors and for $x\in\mathcal{X}$, $e_{x}$ is the function defined by $x^{\prime}\in\mathcal{X}$, $e_{x}(x^{\prime})=\delta\left[x=x^{\prime}\right]$.
+Real functions on $\mathcal{X}$ are written as infinite row vectors and for $x\in\mathcal{X}$, $e_{x}$ is the function defined by $x^{\prime}\in\mathcal{X}$, $e_{x}(x^{\prime})=\delta\left[x=x^{\prime}\right]$, where $\delta[\cdot]$ is the predicate function taking values in $\left\{0,1\right\}$.
 For $p\in(0,\infty)$, and a function $v\in\mathbb{R}^{\mathcal{X}}$,
@@
 We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$.
-For a finite trajectory of length $n\in\mathbb{N}$, started from an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
+For a finite trajectory of length $n\in\mathbb{N}$, started with an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
 $\mathbb{P}_{\mu}\left(X_{1}=x_{1},\dots,X_{n}=x_{n}\right)=\mu(x_{1})\prod_{t=1}^{n-1}P(x_{t},x_{t+1}).$
@@
 The above should be understood as the total variance distance of the chain to stationarity, averaged with respect to $\pi$ over the starting state.
-The limit $\lim_{t\to\infty}d^{\sharp}(t)=0$ always exists [^1] This follows from a connection between average-mixing time and $\beta$-mixing discussed at Section 2.3, thus for $\xi\in(0,1)$, the average-mixing time of $P$, defined (Münch and Salez, 2022) as
+The limit $\lim_{t\to\infty}d^{\sharp}(t)=0$ always exists [^1] This fact follows from a connection between average-mixing time and $\beta$-mixing discussed at Section 2.3., thus for $\xi\in(0,1)$, the average-mixing time of $P$, defined (Münch and Salez, 2023) as
 $\begin{split}t^{\sharp}_{\mathsf{mix}}(\xi)&\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d^{\sharp}(t)\leq\xi\right\},\end{split}$
@@
 Following their respective definitions, it holds that $d^{\sharp}(t)\leq d(t)$.
-Thus, for any fixed $\xi\in(0,1)$, average-mixing is always faster than mixing.
-The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2022, Remark 1), which in particular include random walks on groups [^2] A chain is called transitive when for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$..
+Thus, for any fixed $\xi\in(0,1)$, average-mixing occurs always faster than mixing.
+The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2023, Remark 1), which in particular include random walks on groups [^2] A chain is called transitive when for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$..
 In general however, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as formalized below.
@@
 Note that $\|\pi\|_{1/2}$ could be infinite, in which case Lemma 2.1 becomes vacuous.
-However, Lemma 2.1 compares favorably with the bound of Münch and Salez (2022, Theorem 5, Corollary 2), which is inversely proportional to the minimum (non-zero) transition probability $P_{\star}\doteq\min_{x,x^{\prime}\in\mathcal{X}}P(x,x^{\prime})$.
+However, Lemma 2.1 compares favorably with the bound of Münch and Salez (2023, Theorem 5, Corollary 2), which is inversely proportional to the minimum (non-zero) transition probability $P_{\star}\doteq\min_{x,x^{\prime}\in\mathcal{X}}P(x,x^{\prime})$.
 Additionally, Lemma 2.1 can be instructively compared with Levin et al.
@@
 where $\sigma(\{X_{t}\}_{t})$ is the $\sigma$-field generated by the random variables $\{X_{t}\}_{t}$, and for two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$,
-$\beta(\mathcal{A},\mathcal{B})\doteq\sup\frac{1}{2}\sum_{i=1}^{I}\sum_{j=1}^{J}\left|\mathbb{P}(A_{i}\cap B_{i})-\mathbb{P}(A_{i})\mathbb{P}(B_{j})\right|,$
+$\beta(\mathcal{A},\mathcal{B})\doteq\sup\frac{1}{2}\sum_{i=1}^{I}\sum_{j=1}^{J}\left|\mathbb{P}(A_{i}\cap B_{j})-\mathbb{P}(A_{i})\mathbb{P}(B_{j})\right|,$
 where the supremum is taken over all pairs of finite partitions $\{A_{1},\dots,A_{I}\}$ and $\{B_{1},\dots,B_{J}\}$ of $\mathcal{X}$ such that for any $1\leq i\leq I$ and any $1\leq j\leq J$, $A_{i}\in\mathcal{A}$ and $B_{j}\in\mathcal{B}$ (Bradley, 2005, (7)).
@@
 When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially.
-When there exists $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially fast when $b\neq 1$, and exponentially fast when $b=1$.
+When there exists $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially when $b\neq 1$, and exponentially when $b=1$.
 In Section 6, we will introduce a class of Markov transition operators, whose flavor of mixing (algebraic, sub-exponential, exponential) can be fully controlled.
@@
 The problem of inferring mixing coefficients from a single trajectory of observations has recently received attention from the community.
-In the most recent work (Khaleghi and Lugosi, 2021), consistent estimators for $\beta$- and $\alpha$-mixing coefficients were proposed.
+In the most recent work (Khaleghi and Lugosi, 2023), consistent estimators for $\beta$- and $\alpha$-mixing coefficients were proposed.
 In Section 3, we will provide more explicit and finer finite-sample bounds for the estimation problem in the specialized Markovian setting.
@@
 Under uniform ergodicity, the mixing time is finite, and the $\beta$-mixing coefficients decays exponentially.
-In this setting, we show how to obtain a sample complexity upper bound for high-probability estimates on the invividual $\beta$-mixing coefficients.
+In this setting, we show how to obtain a sample complexity upper bound for high-probability estimates on the individual $\beta$-mixing coefficients.
 Theorem 3.2.
@@
 We move on to the estimation of the average-mixing time.
-Since the average distance to stationarity $d^{\sharp}=\beta$ does not generally enjoy sub-multiplicativity (it does not hold that $t^{\sharp}_{\mathsf{mix}}(\xi)(1/4)$ can be bounded by $t^{\sharp}_{\mathsf{mix}}$ and a polynomial of $\log\frac{1}{\xi}$), we formalize desired guarantees somewhat differently from Hsu et al.
+Since the average distance to stationarity $d^{\sharp}=\beta$ does not generally enjoy sub-multiplicativity (it does not hold that $t^{\sharp}_{\mathsf{mix}}(\xi)$ can be bounded by $t^{\sharp}_{\mathsf{mix}}(1/4)$ and a polynomial of $\log\frac{1}{\xi}$), we formalize desired guarantees somewhat differently from Hsu et al.
 (2019); Wolfer and Kontorovich (2024).
@@
 We summarize the relations between these conditions in Figure 1.
-{tikzcd}
+Bound on ${\text{Bound on }\mathcal{J}_{p}^{(s)}\text{ uniform over }s}$ Condition 1Condition 2Condition 3Bound on ${\text{Bound on }\mathcal{J}_{p}^{(s)}\text{ sub-linear in }s}$Corollary 4.1Corollary 4.1Lemma 4.2Lemma 4.3Lemma 4.1
 Figure 1: Relation between the various ergodic conditions in Subsection 4.1. In this diagram, $\Rightarrow$ is to be read as ”implies under the assumptions of the result referred to”.
@@
 $\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s},$
-where for any $v\in\mathbb{R}^{\mathcal{X}}$, we wrote
-$\left\|\mu\right\|_{V}\doteq\sup_{x\in\mathcal{X}}V(x)v(x).$
+where for any $\nu\in\mathbb{R}^{\mathcal{X}}$, we wrote
+$\left\|\nu\right\|_{V}\doteq\sup_{x\in\mathcal{X}}V(x)\nu(x).$
 Condition 2 (Existence of a dominating distribution).
@@
 is the set of directed edges.
-The growth of a graph (Trofimov, 1985; Campbell et al., 2022) is defined as a function $g\colon\mathbb{N}\to\mathbb{N}\cup\left\{\infty\right\}$, where for $s\in\mathbb{N}$,
+The growth of a graph (Trofimov, 1985; Campbell et al., 2023) is defined as a function $g\colon\mathbb{N}\to\mathbb{N}\cup\left\{\infty\right\}$, where for $s\in\mathbb{N}$,
 $g(s)\doteq\sup_{\mathcal{S}\subset\mathcal{X},\rho\left((\mathcal{X},\mathcal{E})|_{\mathcal{S}}\right)\leq s}\left|\mathcal{S}\right|,$
-where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the of a directed graph, which is the maximum eccentricity of any vertex in the graph [^6] With the convention that a disconnected graph has infinite diameter., $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.
+where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the diameter of a directed graph, which is the maximum eccentricity of any vertex in the graph [^6] With the convention that a disconnected graph has infinite diameter., $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.
 Condition 4 (Connection graph with polynomial growth).
@@
 $\pi_{p,q}^{\intercal}=\frac{1}{p+q}\begin{pmatrix}q\\ p\end{pmatrix}.$
-It follows from a direct computation (see e.g. the proof of Wolfer (2023+, Lemma 8.1)) that
+It follows from a direct computation (see e.g. the proof of Wolfer (2024, Lemma 8.1)) that
 $\begin{split}P_{p,q}^{t}&=1^{\intercal}\pi_{p,q}+\frac{(1-p-q)^{t}}{p+q}\begin{pmatrix}p&-p\\ -q&q\end{pmatrix},\\ \end{split}$
@@
 In other words, we decompose the observed trajectory into $B$ even and $B$ odd blocks of size $s$, as illustrated in Figure 2.
-{tikzpicture}
+[…]
 Figure 2: Decomposing the process into blocks.
@@
 $\{E_{t}\}_{t\in\mathbb{N}}\doteq\left\{\left(X_{t},X_{t+1}\right)\right\}_{t\in\mathbb{N}}$
-forms a Markov chain over the state space =$\mathcal{E}=\left\{x,x^{\prime}\in\mathcal{X}\colon P(x,x^{\prime})>0\right\}$, called the Hudson expansion (Kemeny and Snell, 1983; Wolfer and Watanabe, 2022) of the original chain, with transition operator
+forms a Markov chain over the state space =$\mathcal{E}=\left\{x,x^{\prime}\in\mathcal{X}\colon P(x,x^{\prime})>0\right\}$, called the Hudson expansion (Kemeny and Snell, 1983; Wolfer and Watanabe, 2024) of the original chain, with transition operator
 $\overline{P}(e=(x_{1},x_{2}),e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime}))=\delta\left[x_{2}=x_{1}^{\prime}\right]P(x_{2},x_{2}^{\prime}),$
@@
 $\beta(t)\leq\sum_{x=1}^{K}2\pi_{T}(x)\left[(1-\nu_{K})^{t}+3\sum_{x>K}\pi_{T}(x)+qt\sum_{x>K}\mu_{x}\right]+\sum_{x>K}\pi_{T}(x)$
-which ends the proof.
+which ends the proof. ∎
 Proof of Lemma 6.3.
@@
 Funding
-GW is supported by the Special Postdoctoral Researcher Program (SPDR) of RIKEN.
-The visit of Pierre Alquier to RIKEN in October 2023, during which part of this work was conducted was sponsored by the SPDR program.
+The work of GW was supported in part by the SPDR Program of RIKEN and by Waseda University Grants for Special Research Projects (Project number: 2024C-667).
+The visit of PA to RIKEN in October 2023, during which part of this work was conducted, was sponsored by the SPDR program of RIKEN.

### v4

@@
 We further illustrate its applicability across a range of settings, from two-point to countable spaces, and discuss some practical implications.
-Keywords— Ergodic Markov chain; average mixing time; countable state space; $\beta$-mixing.
+Keywords— Ergodic Markov chain; average-mixing time; countable state space; $\beta$-mixing.
 1  Introduction
@@
 Two observations on the chain separated in time by about $t_{\mathsf{mix}}$ will be approximately independent and sampled from $\pi$.
-Estimates on the mixing time yield exponential concentration inequalities (Paulin, 2015), and in turn, generalization bounds for stable machine learning algorithms trained with Markovian data (Mohri and Rostamizadeh, 2010; Agarwal and Duchi, 2012; Oliveira et al., 2022), regret bounds for reinforcement learning algorithms (Ortner, 2020), and convergence diagnostics for Markov Chain Monte Carlo methods, which are widely used in mathematical physics (Metropolis et al., 1953; Hastings, 1970), computer science (Sinclair, 2012) and statistics (Chib and Greenberg, 1995).
+Estimates on the mixing time yield exponential concentration inequalities (Chung et al., 2012; Paulin, 2015), and in turn, generalization bounds for stable machine learning algorithms trained with Markovian data (Mohri and Rostamizadeh, 2010; Agarwal and Duchi, 2012; Oliveira et al., 2022), regret bounds for reinforcement learning algorithms (Ortner, 2020), and convergence diagnostics for Markov Chain Monte Carlo methods, which are widely used in mathematical physics (Metropolis et al., 1953; Hastings, 1970), computer science (Sinclair, 2012) and statistics (Chib and Greenberg, 1995).
 However, $t_{\mathsf{mix}}$ is an intrinsically pessimistic notion, defined in terms of the worst-case initial distribution.
@@
 for estimation down to constant multiplicative error (Hsu et al., 2019, Theorem 3.1, Theorem 3.2), where $\left|\mathcal{X}\right|$ is the size of the space and $\pi_{\star}$ is the minimum stationary probability (refer to Section 1.3).
-The sample complexity lower bound in (1) is prohibitively large for vast state spaces, and even vacuous in the infinite case.
-In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in (Díaz et al., 2024; Münch and Salez, 2023).
-Specifically, we will confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space.
-Furthermore, we initiate the program of estimating $t_{\mathsf{mix}}^{\sharp}$ from a single trajectory and we demonstrate that it is statistically less costly than the problem pertaining to $t_{\mathsf{mix}}$.
+The sample complexity lower bound in (1) is prohibitively large when $\pi_{\star}\propto\exp(-\left|\mathcal{X}\right|)$ for vast state spaces, and even vacuous in the infinite case.
+In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in Díaz et al.
+(2024); Münch and Salez (2023).
+Specifically, we will first confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space.
+We then initiate the program of estimating $t_{\mathsf{mix}}^{\sharp}$ from a single trajectory of observations and we demonstrate that the problem is statistically less costly than the one pertaining to $t_{\mathsf{mix}}$.
 In fact, contrary to the worst-case problem, our analysis carries to a large subset of countable-state Markov chains.
-One should rightfully expect $t_{\mathsf{mix}}^{\sharp}$ to be less informative than $t_{\mathsf{mix}}$ since it quantifies a weaker notion of mixing.
-However, we confirm that knowledge of $t_{\mathsf{mix}}^{\sharp}$ remains of practical value in the sense where it can be used in known powerful decoupling techniques (Yu, 1994) and appear as a natural parameter controlling the large deviation rate for additive functionals.
+One should rightfully expect $t_{\mathsf{mix}}^{\sharp}$ to be less informative than $t_{\mathsf{mix}}$ since it measures a weaker notion of mixing.
+However, we confirm that knowledge of $t_{\mathsf{mix}}^{\sharp}$ remains of practical value in the sense where it can be used in known powerful decoupling techniques (Yu, 1994) and appears as a natural parameter controlling the large deviation rate for additive functionals.
 Along the way, we explore the connections between the average-mixing time and $\beta$-mixing, argued by Vidyasagar (2013); Mohri and Rostamizadeh (2010) to be “just the right” assumption for the analysis of machine learning algorithms with weakly-dependent data.
 1.1  Related work
+Concentration inequalities for Markov chains typically involve a parameter, such as a spectral gap or the mixing time, which accounts for dependencies within the sample (Lezaud, 2001; A León and Perron, 2004; Dedecker and Gouëzel, 2015; Havet et al., 2020).
+Estimates on this parameter yield corresponding deviation bounds.
+However, this parameter is often unknown and its definition is restricted to geometric ergodic settings.
+1.1.1 Estimation of mixing time in Markov chains
+The program of estimating the (worst-case) mixing time of a Markov chain from a single trajectory of observations was initiated by Hsu et al.
+(2015).
+In the time-reversible setting, they obtained the first upper and lower bounds on the sample complexity of estimating the absolute spectral gap of a chain in multiplicative error, and constructed fully empirical confidence intervals for the quantity.
+The analysis in the reversible setting was later complemented in Hsu et al.
+(2019), introducing an amplification method to obtain a generally unimprovable upper bound on the problem.
+In the more challenging non-reversible setting, the absolute spectral gap is no longer a good proxy for the mixing time.
+In Wolfer and Kontorovich (2024), the authors propose to estimate Paulin’s pseudo-spectral gap (Paulin, 2015) instead, and devise a method which enjoys the same sample complexity as in the reversible setting, modulo logarithmic factors.
+In parallel, Wolfer (2024) put forward a contraction approach, using a generalization of Dobrushin’s coefficient, instead of spectral methods to construct fully empirical confidence intervals for the mixing time.
+1.1.2 Average-mixing time
 Average variants of the mixing time have recently made their apparition in the literature.
@@
 In (Lovász and Kannan, 1999), the authors introduce the notion of a conductance averaged over subsets of states with various sizes, and demonstrate that it can lead to tighter convergence bounds for counting and volume computation problems.
-Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered in (Münch and Salez, 2023) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
-In particular, we stress that the notion of average-mixing treated in this paper is different from the one discussed in Peres and Sousi (2015); Anderson et al.
-(2021), defined by taking an average over two successive steps on the chain.
+Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered by Münch and Salez (2023) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
+In particular, we emphasize that the notion of average-mixing discussed in this paper differs from that in Peres and Sousi (2015); Hermon and Peres (2017, 2018); Anderson et al.
+(2021).
+In these references, averaging is performed by taking an average over two successive steps of the chain to address periodicity issues.
+In contrast, our analysis involves averaging over the starting state, similar to the definition provided by Münch and Salez (2023).
 1.2  Main contributions
-We direct the reader to Section 1.4 for the notation.
+We only introduce the setup and minimal notation in this section.
+We direct the reader to Section 1.3 for the comprehensive set of definitions used in this paper.
+Let $X_{1},\dots,X_{n}$ be a time-homogeneous time-discrete $\pi$-stationary ergodic Markov chain with transition operator $P$ over a countable state space $\mathcal{X}$.
+Assume that we are interested in approximating the stationary mean of a real function $f$ by the sample mean computed over the trajectory $X_{1},\dots,X_{n}$.
+More specifically, we seek a bound on the deviation $\left|f-\mathbb{E}_{\pi}f\right|$.
+To achieve a bound in the Markovian setting similar to those for iid data, the required sample size is typically multiplied by the worst-case mixing time (Chung et al., 2012; Paulin, 2015).
+The underlying idea is that one must allow the chain sufficient time to reach its steady state before regarding the observation as approximately independent and drawn from the stationary distribution.
+However, after the chain has reached steady state for the first time, waiting for an additional mixing time addresses the fact that we may have ended up in the worst possible state.
+While this approach is conservative, it may be unnecessarily cautious.
 1.2.1 Highlight the significance of the average-mixing time
-Observing that the average-mixing time $t^{\sharp}_{\mathsf{mix}}$ recently analyzed by Münch and Salez (2023) corresponds to the stationary $\beta$-mixing time of the Markov chain,
-$t^{\sharp}_{\mathsf{mix}}(\xi)=\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\beta(s)\leq\xi\right\},$
-we briefly demonstrate how the average-mixing time, together with an a priori known order of convergence, can be used to quantitatively capture the convergence rate of a Markov chain in lieu of the worst-case mixing time.
-Specifically, we show (Lemma 2.3) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ naturally appears in the sample complexity for controlling the deviation probability from zero of a centered function $f\colon\mathcal{X}\to[-1,1]$.
-Namely, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$, it is sufficient to have a trajectory of
-$n\geq\frac{\left(e/b\right)^{1/b}}{b\varepsilon^{2}}t^{\sharp}_{\mathsf{mix}}\left(\delta/e\right)\operatorname{polylog}\left(\frac{1}{\delta},\frac{1}{\varepsilon}\right),$
+Once the chain has reached steady state, since the state we reach should be distributed according to the stationary distribution, it might be preferable to consider the next time to stationarity in a more probabilistic way.
+With this in mind, the question arises whether we can replace the worst-case mixing time with the following more intuitive average version,
+$t^{\sharp}_{\mathsf{mix}}(\xi)=\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}}\leq\xi\right\},$
+recently analyzed by Münch and Salez (2023).
+Under this criterion, convergence is achieved faster.
+In fact convergence can occur arbitrarily faster, for denumerable spaces—in which the worst-case mixing may not even be defined—and even for binary state spaces.
+We illustrate how together with an a priori known order of convergence, $t^{\sharp}_{\mathsf{mix}}$ can be used to summarize the convergence rate of a Markov chain in lieu of its worst-case mixing time.
+Specifically, we show (refer to Lemma 2.1) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ can be naturally plugged into the sample complexity for controlling the deviation probability from zero of a $1$-bounded centered function $f$.
+Indeed, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$ (Definition 2.1), it is sufficient to have a trajectory of observations of length
+$n\geq\frac{\left(e/b\right)^{1/b}}{b\varepsilon^{2}}t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\operatorname{polylog}\left(\frac{1}{\delta},\frac{1}{\varepsilon}\right),$
 where $\operatorname{polylog}$ is a poly-logarithmic function of its parameters and $e$ is the Euler number.
-Similarly, in the polynomial $\beta$-mixing setting of order $b$, we show that for some natural constant $C\in\mathbb{R}_{+}$,
+It is instructive to note that $b=1$ corresponds to the assumption that the chain is uniformly ergodic, and that the bound continues to hold for $b<1$, in which case the rate $b$ affects only the constant in the above expression.
+Similarly, in the polynomial $\beta$-mixing setting of order $b$ (Definition 2.1), we show that for some natural constant $C\in\mathbb{R}_{+}$,
 $n\geq C\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(2t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\right)^{\frac{1}{b}},$
-is sufficient.
-1.2.2 Estimation of $\beta$-mixing coefficients
-For a general time-discrete, time-homogeneous Markov chain over a countable state space, we establish a parametric rate of convergence for its individual $\beta$-coefficients from a single stationary trajectory of observations $X_{1},\dots,X_{n}$.
-Namely, for $s\in\mathbb{N}$, we construct an estimator $\widehat{\beta}(s)$ for $\beta(s)$, and we show (refer to Theorem 3.1) that the absolute mean deviation can be controlled as follows,
-$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
-with
-$\mathcal{B}_{p}^{(s)}\leq\left\|\beta^{(s)}\right\|_{1/p}^{1/p},\qquad\mathcal{J}_{p}^{(s)}=\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
+is sufficient. The trajectory lengths listed above both assume a stationary start. However, this assumption can be relaxed (refer to Section A).
+1.2.2 Estimation of the average-mixing time
+In the case where $\mathcal{X}$ is a finite space, for $\xi\in(0,1)$ and $\delta\in(0,1)$, our chief technical contribution is to show (refer to Corollary 4.1) that a trajectory $X_{1},X_{2},\dots,X_{n}$ of observations of length
+$n\geq\frac{t_{\mathsf{mix}}}{\xi^{2}}\left|\mathcal{X}\right|^{2}\operatorname{polylog}\left(t_{\mathsf{mix}},\frac{1}{\xi},\frac{1}{\delta}\right),$
+where $t_{\mathsf{mix}}$ is the mixing time of $P$, is sufficient in order to obtain an estimate $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ of the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\xi)$ which satisfies
+$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(3\xi/2),t_{\mathsf{mix}}^{\sharp}(\xi/2)\right],\end{split}$
+with probability at least $1-\delta$.
+In other words, one can demonstrably estimate with high probability the average-mixing time linearly in the number of parameters of the transition operator, answering a question posed by Luca Zanetti and John Sylvester (Zanetti and John, 2023).
+This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024) (refer to Section 4.3).
+In fact, contrary to the worst-case mixing time estimation problem, we show that it is possible to extend the problem and upper bounds on the sufficient trajectory length to denumerable spaces under additional suitable conditions.
+Namely, when $\mathcal{X}\cong\mathbb{N}$, we show (refer to Theorem 3.5.2) that in the uniformly ergodic setting, a single path of length
+$n\geq\frac{t_{\mathsf{mix}}}{\xi^{2}}\mathcal{J}_{\infty,\xi/2}\operatorname{polylog}\left(t_{\mathsf{mix}},\frac{1}{\xi},\frac{1}{\delta}\right),$
+is sufficient, where for any $p\in[1,\infty]$, $\mathcal{J}_{p,\xi}$ is an entropic term defined as
+$\mathcal{J}_{p,\xi}\doteq\sup\left\{\left\|Q^{(s)}\right\|_{(1-1/p)/2}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$
 and where for $x,x^{\prime}\in\mathcal{X}$, $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$.
-Here, $\mathcal{B}_{p}^{(s)}$ is a mixing parameter roughly controlled from above by the $\ell_{1/p}$ quasi-norm of the sequence of $\beta$-mixing coefficients of the $s$-skipped Markov chain, and $\mathcal{J}_{p}^{(s)}$ is an entropic quantity pertaining to the $s$-skipped Markov chain.
-Crucially, the parameter $p\in[1,\infty)$ is free and will allow us to balance the two factors in the upper bound.
-1.2.3 Estimation of the average-mixing time
-In the uniformly ergodic setting, for $\xi\in(0,1)$ and $\delta\in(0,1)$, we show (Theorem 3.4) that a single path length of
-$n\geq\frac{t_{\mathsf{mix}}}{\xi^{2}}\inf_{p\geq 1}\left\{p\mathcal{J}_{p,3\xi/4}\right\}\operatorname{polylog}\left(\frac{1}{\xi},\frac{1}{\delta},t_{\mathsf{mix}}\right),$
-where $t_{\mathsf{mix}}$ is the mixing time of $P$ and
-$\mathcal{J}_{p,\xi}\doteq\sup\left\{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$
-is sufficient in order to obtain an estimate $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ of the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\xi)$ which satisfies
-$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(3\xi/2),t_{\mathsf{mix}}^{\sharp}(\xi/2)\right]\end{split}$
-with probability at least $1-\delta$. Furthermore, for finite spaces, setting $\xi=1/4$ for simplicity, the parameter $p$ can be chosen so as to obtain an upper bound (refer to Corollary 4.2) of
-$n\geq Ct_{\mathsf{mix}}\log\max\left\{\left|\mathcal{X}\right|,\frac{t_{\mathsf{mix}}}{\delta}\right\},$
-where $C\in\mathbb{R}_{+}$ is a universal constant.
-One can thus demonstrably estimate with high probability the average-mixing time sub-linearly in the number of states, answering a question posed by Luca Zanetti and John Sylvester.
-This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024).
-1.3   Outline
-Section 2 introduces the average-mixing time, illustrates its significance and analyzes some of its properties.
-In Section 2.2, under geometric ergodicity and time-reversibility assumptions, we establish an upper bound of $t^{\sharp}_{\mathsf{mix}}$ in terms of the relaxation time of the chain, which remains non-vacuous over countable state spaces and under suitable conditions on the stationary distribution.
-In Section 2.3 we recall some important connection between the average-mixing time and $\beta$-mixing properties of the chain.
-Section 3 is dedicated to the problem of inferring the $\beta$-mixing coefficients and average-mixing time from a single trajectory of observations.
-In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain; first individually and then as a sequence, with respect to the uniform norm.
-From Section 3.3, we specialize our analysis to the uniformly ergodic setting, and in Section 3.4 we deduce an upper bound for the problem of estimating the average-mixing time from a single trajectory.
-In Section 4, we demonstrate that our contribution holds implications at various scales of state spaces and provide more explicit rates under additional structural assumptions on the underlying chain.
-Technical proofs have been deferred to Section 5 for readability.
-Finally, Section 6 introduces a class of Markov chain which illustrates some of our results.
+We show that one can even analyze and obtain a bound on the sufficient trajectory length in the non uniformly ergodic setting (refer to Theorem 3.5.1).
 1.3  Notation and setting
@@
 and for any $p\in[0,\infty)$,
-$\ell_{p}\doteq\left\{s\in\mathbb{R}^{\mathcal{X}}\colon\left\|s\right\|_{p}<\infty\right\}.$
-The set of all probability distributions over $\mathcal{X}$, written $\mathcal{P}(\mathcal{X})$, corresponds to non-negative functions summing to unity. A time-discrete, time-homogeneous Markov chain
-$X_{1},X_{2},X_{3}\dots$
-on $\mathcal{X}$ is defined by a transition operator represented by an infinite stochastic matrix $P$
-$P(x,x^{\prime})=\mathbb{P}\left(X_{t+1}=x^{\prime}|X_{t}=x\right).$
-We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$.
+$\ell_{p}\doteq\left\{v\in\mathbb{R}^{\mathcal{X}}\colon\left\|v\right\|_{p}<\infty\right\}.$
+The set of all probability distributions over $\mathcal{X}$, written $\mathcal{P}(\mathcal{X})$, corresponds to non-negative functions summing to unity.
+The dynamics of a time-discrete, time-homogeneous Markov chain
+$X_{1},X_{2},X_{3},\dots$
+on $\mathcal{X}$ are governed by a transition operator represented by an infinite stochastic matrix $P$
+$P(x,x^{\prime})=\mathbb{P}\left(X_{t+1}=x^{\prime}|X_{t}=x\right),\forall x,x^{\prime}\in\mathcal{X}.$
+We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$—and we also identify $\mathcal{W}(\mathcal{X})$ with Markov chains whose transition operator belong to $\mathcal{W}(\mathcal{X})$.
 For a finite trajectory of length $n\in\mathbb{N}$, started with an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
 $\mathbb{P}_{\mu}\left(X_{1}=x_{1},\dots,X_{n}=x_{n}\right)=\mu(x_{1})\prod_{t=1}^{n-1}P(x_{t},x_{t+1}).$
-By abuse of notation, matrix multiplication will retain its meaning in the countable setting. Namely, for $\mu\in\mathcal{P}(\mathcal{X})$ and $P\in\mathcal{W}(\mathcal{X})$, we will write,
+Matrix multiplication retains its meaning in the countable setting. Namely, for $\mu\in\mathcal{P}(\mathcal{X})$ and $P\in\mathcal{W}(\mathcal{X})$, we write,
 $\mu P=\sum_{x,x^{\prime}\in\mathcal{X}}\mu(x)P(x,x^{\prime})e_{x^{\prime}}.$
-Extensions of notation to $Pf^{\intercal}$ with $f\in\ell_{1}$, or for $P^{t}$ are similarly defined.
+Extensions of notation to $Pf^{\intercal}$ with $f\in\ell_{1}$, or for $P^{t}$, are similarly defined.
 We say that $P$ is ergodic when $P$ is irreducible, aperiodic and positive-recurrent (Levin et al., 2009).
-In that case, there exists a unique stationary distribution $\pi$ verifying $\pi P=\pi$, and we can define the distance to stationarity
+In that case, there exists a unique stationary distribution $\pi$ verifying $\pi P=\pi$, and we can define the distance to stationarity as
 $d(t)\doteq\sup_{x\in\mathcal{X}}\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}},$
@@
 $d(t)\leq M\rho^{t}.$
-We write $\mathcal{W}^{\star}(\mathcal{X})$ the set of uniformly ergodic Markov chains. We recall that the (worst-case) mixing time is defined as
-$\begin{split}t_{\mathsf{mix}}(\xi)\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d(t)\leq\xi\right\}.\end{split}$
-It is furthermore customary to define
+We write $\mathcal{W}^{\star}(\mathcal{X})$ the set of uniformly ergodic Markov chains—identified with their transition operators.
+We recall that when $P\in\mathcal{W}^{\star}(\mathcal{X})$ the (worst-case) mixing time, defined for any $\xi\in(0,1)$ as
+$\begin{split}t_{\mathsf{mix}}(\xi)\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d(t)\leq\xi\right\},\end{split}$
+is finite. It is furthermore customary to define
 $\qquad t_{\mathsf{mix}}\doteq t_{\mathsf{mix}}(1/4),$ (2)
@@
 $t_{\mathsf{mix}}(\xi)\leq\left\lceil\log_{2}\frac{1}{\xi}\right\rceil t_{\mathsf{mix}}.$ (3)
-Elements related directly to the notion of average-mixing time will be introduced in the next section for clarity of the exposition.
+Elements related to the notion of average-mixing time are deferred to the Section 2 for clarity.
+1.4  Outline
+Section 2 introduces the average-mixing time and illustrates its significance.
+Section 3 is dedicated to the problem of inferring the $\beta$-mixing coefficients and average-mixing time from a single trajectory of observations.
+In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain individually, first under general ergodicity condition and then under uniform ergodicity.
+In Section 3.2 we deduce upper bounds for the problem of estimating the average-mixing time from a single trajectory.
+In Section 4, we demonstrate that our contribution holds implications at various scales of state spaces and provide more explicit rates under additional structural assumptions on the underlying chain.
+Technical proofs, extensions to non-stationary chains and a discussion on the connection with $\beta$-mixing have been deferred to the supplementary material for readability.
 2  The average-mixing time
-In this section, we first recall the definition of the average-mixing time, and subsequently explore some of its basic properties.
-For $P\in\mathcal{W}(\mathcal{X})$, the average distance to stationarity of the chain induced from $P$ at time $t\in\mathbb{N}$ is expressed as
-$d^{\sharp}(t)\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}},$
-where $\pi$ is the stationary distribution of $P$.
-The above should be understood as the total variance distance of the chain to stationarity, averaged with respect to $\pi$ over the starting state.
-The limit $\lim_{t\to\infty}d^{\sharp}(t)=0$ always exists [^1] This fact follows from a connection between average-mixing time and $\beta$-mixing discussed at Section 2.3., thus for $\xi\in(0,1)$, the average-mixing time of $P$, defined (Münch and Salez, 2023) as
-$\begin{split}t^{\sharp}_{\mathsf{mix}}(\xi)&\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d^{\sharp}(t)\leq\xi\right\},\end{split}$
-always verifies $t^{\sharp}_{\mathsf{mix}}(\xi)<\infty$.
+In this section, we first recall the definition of the average-mixing time and then explore some of its basic properties.
+For a $\pi$-stationary transition matrix $P\in\mathcal{W}(\mathcal{X})$ and for a proximity parameter $\xi\in(0,1)$, the average-mixing time of $P$ is defined (Münch and Salez, 2023) as
+$\begin{split}t^{\sharp}_{\mathsf{mix}}(\xi)&\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{\beta(t)\leq\xi\right\},\end{split}$
+where $\beta\colon\mathbb{N}\to[0,1]$ is a sequence defined by
+$\beta(t)\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}}.$ (4)
+Observe that instead of considering the distance to stationarity from the worst-case starting state, as the distance $d$ does, $\beta$ takes an average distance over starting states with respect to the long term probability of being in each state.
+This choice naturally reduces the contribution of states that are difficult to reach, resulting in a more optimistic convergence criterion.
 Remark 2.1.
-Similar to $t_{\mathsf{mix}}$ in (2), it is tempting to introduce $t_{\mathsf{mix}}^{\sharp}\doteq t_{\mathsf{mix}}^{\sharp}(1/4)$.
-However, since $d^{\sharp}$ does not enjoy the same sub-multiplicativity properties as $d$ (3) (refer to (Levin et al., 2009, Lemma 4.11)), it is generally not enough to consider a fixed value of $\xi\in(1/2)$, and in this manuscript we will keep $\xi$ to be a user-fixed value.
+Similar to the definition of $t_{\mathsf{mix}}$ in (2), one may wish to introduce $t_{\mathsf{mix}}^{\sharp}\doteq t_{\mathsf{mix}}^{\sharp}(1/4)$.
+However, since $\beta$ does not enjoy the same sub-multiplicativity properties as $d$ —refer to (3) and (Levin et al., 2009, Lemma 4.11)—it is generally not enough to consider an arbitrary value of $\xi\in(1/2)$, and in this manuscript we will keep $\xi$ as a user-fixed value.
+Remark 2.2 (Average-mixing time is stationary $\beta$-mixing time).
+The astute reader will have noticed that we choose $\beta$ to denote the average distance to stationarity instead of $d^{\sharp}$, as in (Münch and Salez, 2023).
+This choice is motivated by our observation that the average-mixing time corresponds to the stationary $\beta$-mixing time of the Markov chain, which we discuss in detail in Section B.
+The sequence $\beta$ defined at (4) will henceforth be referred to as the sequence of (stationary) $\beta$-mixing coefficients.
+An immediate consequence of Remark 2.2 is that the limit $\lim_{t\to\infty}\beta(t)=0$ always exists (Bradley, 2005), or in other words, $t^{\sharp}_{\mathsf{mix}}(\xi)<\infty$ for any $\xi\in(0,1)$ and any ergodic $P\in\mathcal{W}(\mathcal{X})$ over countable $\mathcal{X}$.
+However, the rate of convergence of the sequence of $\beta$-mixing coefficients can be arbitrarily slow.
+Definition 2.1.
+Rates of convergence for $\beta$-mixing.
+  1. 1.
+    (Sub-exponential, exponential) When there exists $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$ and $b\in(0,1]$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially when $b\neq 1$, and exponentially when $b=1$.
+  2. 2.
+    (Polynomial) When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially.
+Notably, when $P$ is uniformly ergodic with mixing time $t_{\mathsf{mix}}$, it holds that
+$\beta(s)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$ (5)
+hence the process is exponentially $\beta$-mixing with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$.
+However, nothing a priori precludes the $\beta$-mixing rate from being much faster than (5).
 2.1  Average-mixing versus worst-case mixing
-Following their respective definitions, it holds that $d^{\sharp}(t)\leq d(t)$.
-Thus, for any fixed $\xi\in(0,1)$, average-mixing occurs always faster than mixing.
-The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2023, Remark 1), which in particular include random walks on groups [^2] A chain is called transitive when for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$..
-In general however, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as formalized below.
+From their respective definitions, it is immediate that $\beta(t)\leq d(t)$.
+In other words, for any fixed $\xi\in(0,1)$, average-mixing never occurs slower than mixing.
+The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2023, Remark 1), which are chains such that for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$ (Levin et al., 2009, Section 2.6.2).
+Transitive chains include, but are not limited to, random walks on groups (McKay and Praeger, 1996).
+For non-transitive chains, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as made precise below.
 Proposition 2.1.
@@
 Lemma 4.2 establishes this fact in the case where $\mathcal{X}=\left\{0,1\right\}$. ∎
-2.2   Spectral methods under reversibility and geometric ergodicity
-In this section, we endow $\mathbb{R}^{\mathcal{X}}$ with the inner product $\langle f,g\rangle_{\pi}\doteq\sum_{x\in\mathcal{X}}f(x)g(x)\pi(x)$, where $\pi$ is a positive distribution.
-We regard $P$ as a linear operator on the resulting Hilbert space $\ell_{2}(\pi)\doteq(\mathbb{R}^{\mathcal{X}},\langle\cdot,\cdot\rangle_{\pi})$ and further assume $P$ to be geometrically ergodic.
+In fact, as we will see in the next section, a regime of particular interest to us is when the chain $\beta$-mixes much faster than $\Theta(t_{\mathsf{mix}})$.
+We begin by recalling the well-studied class of birth and death Markov chains, which will serve as our running example.
+Example 2.1 (Birth and death Markov chains).
+We let $\mathcal{X}=\mathbb{N}$, and for three sequences $u,v,w\in[0,1]^{\mathbb{N}}$ such that $u(1)=0$ and $\forall n\in\mathbb{N},u(n)+v(n)+w(n)=1$, we define the transition operator,
+$P_{u,v,w}=\begin{pmatrix}v(1)&w(1)&0&0&0&\ldots\\ u(2)&v(2)&w(2)&0&0&\vdots\\ 0&u(3)&v(3)&w(3)&0&\ddots\\ 0&0&u(4)&v(4)&\ddots&\ddots\\ \vdots&\ldots&\ddots&\ddots&\dots&\ddots\\ \end{pmatrix}$.
+Due to their ubiquity, ergodicity properties of birth and death chains have been extensively studied in the literature (Miclo, 1999; Ding et al., 2010; Kovchegov, 2010; Chen and Saloff-Coste, 2013).
+In the special case where $v(1)=0$ and for $n\geq 2$ the sequences $u,v,w$ are constant with $u>w$, it can be shown that the family of chains
+$P_{u,v,w}=\begin{pmatrix}0&1&0&0&0&\ldots\\ u&v&w&0&0&\vdots\\ 0&u&v&w&0&\ddots\\ 0&0&u&v&\ddots&\ddots\\ \vdots&\ldots&\ddots&\ddots&\dots&\ddots\\ \end{pmatrix}$
+is geometrically ergodic (Meitz and Saikkonen, 2021; Kovchegov, 2010), that is, there exists $M\colon\mathcal{X}\to\mathbb{R}_{+}$ and $\rho\in(0,1)$ such that for any $t\in\mathbb{N}$ and any $n_{0}\in\mathcal{X}$,
+$\left\|e_{n_{0}}P_{u,v,w}^{t}-\pi\right\|_{\mathsf{TV}}\leq M(n_{0})\rho^{t}.$ (6)
+In particular, the geometric rate of the above family is given by
+$\rho=\max\left\{v+2\sqrt{uw},\frac{w}{w+v}\right\}.$
+While geometric ergodicity is weaker than uniform ergodicity in as much as the distance to the stationary distribution depends on the starting state, convergence remains exponential and leads to a powerful convergence analysis based on spectral methods (refer also to Section 2.3.1).
+But we now turn our attention to the even less favorable case of sub-geometric convergence, where the aforementioned spectral methods break down.
+This more challenging setting can arise even in for the seemingly simple class of birth and death chains in this example.
+Indeed, let us borrow the following instantiation of a Chebyshev–type random walk, proposed by Kovchegov and Michalowski (2013, Section 4).
+Let $\theta>0$ and $\lambda\geq\frac{2\theta^{2}}{(1+\theta)(1+3\theta)}$.
+We define
+[…]
+while for $n\geq 2$,
+[…]
+and v(n)=1u(n)w(n)v(n)=1-u(n)-w(n).
+For convenience, we denote $P_{\theta,\lambda}=P_{u,v,w}$ for the above-defined sequences.
+It can be verified that the stationary distribution of $P_{\theta,\lambda}$ satisfies for any $n\in\mathbb{N}$,
+$\pi_{\theta}(n)=\frac{2(1+\theta)\theta}{(1+(2n-1)\theta)(1+(2n-3)\theta)}\asymp\frac{1}{n^{2}}.$
+One interesting feature is that the chain does not have a spectral gap—or in other words, it “[lies] outside the scope of geometric ergodicity theory” (Kovchegov and Michalowski, 2013).
+In fact, for the random walk originating at state $n_{0}\in\mathbb{N}$, it holds (Kovchegov and Michalowski, 2013, Theorem 2) that for any $t\in\mathbb{N}$,
+$\frac{c}{\sqrt{t}}\leq\left\|e_{n_{0}}P_{\theta,\lambda}^{t}-\pi_{\theta}\right\|_{\mathsf{TV}}\leq\frac{Cn_{0}\log(t+n_{0}+2)}{\sqrt{t}},$
+where $c$ is a constant depending on $\theta$, $\lambda$ and the starting state $n_{0}$, while $C$ is a constant which depends only on $\theta$ and $\lambda$. Let $\xi\in(0,1)$ and denote
+$n_{\star}=n_{\star}(\theta,\xi/2)=\min_{n\in\mathbb{N}}\left\{\sum_{k=n+1}^{\infty}\pi_{\theta}(k)<\xi/2\right\}.$
+We can thus bound the $\beta$-mixing coefficient as follows,
+[…]
+where $C^{\prime}$ is a constant which depends only on $\theta$ and $\lambda$ and for any $k\in\mathbb{N}$, $H_{k}$ denotes the $k$th harmonic number which is bounded from above as $H_{k}\leq 1+\log(k)$.
+Furthermore, one can verify that $n_{\star}\leq 2+\sqrt{(1+\theta)/(\xi\theta)}$.
+As a result, there exists a constant $C^{\prime\prime}$ which depends only on $\theta$ and $\lambda$ such that,
+[…]
+and we obtain that
+$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\frac{1}{\xi^{2}}\operatorname{polylog}(1/\xi),$
+where $\operatorname{polylog}$ is a poly-logarithmic function of its argument, which may also depend on $\theta$ and $\lambda$.
+Our example demonstrates how the contributions to $\beta(s)$ starting from the tail of the stationary distribution tend to be largely absorbed.
+2.2  Deviation of additive functionals of Markov chains
+We now illustrate how the average-mixing time, through its connection to $\beta$-mixing naturally appears as a quantity of interest when bounding deviations of functions evaluated on a trajectory sampled from a Markov chain.
+More formally, let $X_{1},\dots,X_{n}$ be a Markov chain over a countable space $\mathcal{X}$, and whose dynamics are governed by $P\in\mathcal{W}(\mathcal{X})$ with stationary distribution $\pi$.
+For simplicity, we consider a real bounded function $f$ on $\mathcal{X}$, which is centered in the sense where its stationary expected value vanishes, that is $\mathbb{E}_{\pi}f=0$.
+The ergodic theorem states that the empirical mean taken on an infinite trajectory will similarly vanish,
+$\lim_{n\to\infty}\frac{1}{n}\sum_{t=1}^{n}f(X_{t})=0.$
+In practice however, we are more likely to be interested in controlling the deviation of the sample mean from its expected value non-asymptotically.
+The following lemma illustrates how to recover a finite sample bound on the deviation probability, given a bound on the average-mixing time and the rate of convergence of the sequence of $\beta$-mixing coefficients.
+Lemma 2.1 (Bounding deviation of additive functionals evaluated on Markov chains).
+Let $f\colon\mathcal{X}\to[-1,1]$ with $\mathbb{E}_{\pi}f=0$, let $\varepsilon,\delta\in(0,1)$.
+For any $\pi$-stationary ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, and average-mixing time $t^{\sharp}_{\mathsf{mix}}\colon(0,1)\to\mathbb{N}$, the following statements hold.
+  1. 1.
+    (Sub-exponential) If there exist $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ with $\beta_{0}\geq 1$, $b\in(0,1]$, such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$, then there exists a universal constant $C\in\mathbb{R}_{+}$ such that for
+    $n\geq\frac{C}{\varepsilon^{2}}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right)\log\left(\frac{1}{\delta}\right),$
+    with ξ(ε, δ) δε216 log(4δ), it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$. The above upper on the path length can be further streamlined as
+    $n\geq\frac{C}{\varepsilon^{2}}\frac{1}{b}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\delta/e\right)\log\left(\frac{1}{\delta}\right)\log\left(\frac{\log(1/\delta)}{\varepsilon}\right),$
+    for some universal constant $C\in\mathbb{R}_{+}$.
+  2. 2.
+    (Polynomial) If there exist $\beta_{1},b\in\mathbb{R}_{+}$ with $\beta_{1}\geq 1$, such that for any $s\in\mathbb{N}$, $\beta(s)\leq\frac{\beta_{1}}{s^{b}}$, then there exists a universal constant $C\in\mathbb{R}_{+}$ such that for
+    $n\geq C\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(2t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\right)^{\frac{1}{b}},$
+    it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$.
+Proof sketch.
+We adopt a traditional method described in Yu (1994).
+This approach combines a blocking technique, credited to Bernstein (1927) in the context of $m$-dependent processes, together with an application of Hoeffding’s inequality to control the deviation probability of a sum of independent observations.
+We further optimize for the block size and reformulate the results to obtain an expression involving the average-mixing time under various rates of $\beta$-mixing.
+Refer to the supplementary material for the complete proof.
+∎
+As a consequence of Lemma 2.1, in both the sub-exponential and polynomial settings, provided an assumption on the convergence order $b$, estimates on the average-mixing time to proximity $\Theta(\delta)$ yield corresponding deviation bounds.
+In the uniformly ergodic setting, corresponding to the assumption of the exponential setting ($b=1$), we obtain a deviation inequality involving the average-mixing time in lieu of the mixing time (Chung et al., 2012, Theorem 3.1), (Paulin, 2015, Corollary 2.10).
+The lemma generalizes to non-centered functions, lower deviations and arbitrarily bounded functions.
+It is also noteworthy that in the sub-exponential setting, the convergence order $b$ solely affects the constant in the sample complexity.
+Our formulation above serves as motivation for analyzing and inferring the average-mixing time.
+Remark 2.3.
+The astute reader will observe that, in the uniformly ergodic setting, the logarithmic dependence in the confidence parameter $\delta$ is slightly sharper for bounds involving the worst-case mixing time.
+(Chung et al., 2012, Theorem 3.1), (Paulin, 2015, Corollary 2.10).
+This is a consequence of our proof technique relying on $\beta$-mixing.
+2.3  Under geometric ergodicity
+In this section, we briefly examine the average-mixing time under geometric ergodicity, first using spectral methods under time-reversibility, and then through more broadly applicable Lyapunov-type methods.
+2.3.1 Spectral methods under reversibility
+Spectral methods are known to provide a direct and concise framework for analyzing the worst-case mixing time.
+In this section, we briefly examine their applicability to the study of the average-mixing time.
+We endow $\mathbb{R}^{\mathcal{X}}$ with the inner product $\langle f,g\rangle_{\pi}\doteq\sum_{x\in\mathcal{X}}f(x)g(x)\pi(x)$, where $\pi$ is a positive distribution.
+We regard $P$ as a linear operator on the resulting Hilbert space $\ell_{2}(\pi)\doteq(\mathbb{R}^{\mathcal{X}},\langle\cdot,\cdot\rangle_{\pi})$ and further assume $P$ to be geometrically ergodic—refer to (6).
 Recall that $P\in\mathcal{W}^{\star}(\mathcal{X})$ is called $\pi$-reversible when the following detailed-balance equation holds,
@@
 Under reversibility, the spectrum of $P$, denoted $\operatorname{Spec}(P)$, is contained in the real line.
-The absolute spectral gap of $P$ is then defined as
-$\gamma_{\star}\doteq 1-\sup\left\{\left|\lambda\right|\colon\lambda\in\operatorname{Spec}(P),\lambda\neq 1\right\}.$
+The absolute spectral gap of $P$ is defined as
+$\gamma_{\star}\doteq 1-\sup\left\{\left|\lambda\right|\colon\lambda\in\operatorname{Spec}(P),\lambda\neq 1\right\}.$ (7)
 Geometric ergodicity and existence of a positive spectral gap are known to be equivalent (Kontoyiannis and Meyn, 2012).
@@
 $t^{\sharp}_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{2\xi},$
-where $t_{\mathsf{rel}}$ is the relaxation time of $P$.
+where $t_{\mathsf{rel}}$ is the relaxation time of $P$ defined by $t_{\mathsf{rel}}=1/\gamma_{\star}$ with $\gamma_{\star}$ the absolute spectral gap of $P$ defined in (7).
 The half quasi-norm of the stationary distribution $\|\pi\|_{1/2}$, making its first appearance in our study, is a natural measure of complexity of the stationary distribution $\pi$, directly related to its Rényi entropy of order $1/2$.
@@
 $t_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{1}{\xi\pi_{\star}},$
-where $\pi_{\star}\doteq\min_{x\in\mathcal{X}}\pi(x)$, which is only informative [^4] We note in passing that the bound at Levin et al.
-(2009, Theorem 12.4) could be improved as $t_{\mathsf{mix}}(\xi)\leq\frac{1}{2}t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{\xi\pi_{\star}}$.
-when $\left|\mathcal{X}\right|<\infty$.
-2.3   Connection with $\beta$-mixing
+where $\pi_{\star}\doteq\min_{x\in\mathcal{X}}\pi(x)$, which is only informative when $\left|\mathcal{X}\right|<\infty$.
+2.3.2 Lyapunov methods
+Following a classical Banach space structure, the definitions of the $V$-norm of a function $f$ on $\mathcal{X}$ and a signed measure $\mu$ on $\mathcal{X}$ are given (Douc et al., 2018, Definition D.3.1) respectively as ,
+$\left\|f\right\|_{V}\doteq\sup_{x\in\mathcal{X}}\frac{\left|f(x)\right|}{V(x)}\qquad\left\|\mu\right\|_{V}\doteq\sum_{x\in\mathcal{X}}\left|\mu(x)\right|V(x)=\sup_{\left\|f\right\|_{V}\leq 1}\left|\sum_{x\in\mathcal{X}}f(x)\mu(x)\right|.$
+We say that $P$ satisfies $V$-geometrical ergodicity (Hervé and Ledoux, 2020) when there exists and $\rho_{V}\in(0,1)$ such that for all $x\in\mathcal{X}$ and all $t\in\mathbb{N}$,
+$\left\|e_{x}P^{t}-\pi\right\|_{V}\leq V(x)\rho_{V}^{t}.$
+In particular, $P$ is known to be $V$-geometrically ergodic under irreducibility, aperiodicity, drift and minorization conditions (Hervé and Ledoux, 2020).
+In this case, observe that we can bound the stationary $\beta$-coefficient as follows,
+[…]
+As a result,
+$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\cfrac{\log\left\|\pi\right\|_{V}+\log\frac{1}{2\xi}}{\log\frac{1}{\rho_{V}}},$
+which is finite whenever $\left\|\pi\right\|_{V}=\mathbb{E}_{\pi}\left[V\right]$ is finite.
+3  Estimation of average convergence from a single trajectory
+To apply the results in Section 2, one needs to have an upper bound on the average-mixing time.
+However, such bound may not be known a priori, and may need to be estimated from the data.
+In this section, we consider the problem of estimating the average-mixing time $t^{\sharp}_{\mathsf{mix}}$ from a single trajectory of observations.
+Specifically, for a user-fixed proximity parameter $\xi\in(0,1)$ and a confidence parameter $\delta\in(0,1)$, our goal is to construct an interval $I_{\delta}\subset[1,\infty)$, such that given a trajectory $X_{1},\dots,X_{n}$ sampled from an unknown transition matrix $P$, and for $n$ sufficiently large, it holds that $t_{\mathsf{mix}}^{\sharp}(\xi)\in I_{\delta}$ with probability at least $1-\delta$.
+We first analyze the intermediary problem of estimating the $\beta$-mixing coefficients of the process.
+We subsequently show how to convert such results into estimation procedures for the average-mixing time both in the uniformly ergodic and non-uniformly ergodic settings (refer to Figure 1).
+[…]
+Figure 1: Logical flow between the estimation results of Section 3. MAD: Mean Absolute Deviation; PAC: Probably Approximately Correct.
+3.1  Estimation of $\beta$-mixing coefficients
+We begin by analyzing a sequence of plug-in estimators for the $\beta$-mixing coefficients. Our analysis relies on stationary skipped trajectories, where for $s\in\mathbb{N},s<n$, we denote
+$\begin{split}X^{(s)}\doteq X_{1},X_{1+s},X_{1+2s},\dots,X_{1+\lfloor(n-1)/s\rfloor s},\end{split}$
+with transition operator $P^{s}$ and initial distribution $\pi$. It will be convenient to define the following counting random variables,
+$\begin{split}N_{x}^{(s)}&\doteq\sum_{t=1}^{\lfloor(n-1)/s\rfloor}\delta\left[X_{1+s(t-1)}=x\right],\\ N_{xx^{\prime}}^{(s)}&\doteq\sum_{t=1}^{\lfloor(n-1)/s\rfloor}\delta\left[X_{1+s(t-1)}=x,X_{1+st}=x^{\prime}\right].\end{split}$
+We use the shortcut notation $N_{x}=N_{x}^{(1)}$ and $N_{xx^{\prime}}=N_{xx^{\prime}}^{(1)}$ to respectively denote number of visits to state $x$ and transitions from $x$ to $x^{\prime}$ on the original chain.
+For $s<n$, we define the following estimator $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ for $\beta(s)$,
+$\widehat{\beta}(s)\doteq\frac{1}{2\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|,$ (8)
+and by convention, we set $\widehat{\beta}(s)\doteq 0$ for $s\geq n$.
+3.1.1 Under general ergodicity condition
+Under general—possibly non uniform—ergodicity, the mixing time may be infinite, and the $\beta$-mixing coefficients may decay at arbitrary slow rate.
+We first analyze the convergence properties of our estimator $\widehat{\beta}(s)$ under an arbitrary $\beta$-mixing rate, and will subsequently focus on exponential, sub-exponential and polynomial rates of convergence (refer to Definition 2.1).
+Theorem 3.1 (Mean Absolute Deviation).
+Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+}$, with $p\geq 1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (8).
+For any $\pi$-stationary ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, it holds that
+$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq 3\sqrt{\frac{\left(1/2+\mathcal{B}_{p}^{(s)}\right)\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
+with
+$\mathcal{B}_{p}^{(s)}\doteq\sum_{t=0}^{\lfloor(n-1)/s\rfloor-1}\beta(st)^{1/p},\qquad\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|^{1-1/p}_{(1-1/p)/2},$
+where for $s\in\mathbb{N}$, $\beta(s)$ follows Definition 4, for $x,x^{\prime}\in\mathcal{X}$, we wrote $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$, and for $q\in\mathbb{R}_{+}$,
+$\left\|Q^{(s)}\right\|^{q}_{q}\doteq\sum_{(x,x^{\prime})\in\mathcal{X}^{2}}Q^{(s)}(x,x^{\prime})^{q}.$
+Theorem 3.2 ($(\varepsilon,\delta)$-PAC Bound).
+Let $\varepsilon,\delta\in(0,1)$, $s,n\in\mathbb{N}$ with $n>2s+1$, $p\in\mathbb{R}_{+}$, with $p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (8).
+There exists a universal constant $C_{\mathsf{erg.}}\in\mathbb{R}_{+}$ such that for any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, when
+$n\geq 1+C_{\mathsf{erg.}}\frac{\log(1/\delta)}{\varepsilon^{2}}\max\left\{\frac{s}{\varepsilon^{2}}\mathcal{B}_{p}^{(s)}\mathcal{J}^{(s)}_{p},t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right)\right\},\qquad\xi(\varepsilon,\delta)=\frac{\varepsilon^{2}\delta}{C_{\mathsf{erg.}}\log(1/\delta)},$
+with probability at least $1-\delta$, it holds that
+$\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon.$
+In Theorem 3.1 and Theorem 3.2, we observe that the two quantities $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$, discussed below, control the estimation rate.
+In infinitely denumerable state spaces, for any $s\in\mathbb{N}$, the well-definition of the estimation rate therefore hinges on the existence of a $p\in[1,\infty)$ such that both $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$ are finite.
+The $\beta$-mixing parameter $\mathcal{B}_{p}^{(s)}$.
+This parameter summarizes the $\beta$-mixing properties of the process and is controlled from above by the $\ell_{1/p}$ pseudo-norm of the sequence of the $\beta^{(s)}$-mixing coefficients,
+$\mathcal{B}_{p}^{(s)}\leq\left\|\beta^{(s)}\right\|_{1/p}^{1/p}.$
+It will be convenient to introduce the shorthand notation $\mathcal{B}_{p}\doteq\mathcal{B}_{p}^{(1)}$ to refer to the sum of the $\beta$-mixing coefficients.
+It is instructive to observe that since $\beta(s)\leq 1$ for any $s\in\mathbb{N}$, $\mathcal{B}_{p}^{(s)}$ increases together with $p$, and
+$\lim_{p\to\infty}\mathcal{B}_{p}^{(s)}\asymp\frac{n}{s}.$
+Additionally,
+$r\leq s\implies\mathcal{B}_{p}^{(r)}\geq\mathcal{B}_{p}^{(s)}.$
+In the following Lemma 3.1, we readily provide more explicit upper bounds on $\mathcal{B}_{p}^{(s)}$ under additional assumptions on the mixing rate.
+Lemma 3.1.
+Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$. For any $\pi$-stationary ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$, the following statements hold.
+  1. 1.
+    (Exponential) If $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s)$, with $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$ and $t\in\mathbb{N}\cup\left\{0\right\}$,
+[…]
+  2. 2.
+    (Sub-exponential) If $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$, with $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$, $b\in(0,1)$, and $s\in\mathbb{N}\cup\left\{0\right\}$.
+    $\begin{split}\mathcal{B}_{p}^{(s)}&\leq 1+\exp(-\beta_{1}s^{b}/p)+\frac{\Gamma\left(\frac{1}{b},s^{b}\beta_{1}/p\right)}{bs(\beta_{1}/p)^{1/b}},\end{split}$
+    where $\Gamma$ is the (upper) incomplete Euler gamma function, given by $\Gamma(a,x)=\int_{x}^{\infty}t^{a-1}e^{-t}dt$.
+  3. 3.
+    (Polynomial) If $\beta(s)\leq\beta_{1}/s^{b}$, with $\beta_{1}\in\mathbb{R}_{+}$, $b\in(1,\infty)$, $b/p\in(1,\infty)$, $s\in\mathbb{N}$ and $\beta(0)=\beta_{0}\in\mathbb{R}_{+}$,
+    $\begin{split}\mathcal{B}_{p}^{(s)}&\leq\frac{\beta_{0}^{1/p}+\zeta(b/p)\frac{\beta_{1}^{1/p}}{s^{b/p}}}{\lfloor(n-1)/s\rfloor},\end{split}$
+    where $\zeta$ is the Riemann zeta function defined by $\zeta(r)=\sum_{t=1}^{\infty}t^{-r}$.
+The entropic parameter $\mathcal{J}_{p}^{(s)}$.
+For $q<1$, since the $\ell_{q}$ quasi-norm and the $q$th order Rényi entropy $H_{q}$ of a distribution $\mu\in\mathcal{P}(\mathcal{X})$ are related by the expression
+$\frac{1}{1-q}\log\left\|\mu\right\|_{q}^{q}=H_{q}(\mu),$
+the quantity $\mathcal{J}_{p}^{(s)}$ pertains to an entropic property inherent in the Markov process.
+Observe that in contradistinction with $\mathcal{B}_{p}^{(s)}$, the quantity $\mathcal{J}_{p}^{(s)}$ decreases with $p$, leading to a trade-off in the choice of the parameter $p$.
+A necessary condition on $P$ for a rate to exist is readily given by
+$\mathcal{J}_{\infty}^{(s)}\doteq\lim_{p\to\infty}\mathcal{J}_{p}^{(s)}=\left\|Q^{(s)}\right\|_{1/2}<\infty,$ (9)
+that is, the tail of the distribution of stationary pairs of observations separated by a time $s$ must not decay more slowly than an inverse square polynomial.
+A more in-depth analysis of $\mathcal{J}_{p}^{(s)}$ is deferred to Section 4, where we show how to obtain upper bounds under natural assumptions.
+3.1.2 Under uniform ergodicity
+Under uniform ergodicity, the mixing time is finite and the $\beta$-mixing coefficients decay exponentially.
+In this setting, we show how to obtain a stronger sample complexity upper bound for high-probability estimates on the individual $\beta$-mixing coefficients.
+In particular, uniform ergodicity removes the need to consider the trade-off between mixing and entropic parameters mentioned in the previous section.
+Theorem 3.3 (Mean Absolute Deviation and $(\varepsilon,\delta)$-PAC Bound).
+Let $s,n\in\mathbb{N}$ with $n>2s+1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (8).
+There exists a universal constant $C_{\mathsf{u.erg.}}\leq 3\sqrt{2}$ such that for any $\pi$-stationary uniformly ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$,
+[…]
+where $t_{\mathsf{mix}}$ is the mixing time of $P$ and $\mathcal{J}_{\infty}^{(s)}$ is defined in (9).
+Theorem 3.4 ($(\varepsilon,\delta)$-PAC Bound).
+Let $\varepsilon,\delta\in(0,1),s,n\in\mathbb{N}$ with $n>2s+1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (8).
+There exists a universal constant $C_{\mathsf{u.erg.}}\leq 3\sqrt{2}$ such that for any $\pi$-stationary uniformly ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$, for
+$n\geq 1+\frac{4(t_{\mathsf{mix}}+2s)}{\varepsilon^{2}}\max\left\{C_{\mathsf{u.erg.}}^{2}\mathcal{J}_{\infty}^{(s)},576\log\frac{2}{\delta}\right\},$
+with probability at least $1-\delta$, it holds that
+$\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon.$
+Remark 3.1.
+Specializing Theorem 3.1 to obtain a PAC bound in the uniformly ergodic setting would instead yield an inferior PAC upper bound in
+$n\geq\frac{C}{\varepsilon^{2}}\max\left\{\left(s+pt_{\mathsf{mix}}\right)\mathcal{J}_{p}^{(s)},(s+t_{\mathsf{mix}})\log\frac{1}{\delta}\right\},C>0.$
+3.1.3 Comparison with prior results
+In McDonald et al. (2015, Theorem 4), the authors establish the nearly parametric convergence rate for estimating $\beta$-mixing coefficients of continuously valued Markov processes
+$\mathbb{E}{\left|\widehat{\beta}(s)-\beta(s)\right|}\leq\mathcal{O}\left(\sqrt{\frac{W_{0}(n)}{n}}\right),$
+where $W_{0}$ is the Lambert function (Corless et al., 1996).
+They also posit that their findings may be considered conservative when applied to discrete-valued processes.
+Theorem 3.1 validates their remark by recovering the superior parametric rate $\mathcal{O}(\sqrt{1/n})$ under convergence of $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$.
+In the general ergodic setting, our method shares similarities with Khaleghi and Lugosi (2023), as both employ Rio’s covariance inequality Rio (1999, Corollary 1.1).
+However, compared to both Khaleghi and Lugosi (2023) and McDonald et al.
+(2015, Theorem 4), we distinguish ourselves by focusing on the discrete space setting and deriving explicit rates in terms of fine-grained properties of the transition operator.
+We also note that Grünewälder and Khaleghi (2026) address the estimation of $\beta$-mixing coefficients in the geometrically ergodic setting, while we consider a broader class of $\beta$-mixing processes, including sub-exponentially and algebraically $\beta$-mixing ones.
+Finally, under uniform ergodicity, instead of Rio’s covariance inequality, we rely on Paulin (2015) which yields a stronger bound on the MAD and sharper PAC bounds (refer to Remark 3.1).
+3.2  Estimation of the average-mixing time
+We now show how the sequence of estimators for the $\beta$-mixing coefficients, discussed in the previous section, translates into an estimator for the average-mixing time.
+Recall that the average distance to stationarity $\beta$ does not generally enjoy sub-multiplicativity.
+In turn it $t^{\sharp}_{\mathsf{mix}}(\xi)$ cannot generally be bounded by $t^{\sharp}_{\mathsf{mix}}(1/4)$ and a polynomial of $\log(1/\xi)$.
+Consequently, we formalize desired guarantees on our estimator differently from Hsu et al.
+(2019); Wolfer and Kontorovich (2024) for the relaxation time, where the error is defined multiplicatively with respect to the estimator $|\widehat{t_{\mathsf{rel}}}/t_{\mathsf{rel}}-1|\leq\varepsilon$.
+When estimating the average-mixing time, we treat the proximity $\xi$ to stationarity as a free parameter, and construct an estimator that targets a band $\xi(1\pm\varepsilon)$.
+Specifically, we define for $\xi\in(0,1)$,
+$\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\doteq\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\widehat{\beta}(s)\leq\xi\right\},$ (10)
+where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (8), and our next objective is to obtain with high probability an upper bound on the trajectory length such that
+$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
+As illustrated on Figure 2, the interval above becomes thinner as $\varepsilon$ decays to zero.
+Our next lemma explains how to convert a sufficient trajectory length for high-probability control of the error of the sequence of estimators for the $\beta$-mixing coefficients into a sufficient trajectory length in order to estimate $t^{\sharp}_{\mathsf{mix}}(\Theta(\xi))$.
+[…]
+Figure 2: Average-mixing time $\xi(1\pm\varepsilon)$ band.
+Lemma 3.2.
+Let $\delta,\xi,\varepsilon\in(0,1)$ and $n\in\mathbb{N}$.
+Suppose that for any $s\in\mathbb{N}$, there exists an estimator $\widehat{\beta}(s)$ for $\beta(s)$ such that for any $\pi$-stationary ergodic Markov chain $X_{1},\dots,X_{n}$, when $n\geq n_{0}(P,s,\varepsilon,\delta)$ it holds with probability at least $1-\delta$ that
+$\left|\widehat{\beta}(s)-\beta(s)\right|<\varepsilon.$
+Let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (10). Then for
+$n\geq\max_{1\leq s\leq t^{\sharp}_{\mathsf{mix}}(\xi(1-\varepsilon))}n_{0}\left(P,s,\xi\varepsilon,\frac{\delta}{2t^{\sharp}_{\mathsf{mix}}(\xi)}\right),$
+with probability at least $1-\delta$, it holds that
+$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
+Theorem 3.5.
+Let $\delta,\xi,\varepsilon\in(0,1)$, $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (10).
+There exists a universal constant $C\in\mathbb{R}_{+}$ such that the following holds.
+Let $X_{1},X_{2},\dots,X_{n}$ be a $\pi$-stationary uniformly ergodic Markov with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$.
+For any $p\in\mathbb{R}_{+}\cup\{\infty\}$, we write
+$\mathcal{J}_{p,\xi}\doteq\sup\left\{\mathcal{J}_{p}^{(s)}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\}.$
+  1. 1.
+    When $P$ is ergodic, and
+    $n\geq 1+C_{\mathsf{erg.}}^{2}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right)\frac{1}{\xi^{4}\varepsilon^{4}}\min_{p\geq 1}\left\{\mathcal{B}_{p}\mathcal{J}_{p,\xi(1-\varepsilon)}\right\}\log\frac{4t^{\sharp}_{\mathsf{mix}}(\xi)}{\delta},\qquad\xi(\varepsilon,\delta)=\frac{\varepsilon^{2}\delta}{C_{\mathsf{erg.}}\log(1/\delta)},$
+  2. 2.
+    or when $P$ is uniformly ergodic, and
+    $n\geq 1+4t_{\mathsf{mix}}\frac{1+\left\lceil\log_{2}\frac{1}{\xi(1-\varepsilon)}\right\rceil}{\xi^{2}\varepsilon^{2}}\max\left\{C_{\mathsf{u.erg.}}^{2}\mathcal{J}_{\infty,\xi(1-\varepsilon)},576\log\frac{4t_{\mathsf{mix}}\left\lceil\log_{2}1/\xi\right\rceil}{\delta}\right\},$
+then with probability at least $1-\delta$, it holds that
+$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
+Proof.
+The theorem follows from plugging Theorem 3.2 and Theorem 3.4 into Lemma 3.2. The theorem in the uniformly ergodic case follows from
+$t^{\sharp}_{\mathsf{mix}}(\xi)\leq t_{\mathsf{mix}}(\xi)\leq t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil.$
+Corollary 3.1 (Confidence interval for $t^{\sharp}_{\mathsf{mix}}(\xi)$).
+As a consequence of Theorem 3.5, we can construct a confidence interval for $t^{\sharp}_{\mathsf{mix}}(\xi)$. Denoting $n_{0}(\xi,\delta)$ the sample complexity in Theorem 3.5, when
+$n\geq\max\left\{n_{0}\left(\frac{\xi}{1-\varepsilon},\delta/2\right),n_{0}\left(\frac{\xi}{1+\varepsilon},\delta/2\right)\right\},$
+it holds with probability $1-\delta$ that
+$t^{\sharp}_{\mathsf{mix}}(\xi)\in\left[\widehat{t_{\mathsf{mix}}}^{\sharp}\left(\frac{\xi}{1+\varepsilon}\right),\widehat{t_{\mathsf{mix}}}^{\sharp}\left(\frac{\xi}{1-\varepsilon}\right)\right].$
+In the non uniformly ergodic setting, our method introduces a trade-off parameter $p$ to obtain an upper bound of order
+$\widetilde{\mathcal{O}}\left(\frac{t^{\sharp}_{\mathsf{mix}}}{\xi^{4}}\inf_{p\geq 1}\left\{\mathcal{B}_{p}\mathcal{J}_{p,\xi}\right\}\right).$
+This enables the use of versatile bounds on the entropic term, which we explore in Section 4.
+Currently, there are no related work to which we can compare the general results obtained in Section 3.2.
+We defer a comparison to the state of the art in the finite space setting to Section 4.3.
+4  Implications at various scales of state space
+Recall (refer to Theorem 3.1, Theorem 3.2) that we require a bound on the entropic term
+$\mathcal{J}_{p}^{(s)}=\left\|Q^{(s)}\right\|^{1-1/p}_{(1-1/p)/2},$
+where for $s\in\mathbb{N}$ and for $x,x^{\prime}\in\mathcal{X}$, we wrote $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$, and for $q\in\mathbb{R}_{+}$,
+$\left\|Q^{(s)}\right\|^{q}_{q}\doteq\sum_{(x,x^{\prime})\in\mathcal{X}^{2}}Q^{(s)}(x,x^{\prime})^{q},$
+in order to recover an estimation rate for the individual $\beta$-mixing coefficients, whilst upper bounds on the sample complexity for estimating the average-mixing time (Theorem 3.5) require a control which holds uniformly over all skipping rates up to the order of the average-mixing time,
+$\mathcal{J}_{p,\xi}=\sup\left\{\mathcal{J}^{(s)}_{p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\}.$ (11)
+Bounds generally depend on the tuning parameter $p$, which can then be appropriately selected—for instance in Theorem 3.5—in order to optimize the upper bound on the sample complexity.
+Unfortunately, $\mathcal{J}_{p,\xi}$ is generally not easily amenable to analysis; hence, our focus lies in establishing bounds for its evaluation.
+It is not necessarily the case that $\mathcal{J}_{p}^{(s)}$ is monotonous in $s$.
+As illustrated in Figure 3, $\mathcal{J}_{p}^{(s)}$ can exhibit a transient regime.
+In this case, under ergodicity of $P$, the pair-wise limit
+$\lim_{s\to\infty}Q^{(s)}(x,x^{\prime})=\pi(x)\pi(x^{\prime}),\qquad\forall(x,x^{\prime})\in\mathcal{X}^{2},$
+hints at a heuristic lower bound in $\left\|\pi\right\|_{(1-1/p)/2}^{2(1-1/p)}$ on $\mathcal{J}_{p,\xi}$.
+Using the stationary regime as a baseline, it seems natural to examine whether the entropic term undergoes relative explosive growth in the transient regime.
+For instance, for some multiplicative factor $K\geq 1$, an upper bound
+$\mathcal{J}_{p}^{(s)}\leq K\left\|\pi\right\|^{2(1-1/p)}_{(1-1/p)/2}$, (12)
+with a controlled $K$ readily streamlines the sample complexities appearing in Section 3 in terms of properties of the stationary distribution.
+Let us illustrate the approach by assuming the existence of a uniformly dominating distribution over all conditional distributions.
+Formally, let us suppose that there exists a universal constant $C\in\mathbb{R}_{+}$ and a probability measure $\nu\in\mathcal{P}(\mathcal{X})$, such that for all $x^{\prime}\in\mathcal{X}$,
+$\sup_{x\in\mathcal{X}}P(x,x^{\prime})\leq C\nu(x^{\prime}).$
+It follows that
+$\begin{split}Q^{(s)}(x,x^{\prime})&\leq\pi(x)\sum_{x^{\prime\prime}\in\mathcal{X}}C\nu(x^{\prime})P^{s-1}(x,x^{\prime\prime})=C\pi(x)\nu(x^{\prime}),\end{split}$
+and thus
+$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq C^{1-1/p}\left\|\nu\right\|_{(1-1/p)/2}^{1-1/p}\left\|\pi\right\|_{(1-1/p)/2}^{1-1/p}.$
+In the rest of the section, we proceed to discuss bounds on the entropic quantities $\mathcal{J}_{p}^{(s)}$ and $\mathcal{J}_{p,\xi}$ in terms of ergodic properties of the chain, structural properties of the connection graph of the Markov chain, or the alphabet size.
+We conclude with a discussion on the binary state space, where the dominant complexity stems from the mixing bottlenecks, revealing that our approach holds implications at every scale.
+[…]
+Figure 3: Transient regime for the entropic term $\mathcal{J}_{p}^{(s)}$.
+4.1  Ergodic properties
+In this subsection, we derive a control of $\mathcal{J}_{p}^{(s)}$ based on a decomposition and natural ergodic properties of the chain that we introduce.
+Let us first decompose $\sqrt{\mathcal{J}_{p}^{(s)}}$ in terms of a stationary quantity and some measure of distance to stationarity.
+Recall that for $a,b>0,q\in(0,1)$, it holds that $(a+b)^{q}\leq a^{q}+b^{q}$, thus
+$\sqrt{\mathcal{J}_{p}^{(s)}}$ $\leq\beta_{(1-1/p)/2}(s)+\left\|\pi\right\|_{(1-1/p)/2}^{1-1/p},$
+where for $q\in(0,1)$, $\beta_{q}(s)\geq\beta(s)$ is a generalization of the $\beta$-mixing coefficient defined as
+[…]
+Moving forward, for a function $V\colon\mathcal{X}\to[1,\infty)$ we introduce the notion of the $V_{q}$-norm of a signed measure as follows,
+$\left\|\mu\right\|_{q,V}^{q}\doteq\sum_{x\in\mathcal{X}}\left|\mu(x)\right|^{q}V(x),$
+and we say that $P$ is $V_{q}$-geometrically ergodic when there exists $\rho_{q,V}\in(0,1)$ such that for all $s\in\mathbb{N}$,
+$\sup_{x\in\mathcal{X}}\frac{\left\|e_{x}P^{s}-\pi\right\|^{q}_{q,V}}{V(x)}\leq\rho_{q,V}^{s}.$ (13)
+Under this assumption the coefficients $\beta(s)$ and $\beta_{q}(s)$ can be bounded as follows,
+[…]
+thus the average-mixing time satisfies,
+$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\cfrac{\log\left\|\pi\right\|_{q,V}^{q}+\log\frac{1}{2\xi}}{\log\rho_{q,V}^{-1}},$
+As a result, writing for simplicity, $\Pi_{r,V}\doteq\left\|\pi\right\|_{r,V}^{r}$ for $r\in(0,1)$, for any $q\in(0,(1-1/p)/2]$ we can bound the entropic term as follows,
+$\begin{split}\mathcal{J}_{p}^{(s)}\leq\left(\rho_{q,V}^{s}\Pi_{q,V}+\Pi_{(1-1/p)/2,1}^{2}\right)^{2},\end{split}$
+leading to the following theorem.
+Theorem 4.1 (Under $V_{q}$-geometrical ergodicity).
+Let $p\geq 1$. If there exists $q\in(0,(1-1/p)/2]$ such $P$ is $V_{q}$-geometrically ergodic—refer to (13)—for any $s\in\mathbb{N}$, it holds that
+$\begin{split}\mathcal{J}_{p}^{(s)}\leq\left(\rho_{q,V}^{s}\Pi_{q,V}+\Pi_{(1-1/p)/2,1}^{2}\right)^{2}.\end{split}$
+Observe from the above that, as soon as
+[…]
+we obtain a control of the entropic term in terms of a purely stationary quantity
+$\begin{split}\mathcal{J}_{p}^{(s)}\leq 4\Pi_{(1-1/p)/2,1}^{4},\end{split}$
+giving us insight into the exit time of the transient regime illustrated in Figure 3.
+Finally, a similar proof allows us to analyze the entropic term under the following condition, which is much weaker than $V$-geometrical ergodicity.
+Theorem 4.2.
+Suppose that there exist $\rho\in(0,1)$ and $V\colon\mathcal{X}\to[1,\infty)$ such that for any $s\in\mathbb{R}$,
+$\sup_{x,x^{\prime}}\left|P^{s}(x,x^{\prime})-\pi(x^{\prime})\right|\frac{V(x^{\prime})}{V(x)}\leq\rho^{s}.$
+Then for all $p\geq 1$ and for all $s\in\mathbb{N}$, it holds that
+$\begin{split}\mathcal{J}_{p}^{(s)}\leq\left(\rho^{s(1-1/p)/2}\left\|\pi\odot V\right\|_{(1-1/p)/2}^{(1-1/p)/2}\left\|1/V\right\|_{(1-1/p)/2}^{(1-1/p)/2}+\Pi_{(1-1/p)/2,1}^{2}\right)^{2}.\end{split}$
+Remark 4.1.
+Obvious summability conditions on $\left\|1/V\right\|_{(1-1/p)/2}$ and $\left\|\pi\odot V\right\|_{(1-1/p)/2}$ are necessary for the bound in the above theorem to be non-vacuous.
+Additionally, under the condition of Theorem 4.2 a bound on the $\beta$-mixing coefficient can be immediately obtained as follows,
+$\beta(s)\leq\frac{\rho^{s}}{2}\left\|\pi\odot V\right\|_{1}\left\|1/V\right\|_{1}.$
+4.2  Infinite graphs with controlled growth
+Natural structural assumptions on the underlying connection graph of a Markov chain also enable us to control the entropic term.
+Recall that an ergodic transition operator $P\in\mathcal{W}(\mathcal{X})$ induces a fully connected digraph $(\mathcal{X},\mathcal{E})$, where
+$\mathcal{E}\doteq\left\{(x,x^{\prime})\in\mathcal{X}^{2}\colon P(x,x^{\prime})>0\right\}$
+is the set of directed edges.
+The growth of a graph (Trofimov, 1985; Campbell et al., 2023) is defined as a function $g\colon\mathbb{N}\to\mathbb{N}\cup\left\{\infty\right\}$, where for $s\in\mathbb{N}$,
+$g(s)\doteq\sup_{\mathcal{S}\subset\mathcal{X},\rho\left((\mathcal{X},\mathcal{E})|_{\mathcal{S}}\right)\leq s}\left|\mathcal{S}\right|,$
+where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the diameter of a directed graph, which is the maximum eccentricity of any vertex in the graph with the convention that a disconnected graph has infinite diameter, $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.
+Condition 1 (Connection graph with polynomial growth).
+A graph is said to have at most polynomial growth when there exists $G\in\mathbb{R}_{+}$ and $q\geq 1$ such that for any $s\in\mathbb{N}$, it holds that $g(s)\leq Gs^{q}$.
+Then $q$ and $G$ are respectively called the degree and factor of growth.
+Example 4.1 (Birth and death chains).
+It is immediate from its definition that the connection graph associated with the Chebyshev type random walk developed in Example 2.1 has bounded growth with $G=2$ and $q=1$.
+Lemma 4.1.
+Let $p\in\mathbb{R}_{+}\cup\{+\infty\},q\in[1,\infty),\xi\in(0,1)$ and let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator.
+If the connection graph of $P$ has polynomial growth of degree $q$ with factor $G$, it holds that
+$\mathcal{J}_{p,\xi}\leq G^{1+1/p}t^{\sharp}_{\mathsf{mix}}(\xi)^{q(1+1/p)}\left\|\pi\right\|^{1-1/p}_{(1-1/p)/2}.$
+Proof.
+For $x\in\mathcal{X}$, let us denote $\mathcal{S}_{x}^{(s)}\subset\mathcal{X}$ the set of states accessible from $x$ in $s$ steps. We have that
+[…]
+where the first inequality stems from the uniform distribution maximizing the $\ell_{(1-1/p)/2}$ quasi-norm, and the second inequality follows from the polynomial growth assumption. ∎
+Example 4.2 (Graph with polynomial growth and stationary distribution with polynomially decaying tail).
+We let $\mathcal{X}=\mathbb{N}$, and suppose that there exists $C\in\mathbb{R}_{+}$, and $\alpha\in(1,\infty)$ such that $\pi(x)\leq\frac{C}{x^{1+\alpha}}$ for any $x\in\mathcal{X}$. It holds that
+$\sum_{x\in\mathcal{X}}\pi(x)^{(1-1/p)/2}\leq C^{(1-1/p)/2}\zeta((1+\alpha)(1-1/p)/2),$
+where $\zeta$ is the Riemann Zeta function.
+  1. 1.
+    In the uniformly ergodic setting, Lemma 4.1 specializes to
+    $\mathcal{J}_{\infty,\xi}\leq GCt^{\sharp}_{\mathsf{mix}}(\xi)^{q}\zeta((1+\alpha)/2)^{2},$
+    and treating $C,G,\alpha$ as constants, an upper bound on the sample complexity for a $(1/4,\delta)$-PAC bound on $t^{\sharp}_{\mathsf{mix}}(\xi)$ is given by
+    $n\geq\widetilde{\mathcal{O}}\left(t_{\mathsf{mix}}^{q+1}/\xi^{2}\right).$
+  2. 2.
+    In the non-uniformly setting, we may instead optimize the following product
+    $\min_{p\geq 1}\left\{\mathcal{B}_{p}G^{1+1/p}t^{\sharp}_{\mathsf{mix}}(\xi)^{q(1+1/p)}\left\|\pi\right\|^{1-1/p}_{(1-1/p)/2}\right\}.$
+4.3  Finite space
+When the state space $\mathcal{X}$ is finite and $P$ is supported on $(\mathcal{X},\mathcal{E})$, the graph-related methods discussed in the previous section remain applicable.
+Alternatively, a cruder—graph agnostic—approach yields the following bound,
+$\sqrt{\mathcal{J}_{\infty}^{(s)}}\leq\sum_{(x,x^{\prime})\in\mathcal{E}}\left(\frac{1}{\left|\mathcal{X}\right|^{2}}\right)^{1/2}=\left|\mathcal{X}\right|.$
+Note that (4.3) matches the heuristic lower bound $\left\|\pi\right\|_{(1-1/p)/2}^{2(1-1/p)}$ in the worst-case where $P$ is bistochastic.
+Corollary 4.1 (of Theorem 3.5 for $\left|\mathcal{X}\right|<\infty$).
+Let $\mathcal{X}$ be a finite space.
+Let $\varepsilon\in(0,1/4)$, and $\delta,\xi\in(0,1)$.
+Let $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (10).
+For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$.
+When
+$n\geq 1+4t_{\mathsf{mix}}\frac{1+\left\lceil\log_{2}\frac{1}{\xi(1-\varepsilon)}\right\rceil}{\xi^{2}\varepsilon^{2}}\max\left\{C_{\mathsf{u.erg.}}^{2}\left|\mathcal{X}\right|^{2},576\log\frac{4t_{\mathsf{mix}}\left\lceil\log_{2}1/\xi\right\rceil}{\delta}\right\},$
+with probability at least $1-\delta$, it holds that
+$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
+Comparison with related work
+Since the criteria for convergence are different, the problems of estimating the relaxation or worst-case mixing time of a chain to multiplicative accuracy and the problem at hand are not strictly speaking comparable, but it is instructive to inspect the sample complexities.
+The problem of estimating the relaxation time of a Markov chain to multiplicative error has sample complexity of the order of $\widetilde{\Theta}(t_{\mathsf{mix}}/\pi_{\star})$ (Hsu et al., 2019; Wolfer and Kontorovich, 2024), while a known upper bound on the complexity of estimating the mixing time is of the order of $\widetilde{\mathcal{O}}(\left|\mathcal{X}\right|t_{\mathsf{mix}}/\pi_{\star})$ (Wolfer, 2024).
+Crucially, the above-mentioned problem all suffer a dependency on the minimum stationary probability $\pi_{\star}$.
+When $\pi_{\star}\propto\exp(-\left|\mathcal{X}\right|)$, the problem becomes intractable.
+Corollary 4.1 yields the superior upper bound $\widetilde{\mathcal{O}}(t_{\mathsf{mix}}\left|\mathcal{X}\right|^{2}/\xi^{2})$, agnostic of $\pi_{\star}$ and enjoying an at most linear dependency on the number of parameters of the transition operator.
+4.4  The two-point space
+In this section, we analyze the case where $\left|\mathcal{X}\right|=2$, illustrating the possibly stark difference between the worst-case mixing time and the average-mixing time.
+For $p,q\in(0,1)$, we let
+$P_{p,q}=\begin{pmatrix}1-p&p\\ q&1-q\end{pmatrix}$.
+Quantities such as $d$ or $d^{\sharp}$ subscripted with $p,q$ will be understood as pertaining to $P_{p,q}$. The matrix $P_{p,q}$ is irreducible, and its stationary distribution is readily given by
+$\pi_{p,q}^{\intercal}=\frac{1}{p+q}\begin{pmatrix}q\\ p\end{pmatrix}$.
+It follows from a direct computation (see e.g. the proof of Wolfer (2024, Lemma 8.1)) that
+[…]
+and as a result,
+$\begin{split}d_{p,q}^{\sharp}(t)&=2\frac{pq}{(p+q)^{2}}\left|1-p-q\right|^{t},\hskip 18.49988ptd_{p,q}(t)=\frac{p\lor q}{p+q}\left|1-p-q\right|^{t}.\end{split}$
+When $p=q$, the stationary distribution is uniform, $d_{p,q}^{\sharp}(t)=d_{p,q}(t)$, and
+$t_{\mathsf{mix}}(\xi)=t_{\mathsf{mix}}^{\sharp}(\xi).$
+However, for $\eta\doteq p/q<1$, the ratio of the distances is a function of the parameter $\eta$,
+$\frac{d_{p,q}(t)}{d^{\sharp}_{p,q}(t)}=\frac{1+\eta}{2\eta}\doteq r_{\eta}(t),$
+whence the following lemma holds.
+Lemma 4.2.
+There exists a family of Markov chains
+$\left\{P_{p,q}\colon(p,q)\in(0,1)^{2}\right\}\in\mathcal{W}^{\star}(\left\{0,1\right\})$
+such that for any $\xi\in(0,1)$ and any $M>0$, there exist $p,q\in(0,1)$ such that
+$t_{\mathsf{mix}}(\xi)>Mt_{\mathsf{mix}}^{\sharp}(\xi).$
+5  Conclusion
+To assess convergence in Markov chains from a single trajectory of observations, recent literature has focused on estimating the worst-case mixing time or the spectral gaps associated with the transition operator.
+These mixing parameters lead to concentration inequalities that are almost as powerful as those in the iid setting.
+However, their existence requires strong assumptions on the chain such as uniform or geometric ergodicity, and they are notoriously hard to estimate from the data, with known lower bounds inversely proportional to the minimum stationary probability $\pi_{\star}$.
+As a result, the estimation problem is only statistically tractable for moderately small space sizes and when the tail of the stationary distribution does not decay too rapidly.
+In this paper, we relaxed these necessary conditions and limitations by considering the average-mixing time instead.
+Over finite spaces, our analysis showed that the average-mixing time is easier to estimate than its worst-case counterpart, and we could obtain an upper bound on the sample complexity of the problem that is independent of the rate of decay of the stationary distribution.
+Moreover, in contradistinction to worst-case mixing time estimation, our analysis extends to countable state spaces and sub-geometrically $\beta$-mixing Markov chains under natural structural assumptions.
+The trade-off for accessing an easier convergence estimation rate is a logarithmic degradation in the available concentration inequalities.
+Appendix A Extension to non-stationary Markov chains
+Although our results are generally stated assuming a stationary start, it is possible to generalize them to non-stationary chains as follows.
+Let $\mu\in\mathcal{P}(\mathcal{X})$ absolutely continuous with respect to $\pi$.
+Inspired by the proof of Paulin (2015, Proposition 3.10), an application of Hölder’s inequality with $q,r\geq 1$ with $1/q+1/r=1$ yields,
+$\mathbb{E}_{\mu}f(X_{1},\dots,X_{n})\leq\left\|\mu(X)/\pi(X)\right\|_{\pi,q}\left\|f(X_{1},\dots,X_{n})\right\|_{\pi,r},$
+and in particular, for any $\varepsilon>0$, we can bound the probability of an $\varepsilon$-deviation assuming a non-stationary start as
+$\mathbb{P}_{\mu}\left(f(X_{1},\dots,X_{n})>\varepsilon\right)\leq\left\|\mu(X)/\pi(X)\right\|_{\pi,q}\mathbb{P}_{\pi}\left(f(X_{1},\dots,X_{n})>\varepsilon\right)^{1/r}.$
+Setting $q=r=2$ recovers Paulin (2015, Proposition 3.10). The reader is invited to consult (Paulin, 2015, Section 3.3) for additional methods involving “burning” a prefix of the observed trajectory.
+Appendix B Connection with $\beta$-mixing
 Recall that for a process $\{X_{t}\}_{t\in\mathbb{Z}}$ on $\mathcal{X}$, for any $s\in\mathbb{Z}$, the $\beta$-mixing coefficient (Bradley, 2005; Doukhan, 2012) is defined by
@@
 $\lim_{s\to\infty}\beta(s)=0.$
-Recall that a stationary, countable-state Markov chain $X$ is $\beta$-mixing if and only if it is irreducible and aperiodic (Bradley, 2005, Theorem 3.2).
-Lemma 2.2 (Davydov (1974, Proposition 1)).
+In particular, a stationary countable-state Markov chain $X$ is $\beta$-mixing if and only if it is irreducible and aperiodic (Bradley, 2005, Theorem 3.2).
+Lemma B.1 (Davydov (1974, Proposition 1)).
 We let $\{X_{t}\}_{t\in\mathbb{N}}$ be a countable-state Markov chain with transition operator $P$.
 Writing $\mu^{(0)}$ for the initial distribution, and for $t>1$, $\mu^{(t)}=\mu^{(0)}P^{t}$ for the marginal distribution after time step $t$, it holds that
-$\beta(s)=\sup_{t\in\mathbb{N}}\sum_{x\in\mathcal{X}}\mu^{(t)}(x)\left\|e_{x}P^{s}-\mu^{(s+t)}\right\|_{\mathsf{TV}}.\$
+$\beta(s)=\sup_{t\in\mathbb{N}}\sum_{x\in\mathcal{X}}\mu^{(t)}(x)\left\|e_{x}P^{s}-\mu^{(s+t)}\right\|_{\mathsf{TV}}.\\$
 It follows from the above lemma that when the chain is stationary, $\mu^{(0)}=\pi=\mu^{(t)}$ for any $t\in\mathbb{N}$, and
-$\begin{split}\beta(s)=\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{s}-\pi\right\|_{\mathsf{TV}}=d^{\sharp}(s),\\ \end{split}$
-thus the average-mixing time is essentially the “stationary $\beta$-mixing time” of the chain555 Similarly, the worst-case mixing time can be connected to the notion of $\phi$-mixing (Davydov, 1974).
+[…]
+thus the average-mixing time is essentially the “stationary $\beta$-mixing time” of the chain [^2] Similarly, the worst-case mixing time is directly connected to the notion of $\phi$-mixing (Davydov, 1974).
 .
@@
 $\beta(0)=1-\left\|\pi\right\|_{2}^{2},$
-which depends only on the stationary properties of the chain.
-When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially.
-When there exists $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially when $b\neq 1$, and exponentially when $b=1$.
-In Section 6, we will introduce a class of Markov transition operators, whose flavor of mixing (algebraic, sub-exponential, exponential) can be fully controlled.
-Notably (refer to the proof of Theorem 3.2), when $P$ is uniformly ergodic with mixing time $t_{\mathsf{mix}}$, it holds that
-$\beta(s)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$ (4)
-hence the process is exponentially $\beta$-mixing with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$.
-However, nothing a priori precludes the $\beta$-mixing rate from being superior to (4).
-The connection between the average-mixing time and $\beta$-mixing provides additional motivation for the problem of estimating the $\beta$-mixing coefficients, whose knowledge is known to lead to powerful decoupling techniques (Yu, 1994).
-The problem of inferring mixing coefficients from a single trajectory of observations has recently received attention from the community.
-In the most recent work (Khaleghi and Lugosi, 2023), consistent estimators for $\beta$- and $\alpha$-mixing coefficients were proposed.
-In Section 3, we will provide more explicit and finer finite-sample bounds for the estimation problem in the specialized Markovian setting.
-A regime of particular interest to us is when the chain $\beta$-mixes much faster than $\Theta(t_{\mathsf{mix}})$.
-Henceforth in this manuscript, when they coincide under stationarity, we may use the notations $\beta$ and $d^{\sharp}$ interchangeably.
-2.4   Bounding the deviation of additive functionals evaluated on Markov chains
-We now illustrate how the average-mixing time, through its connection to $\beta$-mixing naturally appears as a quantity of interest when bounding deviations of functions evaluated on a trajectory sampled from a Markov chain.
-More formally, let $X_{1},\dots,X_{n}$ be a Markov chain over a countable space $\mathcal{X}$, and whose dynamics are governed by $P\in\mathcal{W}(\mathcal{X})$ with stationary distribution $\pi$.
-For simplicity, we consider a real bounded function $f$ on $\mathcal{X}$, which is centered in the sense where its stationary expected value vanishes $\mathbb{E}_{\pi}f=0$.
-The ergodic theorem states that the empirical mean taken on an infinite trajectory will similarly vanish,
-$\lim_{n\to\infty}\frac{1}{n}\sum_{t=1}^{n}f(X_{t})=0.$
-In practice however, we are more likely to be interested in controlling the deviation of the sample mean from its expected value in a non-asymptotic setting.
-Lemma 2.3 (Bounding deviation of additive functionals evaluated on Markov chains).
-Let $f\colon\mathcal{X}\to[-1,1]$ with $\mathbb{E}_{\pi}f=0$, let $\varepsilon,\delta\in(0,1)$, and let $t^{\sharp}_{\mathsf{mix}}\colon(0,1)\to\mathbb{N}$.
-For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, and average-mixing time bounded from above by $t^{\sharp}_{\mathsf{mix}}$, the following statements hold.
-  1. 1.
-    (Sub-exponential) If there exist $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ with $\beta_{0}\geq 1$, $b\in(0,1]$, such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$, then there exists a universal constant $C\in\mathbb{R}_{+}$ such that for
-    $n\geq\frac{C}{\varepsilon^{2}}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right)\log\left(\frac{1}{\delta}\right),$
-    with
-    $\xi(\varepsilon,\delta)\doteq\frac{\delta\varepsilon^{2}}{16\log\left(\frac{4}{\delta}\right)},$
-    it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$. The above upper on the path length can be further streamlined as
-    $n\geq\frac{C}{\varepsilon^{2}}\frac{1}{b}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\delta/e\right)\log\left(\frac{1}{\delta}\right)\log\left(\frac{\log(1/\delta)}{\varepsilon}\right),$
-    for some universal constant $C\in\mathbb{R}_{+}$.
-  2. 2.
-    (Polynomial) If there exist $\beta_{1},b\in\mathbb{R}_{+}$ with $\beta_{1}\geq 1$, such that for any $s\in\mathbb{N}$, $\beta(s)\leq\frac{\beta_{1}}{s^{b}}$, then there exists a universal constant $C\in\mathbb{R}_{+}$ such that for
-    $n\geq C\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(2t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\right)^{\frac{1}{b}},$
-    it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$.
-Proof sketch.
-We adopt the traditional method as described in Yu (1994).
-This approach combines a blocking technique, credited to Bernstein (1927) in the context of $m$-dependent processes, together with an application of Hoeffding’s inequality to control the deviation probability of a sum of independent observations.
-We further optimize for the block size and reformulate the results to obtain an expression involving the average-mixing time under various rates of $\beta$-mixing.
-Refer to Section 5.1 for the proof.
-∎
-As a consequence of Lemma 2.3, in both the sub-exponential and polynomial settings, provided an assumption on the convergence order $b$, estimates on the average-mixing time to proximity $\Theta(\delta)$ yield corresponding deviation bounds.
-In the uniformly ergodic setting, corresponding to the assumption of the exponential setting ($b=1$), we obtain a deviation inequality involving the average-mixing time in lieu of the mixing time (Chung et al., 2012, Theorem 3.1).
-The lemma generalizes to non-centered functions, lower deviations and arbitrarily bounded functions.
-It is also noteworthy that in the sub-exponential setting, the convergence order $b$ solely affects the constant in the sample complexity.
-Our formulation above serves as additional motivation for analyzing and inferring the average-mixing time.
-3   Estimation from a single trajectory
-In this section, we analyze the estimation of $\beta$-mixing coefficients and the average-mixing time from a single trajectory of observations.
-3.1   Estimation of $\beta$-mixing coefficients under general ergodicity condition
-In this section we analyse plug-in estimators for $\beta$-mixing coefficients.
-We recall that in this general setting, the mixing time can be infinite, and the $\beta$-mixing coefficients can decay at arbitrary slow rate.
-We will focus on exponential, sub-exponential and polynomial rates of convergence.
-Our analysis relies on stationary skipped trajectories, where for $s\in\mathbb{N},s<n$, we denote
-$\begin{split}X^{(s)}\doteq X_{1},X_{1+s},X_{1+2s},\dots,X_{1+\lfloor(n-1)/s\rfloor s},\end{split}$
-with transition operator $P^{s}$ and initial distribution $\pi$. It will be convenient to define the following counting random variables,
-$\begin{split}N_{x}^{(s)}&\doteq\sum_{t=1}^{\lfloor(n-1)/s\rfloor}\delta\left[X_{1+s(t-1)}=x\right],\\ N_{xx^{\prime}}^{(s)}&\doteq\sum_{t=1}^{\lfloor(n-1)/s\rfloor}\delta\left[X_{1+s(t-1)}=x,X_{1+st}=x^{\prime}\right].\end{split}$
-We use the shortcut notation $N_{x}=N_{x}^{(1)}$ and $N_{xx^{\prime}}=N_{xx^{\prime}}^{(1)}$ to respectively denote number of visits to state $x$ and transitions from $x$ to $x^{\prime}$ on the original chain.
-We define the following estimator $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ for $\beta(s)$,
-$\widehat{\beta}(s)=\frac{1}{2\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|,$ (5)
-and by convention, we set $\widehat{\beta}(s)=0$ for $s\geq n$.
-Theorem 3.1.
-Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5).
-For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, it holds that
-$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{\mathcal{B}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}}\left(\sqrt{\frac{3}{2}\mathcal{J}_{p}^{(s)}}+2\sqrt{\left\|\pi\right\|_{1-1/p}^{1-1/p}}\right)\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
-where for $x,x^{\prime}\in\mathcal{X}$, we wrote $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$, and
-$\mathcal{B}_{p}^{(s)}\doteq\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p},\qquad\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}.$
-We observe that the two key quantities $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$ control the estimation rate.
-The former essentially captures $\beta$-mixing properties of the process, while the latter is tightly connected to entropic properties of the distribution of (skipped) stationary pairs.
-In the following Corollary 3.1, we readily provide more explicit upper bounds on $\mathcal{B}_{p}^{(s)}$ under additional assumptions on the mixing rate.
-The analysis of $\mathcal{J}_{p}^{(s)}$ is deferred to Section 4.
-Corollary 3.1.
-Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5).
-For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$ with $\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}<\infty$, the following statements hold.
-  1. 1.
-    (Exponential) If $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s)$, with $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$ and $t\in\mathbb{N}\cup\left\{0\right\}$,
-    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}\left(\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}\right)\mathcal{J}_{p}^{(s)}}.\\ \end{split}$
-  2. 2.
-    (Sub-exponential) If $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$, with $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$, $b\in(0,1)$, and $s\in\mathbb{N}\cup\left\{0\right\}$.
-    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{\beta_{0}^{1/p}}{\lfloor(n-1)/s\rfloor}\left(1+\exp(-\beta_{1}s^{b}/p)+\frac{\Gamma\left(\frac{1}{b},s^{b}\beta_{1}/p\right)}{bs(\beta_{1}/p)^{1/b}}\right)\mathcal{J}_{p}^{(s)}},\end{split}$
-[…]
-  3. 3.
-    (Polynomial) If $\beta(s)\leq\beta_{1}/s^{b}$, with $\beta_{1}\in\mathbb{R}_{+}$, $b\in(1,\infty)$, $b/p\in(1,\infty)$, $s\in\mathbb{N}$ and $\beta(0)=\beta_{0}\in\mathbb{R}_{+}$,
-    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{\beta_{0}^{1/p}+\zeta(b/p)\frac{\beta_{1}^{1/p}}{s^{b/p}}}{\lfloor(n-1)/s\rfloor}\mathcal{J}_{p}^{(s)}},\end{split}$
-    where $\zeta$ is the Riemann zeta function defined by $\zeta(r)=\sum_{t=1}^{\infty}t^{-r}$.
-3.2   Comparison with prior results
-In McDonald et al. (2015, Theorem 4), the authors establish the nearly parametric convergence rate for continuously valued Markov processes
-$\mathbb{E}{\left|\widehat{\beta}(s)-\beta(s)\right|}\leq\mathcal{O}\left(\sqrt{\frac{W_{0}(n)}{n}}\right),$
-where $W_{0}$ is the Lambert function (Corless et al., 1996).
-They also posit that their findings may be considered conservative when applied to discrete-valued processes.
-Theorem 3.1 validates their remark by recovering the superior parametric rate $\mathcal{O}(\sqrt{1/n})$ under convergence of $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$.
-Our method shares similarities with Khaleghi and Lugosi (2023), as both employ Rio’s covariance inequality Rio (1999, Corollary 1.1).
-However, compared to both Khaleghi and Lugosi (2023) and McDonald et al.
-(2015, Theorem 4), we distinguish ourselves by focusing on the discrete space setting and deriving explicit rates in terms of fine-grained properties of the transition operator.
-3.3   Estimation of $\beta$-mixing coefficients under uniform ergodicity
-Under uniform ergodicity, the mixing time is finite, and the $\beta$-mixing coefficients decays exponentially.
-In this setting, we show how to obtain a sample complexity upper bound for high-probability estimates on the individual $\beta$-mixing coefficients.
-Theorem 3.2.
-Let $s,n\in\mathbb{N}$ with $n>2s+1$, let $p\in\mathbb{R}_{+},p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5).
-For any uniformly ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$,
-  1. $(i)$
-    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 8\sqrt{\frac{1}{n-1}\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}},\\ \end{split}$
-    where $t_{\mathsf{mix}}$ is the mixing time of $P$.
-  2. $(ii)$
-    Let $\varepsilon,\delta\in(0,1)$. For
-    $n\geq 1+\frac{2^{8}}{\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{2}{\delta}\right\},$
-    with probability at least $1-\delta$, it holds that
-    $\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon.$
-Using a truncated estimator with a cut-off determined from a uniform bound on the complexity terms $\|Q^{(s)}\|_{1-1/p}^{1-1/p}$, we can control the sequence of $\beta$-mixing coefficients with respect to the uniform norm modulo an additional logarithmic factor in the sample complexity.
-Theorem 3.3.
-Let $\varepsilon\in(0,1/4],\delta\in(0,1),p\in\mathbb{R}_{+},p\geq 1$ and $\mathcal{J}_{p}\in\mathbb{R}_{+}$.
-Let $n,S\in\mathbb{N}$, and for $s\in\mathbb{N}$, let $\widehat{\beta}_{[S]}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined as
-$\widehat{\beta}_{[S]}(s)=\begin{cases}\widehat{\beta}(s)&\text{ for }s\leq S\\ 0&\text{ otherwise,}\end{cases}$
-where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (5).
-For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$ such that,
-$\sup_{s\in\mathbb{N}}\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\leq\mathcal{J}_{p},$
-when
-$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\varepsilon}\right\rceil\max\left\{\mathcal{J}_{p},18\log\left(\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\varepsilon}\right\rceil\right)\right\},$
-with probability at least $1-\delta$, it holds that,
-$\left\|\widehat{\beta}_{[S(n,\varepsilon,\delta,p,\mathcal{J}_{p})]}-\beta\right\|_{\infty}\leq\varepsilon,$
-where the cut-off is set as
-$S(n,\varepsilon,\delta,p,\mathcal{J}_{p})\doteq\frac{(n-1)\varepsilon^{2}}{2^{9}\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}}.$
-Remark 3.1.
-We note but do not pursue the fact that a more delicate bound on the sample complexity in Theorem 3.3 could be obtained, replacing a logarithm with a Lambert $W$ function.
-In Section 4, we show how to obtain bounds on the entropy term in several natural settings.
-3.4   Estimation of the average-mixing time
-We move on to the estimation of the average-mixing time.
-Since the average distance to stationarity $d^{\sharp}=\beta$ does not generally enjoy sub-multiplicativity (it does not hold that $t^{\sharp}_{\mathsf{mix}}(\xi)$ can be bounded by $t^{\sharp}_{\mathsf{mix}}(1/4)$ and a polynomial of $\log\frac{1}{\xi}$), we formalize desired guarantees somewhat differently from Hsu et al.
-(2019); Wolfer and Kontorovich (2024).
-Indeed, for the relaxation time $t_{\mathsf{rel}}$, we were seeking an estimator $\widehat{t_{\mathsf{rel}}}$ to multiplicative error $\varepsilon$.
-Namely, the estimators in Hsu et al.
-(2019); Wolfer and Kontorovich (2024) guarantee that with high probability,
-$\left|\frac{\widehat{t_{\mathsf{rel}}}}{t_{\mathsf{rel}}}-1\right|\leq\varepsilon.$
-Here, we keep the distance to stationarity as a free parameter, and construct an estimator for the average-mixing time for proximity parameter $\xi$ up to multiplicative accuracy $\varepsilon$.
-Define for $\xi\in(0,1)$,
-$\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\doteq\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\widehat{\beta}(s)\leq\xi\right\},$ (6)
-where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (5).
-Theorem 3.4.
-Let $\varepsilon\in(0,1/4),\delta,\xi\in(0,1),p\in\mathbb{R}_{+}$ and $\mathcal{J}_{p,\xi}\in\mathbb{R}_{+}$.
-Let $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (6).
-For any uniformly ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$ such that,
-$\sup\left\{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\}\leq\mathcal{J}_{p,\xi},$
-when
-$n\geq 1+\frac{2^{8}t_{\mathsf{mix}}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+\frac{p}{\log 2}\right)\mathcal{J}_{p,\frac{3\xi}{4}},9\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+1\right)\log\left(\frac{4t_{\mathsf{mix}}}{\delta}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil\right)\right\},$
-with probability at least $1-\delta$, it holds that
-$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
-3.5   Extension to non-stationary Markov chains
-Although results in this section are stated assuming a stationary start, it is possible to generalize them to non-stationary chains as follows.
-Let $\mu\in\mathcal{P}(\mathcal{X})$ absolutely continuous with respect to $\pi$.
-Inspired by the proof of Paulin (2015, Proposition 3.10), an application of Hölder’s inequality with $q,r\geq 1$ with $11/q+1/r=$ 1 yields,
-$\mathbb{E}_{\mu}f(X_{1},\dots,X_{n})\leq\left\|\mu(X)/\pi(X)\right\|_{\pi,q}\left\|f(X_{1},\dots,X_{n})\right\|_{\pi,r},$
-and in particular, for any $\varepsilon>0$, we can bound the probability of an $\varepsilon$-deviation assuming a non-stationary start as
-$\mathbb{P}_{\mu}\left(f(X_{1},\dots,X_{n})>\varepsilon\right)\leq\left\|\mu(X)/\pi(X)\right\|_{\pi,q}\mathbb{P}_{\pi}\left(f(X_{1},\dots,X_{n})>\varepsilon\right)^{1/r}.$
-Setting $q=r=2$ recovers Paulin (2015, Proposition 3.10). The reader is invited to consult (Paulin, 2015, Section 3.3) for additional methods involving “burning” a prefix of the observed trajectory.
-3.6   Comparison with an alternative method
-We point out the fact that for the special case $p=2$, by using the techniques in Paulin (2015), one can obtain an upper bound of the order (ignoring logarithmic factors) of
-$\frac{t_{\mathsf{mix}}}{\varepsilon^{2}\xi^{2}}\mathcal{J}_{2,\xi}$
-for the problem of estimating the average-mixing time in uniformly ergodic Markov chains.
-Indeed, the mixing time of the $s$-skipped pairs of observations is upper bounded by $t_{\mathsf{mix}}^{(s)}+s\leq\lceil t_{\mathsf{mix}}/s\rceil+s$.
-From Paulin (2015, Theorem 3.2, Proposition  3.4), it follows that
-$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]&\leq 2\lfloor(n-1)/s\rfloor\lceil t_{\mathsf{mix}}/s\rceil\pi(x),\\ \operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]&\leq 2\lfloor(n-1)/s\rfloor(\lceil t_{\mathsf{mix}}/s\rceil+s)\pi(x)P^{s}(x,x^{\prime}),\\ \end{split}$
-which yields
-$\begin{split}\mathbb{E}\left|\widehat{\beta}(s)-\beta(s)\right|\leq&\sqrt{2\frac{\lceil t_{\mathsf{mix}}/s\rceil+s}{\lfloor(n-1)/s\rfloor}}\left(\left\|Q^{(s)}\right\|_{1/2}^{1/2}+2\left\|\pi\right\|_{1/2}^{1/2}\right).\\ \end{split}$
-Our method is superior in that it introduces a trade-off parameter $p$ to obtain an upper bound of order
-$\frac{t_{\mathsf{mix}}}{\varepsilon^{2}\xi^{2}}\inf_{p\geq 1}\left\{p\mathcal{J}_{p,\xi}\right\}.$
-This enables the use of more versatile bounds on the entropic term, which we explore in Section 4.
-4   Implications at various scales of state space
-In infinitely denumerable state spaces, recall that for any $s\in\mathbb{N}$, the well-definition of the estimation rate for the $\beta(s)$-mixing coefficient (Theorem 3.1) hinges on the finiteness of the following quantity
-$\mathcal{J}^{(s)}_{p}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
-where $p\in[1,\infty)$ is a user-fixed parameter. It is instructive to observe that on one hand
-$\lim_{p\to\infty}\mathcal{J}_{p}^{(s)}=1,$
-while on the other hand, $\mathcal{B}_{p}^{(s)}$ increases with $p$,
-$\lim_{p\to\infty}\mathcal{B}_{p}^{(s)}\asymp\frac{n}{s},$
-leading to a trade-off in the choice of the parameter in Theorem 3.1. In particular, provided a control which holds uniformly over $s$,
-$\mathcal{J}_{p,\xi}\doteq\sup\left\{\mathcal{J}^{(s)}_{p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$ (7)
-the free parameter $p$ can be appropriately selected in Theorem 3.4, in order to optimize the upper bound on the sample complexity of estimating the average-mixing time.
-It is also noteworthy that since the $\ell_{1-1/p}$ quasi-norm and the $(1-1/p)$th order Rényi entropy $H_{1-1/p}$ of a distribution $\mu\in\mathcal{P}(\mathcal{X})$ are related by the expression
-$p\log\left\|\mu\right\|_{1-1/p}^{1-1/p}=H_{1-1/p}(\mu),$
-the quantity $\mathcal{J}_{p,\xi}$ pertains to a specific entropic property inherent in the Markov process.
-Unfortunately, this term is generally not easily amenable to analysis; hence, our focus lies in establishing bounds for its evaluation.
-Under ergodicity of $P$, the pair-wise limit
-$\lim_{s\to\infty}Q^{(s)}(x,x^{\prime})=\pi(x)\pi(x^{\prime}),\qquad\forall x,x^{\prime}\in\mathcal{X},$
-hints at a heuristic lower bound in $\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}$ on $\mathcal{J}_{p,\xi}$.
-Using the stationary regime as a baseline, it seems natural to examine chains where the entropic term does not undergo relative explosive growth in the transient regime.
-For instance, for some multiplicative factor $K\geq 1$, an upper bound
-$\mathcal{J}_{p}^{(s)}\leq K\left\|\pi\right\|^{2(1-1/p)}_{1-1/p},$ (8)
-with a controlled $K$ readily simplifies the sample complexities appearing in Section 3.
-In the present section, we proceed to discuss bounds on the entropic quantities $\mathcal{J}_{p}^{(s)}$ and $\mathcal{J}_{p,\xi}$ in terms of state-wise conditional entropy, structural properties of the connection graph of the Markov chain, or the alphabet size.
-We conclude with a discussion on the binary state space, where the dominant complexity stems from the mixing bottlenecks, revealing that our results hold implications at every scale.
-4.1   Ergodic properties
-In this subsection, we derive various controls of $\mathcal{J}_{p}^{(s)}$ based on ergodic properties of the chain.
-Specifically, we introduce three conditions.
-Condition 1 is the classical notion of $V$-geometric ergodicity, which is known to be weaker than uniform ergodicity.
-Condition 2 imposes a stronger criterion, under which an ergodic chain is necessarily uniformly ergodic.
-Both Conditions 1 and 2 allow to derive bounds on $\mathcal{J}_{p}^{(s)}$ that are uniform over $s$.
-However, these conditions are quite stringent.
-We introduce a natural extension of these two conditions, Condition 3, which also allows to control $\mathcal{J}_{p}^{(s)}$.
-However, under Condition 3, this bound might be growing (sub-linearly) with $s$.
-We summarize the relations between these conditions in Figure 1.
-Bound on ${\text{Bound on }\mathcal{J}_{p}^{(s)}\text{ uniform over }s}$ Condition 1Condition 2Condition 3Bound on ${\text{Bound on }\mathcal{J}_{p}^{(s)}\text{ sub-linear in }s}$Corollary 4.1Corollary 4.1Lemma 4.2Lemma 4.3Lemma 4.1
-Figure 1: Relation between the various ergodic conditions in Subsection 4.1. In this diagram, $\Rightarrow$ is to be read as ”implies under the assumptions of the result referred to”.
-Condition 1 ($V$-geometric ergodicity (Douc et al., 2018, Section 18)).
-Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator.
-We say that $P$ is $V$-geometrically ergodic for some function $V\colon\mathcal{X}\to[1,\infty)$, when there exist constants $\theta\in(0,1)$ and $M\in\mathbb{R}_{+}$ such that for any $x\in\mathcal{X}$,
-$\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s},$
-where for any $\nu\in\mathbb{R}^{\mathcal{X}}$, we wrote
-$\left\|\nu\right\|_{V}\doteq\sup_{x\in\mathcal{X}}V(x)\nu(x).$
-Condition 2 (Existence of a dominating distribution).
-Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator.
-Suppose that there exists a universal constant $C\in\mathbb{R}_{+}$ and a probability measure $\nu\in\mathcal{P}(\mathcal{X})$, such that for all $x^{\prime}\in\mathcal{X}$,
-$\sup_{x\in\mathcal{X}}P(x,x^{\prime})\leq C\nu(x^{\prime}).$
-Condition 3 (Weak $V$-geometric ergodicity).
-Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. We say that $P$ is weakly $V$-geometrically ergodic when there exists functions
-$V\colon\mathcal{X}\to[1,\infty),\qquad D\colon\mathcal{X}\to[1,\infty),$
-and a constant $b\geq 0$ such that for any $s\in\mathbb{N}$,
-$\sup_{x^{\prime}\in\mathcal{X}}V(x^{\prime})(P^{s}(x,x^{\prime})-\pi(x^{\prime}))\leq D(x)+b(s-1).$
-Lemma 4.1.
-Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary weakly $V$-geometrically ergodic transition operator, with $V,D,b$ defined as in Condition 3, and let $p\geq 1$ such that
-$\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty.$
-Then,
-$\mathcal{J}_{p}^{(s)}\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right).$
-Proof.
-Directly from the assumptions,
-$\mathcal{J}_{p}^{(s)}$ $\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left(\pi(x)\left[\pi(x^{\prime})+\frac{1}{V(x^{\prime})}\left(D(x)+b(s-1)\right)\right]\right)^{1-1/p}$
-$\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\sum_{x,x^{\prime}\in\mathcal{X}}\left[\pi(x)\frac{1}{V(x^{\prime})}\left(D(x)+b(s-1)\right)\right]^{1-1/p}$
-$\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right),$
-where the two inequalities stem from the sub-additivity of the function $x\mapsto x^{1-1/p}$. ∎
-Corollary 4.1 (Under $V$-geometric ergodicity).
-Let $P$ a $\pi$-stationary $V$-geometrically ergodic transition operator, with $V,M,\theta$ defined as in Condition 1. Then Condition 3 is satisfied with
-$D=M\theta V\text{ and }b=0.$
-If the following summability conditions are met for some $p\geq 1$,
-$\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\pi\circ V\right\|_{1-1/p}^{1-1/p}<\infty,$
-we obtain a control of $\mathcal{J}_{p}^{(s)}$ uniformly over $s$:
-$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\\ \leq\|\pi\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left\|\pi\circ D\right\|^{1-1/p}_{1-1/p}.$
-Proof.
-We first check Condition 3. For any $x,x^{\prime}\in\mathcal{X}$,
-$V(x)(P^{s}(x,x^{\prime})-\pi(x^{\prime}))$ $\leq\sup_{x^{\prime\prime}\in\mathcal{X}}V(x^{\prime\prime})|P^{s}(x,x^{\prime\prime})-\pi(x^{\prime\prime})|$
-$=\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s}\leq MV(x)\theta,$
-where the second inequality stems from the definition of $V$-geometric ergodicity.
-Thus, Condition 3 is satisfied with $D=M\theta V$ and $b=0$.
-An application of Lemma 4.1 gives the the second claim.
-∎
-Remark 4.1.
-Under the assumption $D\propto V$, we can clarify the summability assumptions in Corollary 4.1.
-  • Let us first assume that $\pi(x)\propto x^{-\alpha}$ for $\alpha\in(1,\infty)$.
-Then, for $V(x)\propto D(x)\propto x^{\beta}$, with $\beta\in\mathbb{R}_{+}$, $\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty$ holds only when $\beta>p/(p-1)$, and $\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty$ holds only when $\alpha-\beta>p/(p-1)$, thus we must have $p/(p-1)<\beta<\alpha-p/(p-1)$, which will only be possible for $\alpha>2p/(p-1)$.
-  • Let us now assume now that $\pi(x)\propto a^{-x}$ for $a\in(0,1)$.
-Then, for $V(x)\propto D(x)\propto x^{\beta}$, $\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty$ always holds but we still require $\beta>p/(p-1)$ in order for $\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty$ to hold.
-Lemma 4.2 (Under the existence of a dominating distribution).
-Assume Condition 2 is satisfied. If moreover, for $p\geq 1$,
-$\left\|\pi\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\nu\right\|_{1-1/p}^{1-1/p}<\infty,$
-then,
-$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq C^{1-1/p}\left\|\nu\right\|_{1-1/p}^{1-1/p}\left\|\pi\right\|_{1-1/p}^{1-1/p}.$
-Proof.
-It directly follows from the assumptions that for any $x,x^{\prime}\in\mathcal{X}$, and for any $s\in\mathbb{N}$,
-$\begin{split}Q^{(s)}(x,x^{\prime})&=\pi(x)\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s-1}(x,x^{\prime\prime})P(x^{\prime\prime},x^{\prime})\\ &\leq C\pi(x)\nu(x^{\prime})\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s-1}(x,x^{\prime\prime})=C\pi(x)\nu(x^{\prime}).\end{split}$
-As a result,
-$\sum_{x,x^{\prime}\in\mathcal{X}}Q^{(s)}(x,x^{\prime})^{1-1/p}\leq C^{1-1/p}\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}\nu(x^{\prime})^{1-1/p}.$
-In practice, Condition 3 can be readily verified through the following lemma.
-Lemma 4.3.
-Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator.
-Assume there exists functions $V\colon\mathcal{X}\to[1,\infty)$, $D\colon\mathcal{X}\to[1,\infty)$ and a constant $b\geq 0$ such that
-$\sup_{x^{\prime}\in\mathcal{X}}V(x^{\prime})(P(x,x^{\prime})-\pi(x^{\prime}))\leq D(x)$
-and, for all $x\in\mathcal{X}$, $PD(x)\leq D(x)+b$. Then, Condition 3 holds.
-In Section 6, we use Lemma 4.3 to check Condition 3 on a whole family of transition operators $(P_{T})_{T\in\mathcal{T}}$.
-Note that Lemma 4.3 also shows that Condition 2 immediately implies Condition 3 with $D=C$, $V=1$ and $b=0$, which completes Figure 1.
-Proof.
-By recursion, for any $t\in\mathbb{N}$,
-$\forall x\in\mathcal{X},\,P^{t}D(x)\leq D(x)+bt.$ (9)
-Indeed, (9) holds for $t=1$ by assumption. Assume it holds for $t-1$, then
-$P^{t}D(x)$ $=P[P^{t-1}D](x)$
-$=\sum_{x^{\prime}\in\mathcal{X}}P(x,x^{\prime})[P^{t-1}D](x^{\prime})$
-$\leq\sum_{x^{\prime}\in\mathcal{X}}P(x,x^{\prime})[D(x^{\prime})+b(t-1)]$
-$=PD(x)+b(t-1)$
-$\leq D(x)+bt.$
-Thus, for any $x^{\prime\prime}\in\mathcal{X}$,
-$V(x^{\prime\prime})[P^{t}(x,x^{\prime\prime})-\pi(x^{\prime\prime})]$ $=\sum_{x^{\prime}\in\mathcal{X}}P^{t-1}(x,x^{\prime})[P(x^{\prime},x^{\prime\prime})-\pi(x^{\prime\prime})]V(x^{\prime\prime})$
-$\leq\sum_{x^{\prime}\in\mathcal{X}}P^{t-1}(x,x^{\prime})D(x^{\prime})$
-$=P^{t-1}D(x)$
-$\leq D(x)+b(t-1),$
-where the last inequality follows from (9). ∎
-4.2   Uniformly bounded conditional entropy
-When for each state $x\in\mathcal{X}$, the conditional distribution $e_{x}P$ has bounded entropy, it is possible to guarantee finiteness of $\mathcal{J}_{p,\xi}$.
-Lemma 4.4.
-Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Suppose that for $p\geq 1$,
-$\mathcal{I}_{p}\doteq\sup_{x\in\mathcal{X}}\left\{\left\|e_{x}P\right\|^{1-1/p}_{1-1/p}\right\}<\infty.$
-Then, for any $s\in\mathbb{N}$, it holds that
-$\mathcal{J}_{p}^{(s)}\leq\mathcal{I}_{p}^{s}\left\|\pi\right\|_{1-1/p}^{1-1/p}.$
-Proof.
-Suppose that the claim holds for $s\geq 1$. Then,
-$\begin{split}\mathcal{J}_{p}^{(s+1)}&=\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)^{1-1/p}\left(\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s}(x,x^{\prime\prime})P(x^{\prime\prime},x^{\prime})\right)^{1-1/p}\\ &\leq\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s}(x,x^{\prime\prime})^{1-1/p}P(x^{\prime\prime},x^{\prime})^{1-1/p}\\ &=\sum_{x,x^{\prime\prime}\in\mathcal{X}}\pi(x)^{1-1/p}P^{s}(x,x^{\prime\prime})^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}P(x^{\prime\prime},x^{\prime})^{1-1/p}\\ &\leq\sum_{x,x^{\prime\prime}\in\mathcal{X}}\pi(x)^{1-1/p}P^{s}(x,x^{\prime\prime})^{1-1/p}\sup_{x\in\mathcal{X}}\left\|e_{x}P\right\|_{1-1/p}^{1-1/p}\\ &\leq\mathcal{J}_{p}^{(s)}\mathcal{I}_{p}.\end{split}$
-The case $s=1$ is easy to verify, thus the lemma holds by induction. ∎
-While a direct consequence of Lemma 4.4 is that
-$\mathcal{J}_{p,\xi}\leq\mathcal{I}_{p}^{t^{\sharp}_{\mathsf{mix}}(\xi)}\left\|\pi\right\|_{1-1/p}^{1-1/p},$
-guaranteeing finiteness of $\mathcal{J}_{p,\xi}$ for any $\xi\in(0,1)$, the above approach is only practical when bounding $\mathcal{J}_{p}^{(s)}$ for restrictively small values of $s$.
-4.3   Infinite graphs with controlled growth
-Natural structural assumptions on the underlying connection graph of a Markov chain also enable us to control the entropic term.
-Recall that an ergodic transition operator $P\in\mathcal{W}(\mathcal{X})$ induces a fully connected digraph $(\mathcal{X},\mathcal{E})$, where
-$\mathcal{E}\doteq\left\{(x,x^{\prime})\in\mathcal{X}^{2}\colon P(x,x^{\prime})>0\right\}$
-is the set of directed edges.
-The growth of a graph (Trofimov, 1985; Campbell et al., 2023) is defined as a function $g\colon\mathbb{N}\to\mathbb{N}\cup\left\{\infty\right\}$, where for $s\in\mathbb{N}$,
-$g(s)\doteq\sup_{\mathcal{S}\subset\mathcal{X},\rho\left((\mathcal{X},\mathcal{E})|_{\mathcal{S}}\right)\leq s}\left|\mathcal{S}\right|,$
-where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the diameter of a directed graph, which is the maximum eccentricity of any vertex in the graph [^6] With the convention that a disconnected graph has infinite diameter., $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.
-Condition 4 (Connection graph with polynomial growth).
-A graph is said to have at most polynomial growth when there exists $G\in\mathbb{R}_{+}$ and $q\geq 1$ such that for any $s\in\mathbb{N}$, it holds that $g(s)\leq Gs^{q}$.
-Then $q$ and $G$ are respectively called the degree and factor of growth.
-Lemma 4.5.
-Let $p,q\in[1,\infty),\xi\in(0,1)$ and let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator.
-If the connection graph of $P$ has polynomial growth of degree $q$ with factor $G$, it holds that
-$\mathcal{J}_{p,\xi}\leq G^{1/p}t^{\sharp}_{\mathsf{mix}}(\xi)^{q/p}\left\|\pi\right\|^{1-1/p}_{1-1/p}.$
-Proof.
-For $x\in\mathcal{X}$, let us denote $\mathcal{S}_{x}^{(s)}\subset\mathcal{X}$ the set of states accessible from $x$ in $s$ steps. We have that
-$\begin{split}\mathcal{J}_{p}^{(s)}&=\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}P^{s}(x,x^{\prime})^{1-1/p}\leq\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{S}_{x}^{(s)}}\left(\frac{1}{\mathcal{S}_{x}^{(s)}}\right)^{1-1/p}\\ &=\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}{\mathcal{S}_{x}^{(s)}}^{1/p}\leq\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}(s^{q}G)^{1/p},\\ \end{split}$
-where the first inequality stems from the uniform distribution maximizing the $\ell_{1-1/p}$ quasi-norm, and the second inequality follows from the polynomial growth assumption. ∎
-Observe that
-$\lim_{p\to\infty}\left\|\pi\right\|_{1-1/p}^{1-1/p}=\lim_{p\to\infty}\left[t^{\sharp}_{\mathsf{mix}}(\xi)\right]^{q/p}=1,$
-that is, for large values of $p$, the upper bound of Lemma 4.5 still approaches unity.
-Example 4.1 (Linear growth and stationary distribution with polynomially decaying tail).
-We let $\mathcal{X}=\mathbb{N}$, and suppose that there exists $C\in\mathbb{R}_{+}$, and $\eta\in(1/8,1/4)$ such that $\pi(x)\leq\frac{C}{x^{1+\eta}}$ for any $x\in\mathcal{X}$.
-Taking $p=2(1+1/\eta)$, it holds that
-$\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\leq C^{1-1/p}\sum_{x\in\mathcal{X}}\frac{1}{x^{(1+\eta)(1-1/p)}}=C^{1-1/p}\sum_{x\in\mathcal{X}}\frac{1}{x^{1+\eta/2}}=C^{2\frac{\eta+2}{\eta+1}}\zeta(1+\eta/2),$
-where $\zeta$ is the Riemann Zeta function.
-When the connection graph of $P$ enjoys linear growth, that is $q=1$, an application of Theorem 3.4, yields that for $\delta\in(0,1/2)$, a trajectory length of
-$n\geq\widetilde{\mathcal{O}}\left(\frac{t_{\mathsf{mix}}^{11/10}}{\delta^{2}}\right),$
-is sufficient to obtain $\widehat{t_{\mathsf{mix}}}^{\sharp}(2\delta/5)\geq t^{\sharp}_{\mathsf{mix}}(\delta/2)$ with probability at least $1-\delta$.
-4.4   General finite space
-When the state space $\mathcal{X}$ is finite, the graph-related methods discussed in the previous section remain applicable. Alternatively, a more direct approach yields the following bound,
-$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left(\frac{1}{\left|\mathcal{X}\right|^{2}}\right)^{1-1/p}=\left|\mathcal{X}\right|^{2/p},$ (10)
-agnostic of the mixing properties of $P$. Note that (10) matches the heuristic lower bound $\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}$ in the worst-case where $P$ is bistochastic.
-Corollary 4.2 (of Theorem 3.4 for $\left|\mathcal{X}\right|<\infty$).
-Let $\mathcal{X}$ be a finite space.
-Let $\varepsilon\in(0,1/4)$, and $\delta,\xi\in(0,1)$.
-Let $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (6).
-For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$.
-When
-$n\geq 1+\frac{2^{8}et_{\mathsf{mix}}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+2\log_{2}\left|\mathcal{X}\right|\right),9\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+1\right)\log\left(\frac{4t_{\mathsf{mix}}}{\delta}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil\right)\right\},$
-with probability at least $1-\delta$, it holds that
-$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
-Proof.
-We plug (10) in Theorem 3.4 and optimize for $p$. The function $p\mapsto p\left|\mathcal{X}\right|^{2/p}$ is minimized for $p=2\log\left|\mathcal{X}\right|$, whence the corollary. ∎
-4.5   The two-point space
-In this section, we analyze the case where $\left|\mathcal{X}\right|=2$, illustrating the possibly stark difference between the worst-case mixing time and the average-mixing time.
-For $p,q\in(0,1)$, we let
-$P_{p,q}=\begin{pmatrix}1-p&p\\ q&1-q\end{pmatrix}.$
-Quantities such as $d$ or $d^{\sharp}$ subscripted with $p,q$ will be understood as pertaining to $P_{p,q}$. The matrix $P_{p,q}$ is irreducible, and its stationary distribution is readily given by
-$\pi_{p,q}^{\intercal}=\frac{1}{p+q}\begin{pmatrix}q\\ p\end{pmatrix}.$
-It follows from a direct computation (see e.g. the proof of Wolfer (2024, Lemma 8.1)) that
-$\begin{split}P_{p,q}^{t}&=1^{\intercal}\pi_{p,q}+\frac{(1-p-q)^{t}}{p+q}\begin{pmatrix}p&-p\\ -q&q\end{pmatrix},\\ \end{split}$
-and as a result,
-$\begin{split}d_{p,q}^{\sharp}(t)&=2\frac{pq}{(p+q)^{2}}\left|1-p-q\right|^{t},\qquad d_{p,q}(t)=\frac{p\lor q}{p+q}\left|1-p-q\right|^{t}.\end{split}$
-When $p=q$, the stationary distribution is uniform, $d_{p,q}^{\sharp}(t)=d_{p,q}(t)$, and
-$t_{\mathsf{mix}}(\xi)=t_{\mathsf{mix}}^{\sharp}(\xi).$
-However, for $\eta\doteq p/q<1$, the ratio of the distances is a function of the parameter $\eta$,
-$\frac{d_{p,q}(t)}{d^{\sharp}_{p,q}(t)}=\frac{(1+\eta)^{2}}{2\eta}\doteq r_{\eta}(t),$
-whence the following lemma holds.
-Lemma 4.6.
-There exists a family of Markov chains
-$\left\{P_{p,q}\colon(p,q)\in(0,1)^{2}\right\}\in\mathcal{W}^{\star}(\left\{0,1\right\})$
-such that for any $\xi\in(0,1)$ and any $M>0$, there exist $p,q\in(0,1)$ such that
-$t_{\mathsf{mix}}(\xi)>Mt_{\mathsf{mix}}^{\sharp}(\xi).$
-5   Proofs
+which depends solely on the stationary properties of the chain.
+Appendix C Proofs
 In this section, we compile technical proofs that were either outlined briefly or omitted in the manuscript. The subsequent technical tool will prove to be convenient.
-Lemma 5.1.
+Lemma C.1.
 Let $A,B\in\mathbb{R}_{+}$, with $A,B\geq e$.
@@
 Proof.
-The function $x\mapsto x/\log x$ is increasing on $[e,\infty)$ and in this range, it holds that $x/\log x\geq e>2$. Hence, for $B\geq 2A\log A$, it holds that
+The function $x\mapsto x/\log x$ is increasing on $e,\infty)$ and in this range, it holds that $x/\log x\geq e>2$. Hence, for $B\geq 2A\log A$, it holds that
 $\frac{B}{\log B}\geq\frac{2A\log A}{\log A+\log(2\log A)}\geq\frac{2A\log A}{\log A+\log A}=A,$
 whence the lemma holds. ∎
-5.1   Proof of Lemma 2.3
-The proof is standard, and based on the blocking technique credited to Bernstein and further developed in the series of papers (Yu, 1993, 1994; Eberlein, 1984; Volkonskii and Rozanov, 1959).
+C.1  Proof of Lemma 2.1
+The proof is standard, and based on the blocking technique credited to Bernstein and further developed in the series of papers (Yu, 1994; Eberlein, 1984; Volkonskii and Rozanov, 1959).
 For simplicity, we assume that $n=2Bs$, where $B,s\in\mathbb{N}$.
 For $b\in[B]$, we write
-$\begin{split}X^{[2b]}&=X^{[2b]}_{1},\dots,X^{[2b]}_{s}=X_{(2b-1)s+1},\dots,X_{2bs},\\ X^{[2b-1]}&=X^{[2b-1]}_{1},\dots,X^{[2b-1]}_{s}=X_{(2b-2)s+1},\dots,X_{(2b-1)s},\\ \end{split}$
+[…]
 and consider the partitioned process
@@
 […]
-Figure 2: Decomposing the process into blocks.
+Figure 4: Decomposing the process into Bernstein blocks.
 Invoking Hoeffding’s inequality (Hoeffding, 1963) and Yu (1994, Corollary 2.7), it holds that
@@
 $s=\left\lceil\left(\frac{1}{b\beta_{1}}W_{0}\left(b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b}\right)\right)^{1/b}\right\rceil,$
-it holds that $22(B-1)\beta(s)\leq\delta/$ 2. As a consequence, the deviation probability at (11) is bounded by $\delta$ when the trajectory length $n$ satisfies the inequality,
+it holds that $2(B-1)\beta(s)\leq\delta/2$. As a consequence, the deviation probability at (14) is bounded by $\delta$ when the trajectory length $n$ satisfies the inequality,
 $n\geq\frac{4}{\varepsilon^{2}}\left\lceil\left(\frac{1}{b\beta_{1}}W_{0}\left(b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b}\right)\right)^{1/b}\right\rceil\log\left(\frac{4}{\delta}\right).$
@@
 $A=\left[\frac{16\beta_{0}}{\delta\varepsilon^{2}}\log\left(\frac{4}{\delta}\right)\right]^{b},\qquad B=b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b},$
-and invoking Lemma 5.1, it is enough that
+and invoking Lemma C.1, it is enough that
 $n\geq\frac{8}{\varepsilon^{2}}\log\left(\frac{4}{\delta}\right)2^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right),$
@@
 ∎
-5.2   Proof of Lemma 2.1
+C.2  Proof of Lemma 2.2
 The proof closely follows the techniques in Levin et al.
@@
 We can write,
-$\begin{split}d^{\sharp}(t)&\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|\delta_{x}P^{t}-\pi\right\|_{\mathsf{TV}}\\ &=\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\left|P^{t}(x,x^{\prime})-\pi(x^{\prime})\right|\\ &=\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\pi(x^{\prime})\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\\ &\stackrel{{\scriptstyle(\spadesuit)}}{{\leq}}\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\pi(x^{\prime})\frac{(1-\gamma_{\star})^{t}}{\sqrt{\pi(x)\pi(x^{\prime})}}\\ &=\frac{1}{2}(1-\gamma_{\star})^{t}\sum_{x,x^{\prime}\in\mathcal{X}}{\sqrt{\pi(x)\pi(x^{\prime})}}\\ &=\frac{1}{2}(1-\gamma_{\star})^{t}\left\|\pi\right\|_{1/2}.\\ \end{split}$
+[…]
 For the inequality ($\spadesuit$), we used compactness of $P$ and the fact Levin et al.
@@
 thus as in Levin et al. (2009, Theorem 12.4),
-$\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\leq\left|\lambda_{\star}\right|^{t}\sqrt{\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x)\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x^{\prime})}.$
+$\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\leq\left|1-\gamma_{\star}\right|^{t}\sqrt{\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x)\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x^{\prime})}.$
 Solving for a proximity parameter $\xi\in(0,1)$, we obtain that
 $t^{\sharp}_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{2\xi}.$
-5.3   Proof of Theorem 3.1
-For $s\in\mathbb{N}$,
+C.3  Proof of Theorem 3.1
+The proof of Theorem 3.1 will rely on the following lemma.
+Lemma C.2.
+Let $X_{1},\dots,X_{n}$ be a Markov chain over a countable state space $\mathcal{X}$ with stationary distribution $\pi$. For any $x\in\mathcal{X}$, and any $p\in[1,\infty)$, it holds that
+$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}\right]&\leq 4(n-1)\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{n-2}\beta(t)^{1/p}\right)\leq 4(n-1)\pi(x)^{1-1/p}\mathcal{B}_{p},\end{split}$
+where $N_{x}=\sum_{t=1}^{n-1}\delta\left[X_{t}=x\right]$ counts the number of visits to state $x\in\mathcal{X}$.
+Proof.
+The sequence $(\delta\left[X_{t}=x\right])_{t\in\mathbb{N}}$ being stationary at second order, we expand the variance of the sum (Rio, 1999, (1.6)) as,
+$\operatorname{Var}_{\pi}\left[N_{x}\right]=(n-1)\operatorname{Var}_{\pi}\left[\delta\left[X_{1}=x\right]\right]+2\sum_{t=1}^{n-2}(n-1-t)\operatorname{Cov}_{\pi}(\delta\left[X_{1}=x\right],\delta\left[X_{1+t}=x\right]).$ (15)
+From Rio (1993, Theorem 1) [^3] This generalization of Ibragimov’s covariance inequality (Ibragimov, 1962) also appears in Davydov (1968) but with a weaker constant.—see also Rio (1999, (1.12b)) and Doukhan (2012, Theorem 3)— for a Hölder triplet $p,r,q\geq 1,\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$ and any $t\in\mathbb{N}\cup\left\{0\right\}$,
+$\begin{split}\operatorname{Cov}_{\pi}(\delta\left[X_{1}=x\right],\delta\left[X_{1+t}=x\right])&\leq 2(2\alpha(t))^{1/p}\left\|\delta\left[X_{1}=x\right]\right\|_{\pi,r}\left\|\delta\left[X_{t}=x\right]\right\|_{\pi,q}\\ &\leq 2\beta(t)^{1/p}\pi(x)^{1-1/p},\end{split}$
+where $\alpha$ is Rosenblatt’s strong mixing coefficient [^4] We note that $\alpha$ is defined as twice Rosenblatt’s coefficient in Rio (1999).
+(Rosenblatt, 1956), we wrote $\left\|X\right\|_{\pi,r}=\left(\mathbb{E}_{\pi}\left|X\right|^{r}\right)^{1/r}$ for a real random variable $X$, and for the second inequality, we relied on the inequality $2\alpha(t)\leq\beta(t)$ (Bradley, 2005, (1.11)).
+Plugging in (15), it follows that
+$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}\right]&\leq 2(n-1)\beta(0)^{1/p}\pi(x)^{1-1/p}+4\sum_{t=1}^{n-2}(n-1-t)\beta(t)^{1/p}\pi(x)^{1-1/p}\\ &\leq 4(n-1)\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{n-2}\beta(t)^{1/p}\right).\end{split}$
+Moving on to the proof of Theorem 3.1, for $s\in\mathbb{N}$, by definition of $\widehat{\beta}(s)$ in (8),
 $\begin{split}2\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|=&\mathbb{E}_{\pi}\left|\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|-\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{s}-\pi\right\|_{1}\right|\\ \leq&\mathbb{E}_{\pi}\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N_{xx^{\prime}}^{(s)}}{\lfloor(n-1)/s\rfloor}-\pi(x)P^{s}(x,x^{\prime})\right|\\ &+\mathbb{E}_{\pi}\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|.\end{split}$
-We further decompose the argument in the second expectation as
-$\begin{split}&\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|\\ &\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\frac{N^{(s)}_{x}\pi(x^{\prime})}{\lfloor(n-1)/s\rfloor}\right|+\left|\frac{N^{(s)}_{x}\pi(x^{\prime})}{\lfloor(n-1)/s\rfloor}-\pi(x)\pi(x^{\prime})\right|\\ &=2\sum_{x\in\mathcal{X}}\left|\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}-\pi(x)\right|.\\ \end{split}$
-Noting that by stationarity,
+With a symmetry argument, we further decompose the argument in the second expectation as
+$\begin{split}\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|&\leq 2\sum_{x\in\mathcal{X}}\left|\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}-\pi(x)\right|.\end{split}$
+By stationarity,
 $\mathbb{E}_{\pi}\left[\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}\right]=\pi(x),\qquad\mathbb{E}_{\pi}\left[\frac{N_{xx^{\prime}}^{(s)}}{\lfloor(n-1)/s\rfloor}\right]=\pi(x)P^{s}(x,x^{\prime}),$
-it follows from Jensen’s inequality that
+and it follows from Jensen’s inequality that
 $\begin{split}2\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq\frac{1}{\lfloor(n-1)/s\rfloor}\left(\sum_{x,x^{\prime}\in\mathcal{X}}\sqrt{\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]}+2\sum_{x\in\mathcal{X}}\sqrt{\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]}\right).\end{split}$ (16)
 We proceed to bound the two variance terms.
-We first treat $\operatorname{Var}_{\pi}\left[N_{x}\right]$ for a fixed $x\in\mathcal{X}$.
-This will in turn yield bounds for general $\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]$ and $\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]$.
-The sequence $(\delta\left[X_{t}=x\right])_{t\in\mathbb{N}}$ being stationary at second order, we expand the variance of the sum (see e.g.
-Rio (1999, (1.6))) as,
-$\operatorname{Var}_{\pi}\left[N_{x}\right]=(n-1)\operatorname{Var}_{\pi}\left[\delta\left[X_{1}=x\right]\right]+2\sum_{t=1}^{n-2}(n-1-t)\operatorname{Cov}_{\pi}(\delta\left[X_{1}=x\right],\delta\left[X_{1+t}=x\right]).$ (13)
-From Rio (1993, Theorem 1) [^7] This generalization of Ibragimov’s covariance inequality (Ibragimov, 1962) also appears in Davydov (1968) but with a weaker constant.
-(see also Rio (1999, (1.12b)) and Doukhan (2012, Theorem 3)), for a Hölder triplet $p,r,q\geq 1,\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$ and any $t\in\mathbb{N}\cup\left\{0\right\}$,
-$\begin{split}\operatorname{Cov}_{\pi}(\delta\left[X_{1}=x\right],\delta\left[X_{1+t}=x\right])&\leq 2(2\alpha(t))^{1/p}\left\|\delta\left[X_{1}=x\right]\right\|_{\pi,r}\left\|\delta\left[X_{t}=x\right]\right\|_{\pi,q}\\ &\leq 2\beta(t)^{1/p}\pi(x)^{1-1/p},\end{split}$
-where $\alpha$ is Rosenblatt’s strong mixing coefficient [^8] We note that $\alpha$ is defined as twice Rosenblatt’s coefficient in Rio (1999).
-(Rosenblatt, 1956), we wrote $\left\|X\right\|_{\pi,r}=\left(\mathbb{E}_{\pi}\left|X\right|^{r}\right)^{1/r}$ for a real random variable $X$, and for the second inequality, we relied on the inequality $2\alpha(t)\leq\beta(t)$ (see e.g.
-(Bradley, 2005, (1.11))).
-Plugging in (13), it follows that
-$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}\right]&\leq 2(n-1)\beta(0)^{1/p}\pi(x)^{1-1/p}+4\sum_{t=1}^{n-2}(n-1-t)\beta(t)^{1/p}\pi(x)^{1-1/p}\\ &\leq 4(n-1)\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{n-2}\beta(t)^{1/p}\right),\end{split}$
-and for a general $s<n$, noting that the mixing coefficient $\beta^{(s)}(t)$ pertaining to the $s$-skipped chain satisfies $\beta^{(s)}(t)=\beta(st)$, we obtain
-$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]&\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p}\right)\\ &\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor\pi(x)^{1-1/p}\mathcal{B}^{(s)}_{p}.\\ \end{split}$
-We verify (see e.g. Wolfer and Kontorovich (2021, Lemma 6.1)) that the process consisting of a sliding window over pairs,
+From Lemma C.2, for a fixed $x\in\mathcal{X}$ and any $s<n$, noting that the mixing coefficient $\beta^{(s)}(t)$ pertaining to the $s$-skipped chain satisfies $\beta^{(s)}(t)=\beta(st)$, we obtain
+[…]
+The process consisting of a sliding window over pairs,
 $\{E_{t}\}_{t\in\mathbb{N}}\doteq\left\{\left(X_{t},X_{t+1}\right)\right\}_{t\in\mathbb{N}}$
-forms a Markov chain over the state space =$\mathcal{E}=\left\{x,x^{\prime}\in\mathcal{X}\colon P(x,x^{\prime})>0\right\}$, called the Hudson expansion (Kemeny and Snell, 1983; Wolfer and Watanabe, 2024) of the original chain, with transition operator
-$\overline{P}(e=(x_{1},x_{2}),e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime}))=\delta\left[x_{2}=x_{1}^{\prime}\right]P(x_{2},x_{2}^{\prime}),$
+forms a Markov chain (Wolfer and Kontorovich, 2021, Lemma 6.1) over the state space
+$\mathcal{E}=\left\{(x,x^{\prime})\in\mathcal{X}^{2}\colon P(x,x^{\prime})>0\right\},$
+called the Hudson expansion (Kemeny and Snell, 1983; Wolfer and Watanabe, 2024) of the original chain, with transition operator
+$\overline{P}(e=(x_{1},x_{2}),e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime}))=\delta\left[x_{2}=x_{1}^{\prime}\right]P(x_{2},x_{2}^{\prime}),\forall(e,e^{\prime})\in\mathcal{E}^{2},$
 and with stationary distribution $\overline{\pi}((x,x^{\prime}))=\pi(x)P(x,x^{\prime})$. Writing $\overline{\beta}(s)$ for the $\beta$-mixing coefficient of $\overline{P}$, we compute
-$\begin{split}\overline{\beta}(0)&=1-\left\|\overline{\pi}\right\|_{2}^{2}\\ \overline{\beta}(1)&=\beta(0)=1-\left\|\pi\right\|_{2}^{2},\end{split}$
-and it follows from the Markov property that for $t\in\mathbb{N},t<n-1$,
+$\begin{split}\overline{\beta}(0)&=1-\left\|\overline{\pi}\right\|_{2}^{2},\\ \overline{\beta}(1)&=\beta(0)=1-\left\|\pi\right\|_{2}^{2}.\end{split}$
+Moreover, it follows from the Markov property that for $t\in\mathbb{N}$,
 $\overline{P}^{t+1}((x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime}))=P(x_{1}^{\prime},x_{2}^{\prime})P^{t}(x_{2},x_{1}^{\prime}),$
-and thus, for $t\geq 1$,
+and thus,
 $\begin{split}\overline{\beta}(t+1)&=\frac{1}{2}\sum_{e=(x_{1},x_{2})\in\mathcal{E}}\overline{\pi}(e)\sum_{e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime})\in\mathcal{E}}\left|\overline{P}^{t+1}(e,e^{\prime})-\overline{\pi}(e^{\prime})\right|\\ &=\frac{1}{2}\sum_{(x_{1},x_{2})\in\mathcal{E}}\pi(x_{1})P(x_{1},x_{2})\sum_{(x^{\prime}_{1},x^{\prime}_{2})\in\mathcal{E}}\left|P^{t}(x_{2},x_{1}^{\prime})-\pi(x_{1}^{\prime})\right|P(x_{1}^{\prime},x_{2}^{\prime})\\ &=\frac{1}{2}\sum_{x_{2}\in\mathcal{X}}\pi(x_{2})\sum_{x^{\prime}_{1}\in\mathcal{X}}\left|P^{t}(x_{2},x_{1}^{\prime})-\pi(x_{1}^{\prime})\right|=\beta(t).\end{split}$
-More generally, for $s\in\mathbb{N},t\in\mathbb{N}\cup\left\{0\right\}$, $1\leq s,\leq n-1$, and $0\leq t<\lfloor(n-1)/s\rfloor$,
+More generally, for $s\in\mathbb{N},t\in\mathbb{N}\cup\left\{0\right\}$, $1\leq s\leq n-1$, and $0\leq t<\lfloor(n-1)/s\rfloor$,
 $\overline{\beta}^{(s)}(t+1)=\beta^{(s)}(t)=\beta(st).$
-As a consequence
-$\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor Q^{(s)}(x,x^{\prime})^{1-1/p}\left(\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\overline{\beta}^{(s)}(t)^{1/p}\right),$
-where we recall that $Q^{(s)}(x,x^{\prime})=\pi(x)^{s}P(x,x^{\prime})$. We further bound
-$\begin{split}\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\overline{\beta}^{(s)}(t)^{1/p}&=\frac{1}{2}\left(1-\left\|Q^{(s)}\right\|_{2}^{2}\right)+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\beta(s(t-1))^{1/p}\\ &\leq\frac{1}{2}\left(1-\left\|\pi\right\|_{2}^{2}\right)+\sum_{t=0}^{\lfloor(n-1)/s\rfloor-3}\beta(st))^{1/p}\\ &\leq\frac{3}{2}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-3}\beta(st)^{1/p}\leq\frac{3}{2}\mathcal{B}_{p}^{(s)},\end{split}$
+An application of Lemma C.2 on the Hudson expanded chain yields that
+$\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor Q^{(s)}(x,x^{\prime})^{1-1/p}\left(\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-1}\overline{\beta}^{(s)}(t)^{1/p}\right),$
+where we recall that $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$. We further bound
+$\begin{split}\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-1}\overline{\beta}^{(s)}(t)^{1/p}&=\frac{1}{2}\left(1-\left\|Q^{(s)}\right\|_{2}^{2}\right)+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-1}\beta(s(t-1))^{1/p}\\ &\leq 1/2+\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p}\leq 1/2+\mathcal{B}_{p}^{(s)},\end{split}$
 and plugging in (16),
-$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}}\left(\sqrt{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\frac{3}{2}\mathcal{B}_{p}^{(s)}}+2\sqrt{\left\|\pi\right\|_{1-1/p}^{1-1/p}\mathcal{B}_{p}^{(s)}}\right),\end{split}$
-which yields the first inequality in the theorem. The second inequality can be obtained from observing that $\left\|Q^{(s)}\right\|_{1-1/p}\geq\left\|\pi\right\|_{1-1/p}$. ∎
-5.4   Proof of Corollary 3.1
+$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}}\left(\left\|Q^{(s)}\right\|_{(1-1/p)/2}^{(1-1/p)/2}\sqrt{1/2+\mathcal{B}_{p}^{(s)}}+2\left\|\pi\right\|_{(1-1/p)/2}^{(1-1/p)/2}\sqrt{\mathcal{B}_{p}^{(s)}}\right).\end{split}$
+Finally, the claim can be obtained from observing that $\left\|Q^{(s)}\right\|_{(1-1/p)/2}\geq\left\|\pi\right\|_{(1-1/p)/2}$. ∎
+C.4  Proof of Lemma 3.1
 Let us first analyze the exponential case, by assuming that $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s)$.
-$\begin{split}\mathcal{B}_{p}^{(s)}=\beta_{0}^{1/p}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\exp\left(-\beta_{1}st/p\right)&\leq\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}.\end{split}$
+$\begin{split}\mathcal{B}_{p}^{(s)}=\beta_{0}^{1/p}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-1}\exp\left(-\beta_{1}st/p\right)&\leq\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}.\end{split}$
 We proceed to address the sub-exponential case. Let us assume that $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$ for $\beta_{0}>0,\beta_{1}>0$ and $b\in(0,1)$. Then,
+$\begin{split}\mathcal{B}_{p}^{(s)}&\leq\beta_{0}^{1/p}+\beta_{0}^{1/p}\sum_{t=1}^{\lfloor(n-1)/s\rfloor-1}\exp\left(-\beta_{1}(st)^{b}/p\right)\\ &\leq\beta_{0}^{1/p}+\beta_{0}^{1/p}\exp(-\beta_{1}s^{b}/p)+\beta_{0}^{1/p}\sum_{t=2}^{\infty}\exp\left(-\beta_{1}s^{b}t^{b}/p\right)\\ &\leq\beta_{0}^{1/p}+\beta_{0}^{1/p}\exp(-\beta_{1}s^{b}/p)+\beta_{0}^{1/p}\int_{1}^{\infty}\exp(-\beta_{1}s^{b}t^{b}/p)dt\\ &=\beta_{0}^{1/p}+\beta_{0}^{1/p}\exp(-\beta_{1}s^{b}/p)+\beta_{0}^{1/p}\left[\frac{-1}{b(\beta_{1}s^{b}/p)^{1/b}}\Gamma\left(\frac{1}{b},\beta_{1}s^{b}t^{b}/p\right)\right]_{1}^{\infty}.\end{split}$
+Finally, in the polynomial case, where for $s\geq 1$, $\beta(s)\leq\beta_{1}/s^{b}$,
 […]
-Finally, in the polynomial case, where for $s\geq 1$, $\beta(s)\leq\beta_{1}/s^{b}$,
-$\begin{split}\mathcal{B}_{p}^{(s)}&\leq\beta_{0}^{1/p}+\beta(s)^{1/p}\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\frac{1}{t^{b/p}}\leq\beta_{0}^{1/p}+\zeta(b/p)\beta(s)^{1/p}.\\ \end{split}$
-5.5   Proof of Theorem 3.2
-Recall that for a uniformly ergodic Markov chain,
-$\beta(s)\leq d(s)\leq d(\lfloor s/t_{\mathsf{mix}}\rfloor t_{\mathsf{mix}})\leq\exp(-\lfloor s/t_{\mathsf{mix}}\rfloor\log 2)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$
-where the third inequality is adapted from Levin et al.
-(2009, (4.33)), i.e.
-$\beta$ is decays exponentially with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$.
-Thus, from the fact that $x>0\implies 1/(1-\exp(-x))\leq 1+1/x,$ and Corollary 3.1, it holds that
-$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{2^{1/p}}{\lfloor(n-1)/s\rfloor}\left(1+\frac{pt_{\mathsf{mix}}}{s\log 2}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}},\\ \end{split}$
-and $(i)$ follows from $n>s+1$ and $x>1\implies\lfloor x\rfloor\geq x/2$ and $2^{1/p}\leq 2$. From $(i)$, when
-$n\geq 1+\frac{2^{8}}{\varepsilon^{2}}\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
+C.5  Proof of Theorem 3.2
+We first focus on the estimation of $\beta(1)$.
+The general result for $\beta(s)$ will be obtained by considering an $s$-skipped chain.
+First, similar to the proof of Theorem 3.1, we decompose using the triangle inequality,
+$\begin{split}2\left|\widehat{\beta}(1)-\beta(1)\right|&\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N_{xx^{\prime}}}{n-1}-Q(x,x^{\prime})\right|+2\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|.\end{split}$
+We treat the two sums separately,
+$\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(1)-\beta(1)\right|>\varepsilon\right)\leq\mathbb{P}_{\pi}\left(\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N_{xx^{\prime}}}{n-1}-Q(x,x^{\prime})\right|>\varepsilon\right)+\mathbb{P}_{\pi}\left(\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|>\varepsilon/2\right).$ (17)
+Let us focus on the second probability statement. For simplicity, we assume that n1=2Brn-1=2Br, where we decompose the process into $2B$ Bernstein blocks (refer to Figure 4) of size $r$.
+[…]
+Therefore, applying Yu (1994, Corollary 2.7),
+$\mathbb{P}_{\pi}\left(\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|>\varepsilon/2\right)\leq\sum_{\sigma\in\{0,1\}}\mathbb{P}_{\pi}\left(\frac{1}{B}\sum_{b=1}^{B}Z_{b}^{(\sigma)}>\varepsilon/2\right)+(B-1)\beta(r),$ (18)
+where
+$Z_{b}^{(\sigma)}\doteq\frac{1}{r}\sum_{x\in\mathcal{X}}\left|\sum_{t=1}^{r}\left(\delta\left[X_{t}^{[2b-\sigma]}=x\right]-\pi(x)\right)\right|$
+are independent random variables.
+First, observing that $Z_{b}^{(\sigma)}$ is a function of a Markov chain $X_{1},\dots,X_{r}$ stationarily sampled from $P$, we upper bound the expected value as in Theorem 3.1,
+$\begin{split}\mathbb{E}_{\pi}Z_{b}^{(\sigma)}&\stackrel{{\scriptstyle(i)}}{{\leq}}\frac{1}{r}\sum_{x\in\mathcal{X}}\sqrt{\operatorname{Var}_{\pi}\left[\sum_{t=1}^{r}\delta\left[X_{t}^{[2b-\sigma]}=x\right]\right]}\\ &\stackrel{{\scriptstyle(ii)}}{{\leq}}\frac{2}{\sqrt{r}}\left\|\pi\right\|_{(1-1/p)/2}^{(1-1/p)/2}\sqrt{\mathcal{B}_{p}},\end{split}$
+where $(i)$ is Jensen’s inequality and $(ii)$ is an application of Lemma C.2. Thus for
+$r\geq\frac{64\left\|\pi\right\|_{(1-1/p)/2}^{1-1/p}\mathcal{B}_{p}}{\varepsilon^{2}},$
+it holds that
+$\mathbb{E}_{\pi}Z_{b}^{(\sigma)}\leq\varepsilon/4,$
+and it remains to bound
+$\mathbb{P}_{\pi}\left(\frac{1}{B}\sum_{b=1}^{B}Z_{b}^{(\sigma)}-\mathbb{E}_{\pi}Z_{b}^{(\sigma)}>\varepsilon/4\right).$
+The random variables $Z_{b}^{(\sigma)}$ are bounded. Indeed,
+[…]
+Therefore, from an application of Hoeffding’s inequality,
+$\mathbb{P}_{\pi}\left(\frac{1}{B}\sum_{b=1}^{B}Z_{b}^{(\sigma)}-\mathbb{E}_{\pi}Z_{b}^{(\sigma)}>\varepsilon/4\right)\leq\exp\left(-\frac{2(B\varepsilon/4)^{2}}{B2^{2}}\right)=\exp\left(-\frac{B\varepsilon^{2}}{32}\right).$
+Plugging in (18),
+$\mathbb{P}_{\pi}\left(\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|>\varepsilon/2\right)\leq 2\exp\left(-\frac{B\varepsilon^{2}}{32}\right)+(B-1)\beta(r).$
+For $B\geq\frac{32\log(8/\delta)}{\varepsilon^{2}}$ and $r\geq t^{\sharp}_{\mathsf{mix}}(\xi(\varepsilon,\delta))$ with
+$\xi(\varepsilon,\delta)=\frac{\varepsilon^{2}\delta}{128\log(8/\delta)},$
+the above probability is smaller than $\delta/2$ In other words, regardless of uniform ergodicity, for
+$n\geq 1+\frac{64\log(8/\delta)}{\varepsilon^{2}}\max\left\{\frac{64}{\varepsilon^{2}}\left\|\pi\right\|_{(1-1/p)/2}^{1-1/p}\mathcal{B}_{p},t^{\sharp}_{\mathsf{mix}}(\xi(\varepsilon,\delta))\right\},\qquad\xi(\varepsilon,\delta)=\frac{\varepsilon^{2}\delta}{128\log(8/\delta)},$
+it holds that
+$\mathbb{P}_{\pi}\left(\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|>\varepsilon/2\right)\leq\delta/2.$
+A similar approach for the sliding-window chain $(X_{t},X_{t+1})_{t\in\mathbb{N}}$ leads to a bound on the first term of (17).
+Finally, considering the $s$-skipped version of a chain, and noting that for any $\xi\in(0,1)$, ${t^{\sharp}_{\mathsf{mix}}}^{(s)}(\xi)\leq\lceil t^{\sharp}_{\mathsf{mix}}(\xi)/s\rceil$ yields the claim for $\beta(s)$.
+∎
+C.6  Proof of Theorem 3.3
+Our starting point for this proof if (16).
+However, instead of Rio’s inequality, we rely on techniques in Paulin (2015) to obtain an upper bound on the two variance terms.
+The mixing time of the $s$-skipped pairs of observations is upper bounded by $t_{\mathsf{mix}}^{(s)}+1\leq\lceil t_{\mathsf{mix}}/s\rceil+1$ (Wolfer and Kontorovich, 2021, Lemma 6.1).
+Combining Paulin (2015, Theorem 3.2, Proposition  3.4), it follows that
+[…]
+which yields
+[…]
+The claim with $C_{\mathsf{u.erg.}}\leq 3\sqrt{2}$ follows from $n>s+1$ and $x>1\implies\lfloor x\rfloor\geq x/2$ and $s(\lceil t_{\mathsf{mix}}/s\rceil+1)\leq t_{\mathsf{mix}}+2s$.
+C.7  Proof of Theorem 3.4
+From Theorem 3.3, when
+$n\geq 1+\frac{4C_{\mathsf{u.erg.}}^{2}}{\varepsilon^{2}}(t_{\mathsf{mix}}+2s)\mathcal{J}_{\infty}^{(s)},$
 it holds that, $\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon/2$ and in this case, the probability of the above deviation can be upper bounded by fluctuations around the mean as follows,
-$\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)\leq\mathbb{P}_{\pi}\left(\left|\left|\widehat{\beta}(s)-\beta(s)\right|-\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\right|>\varepsilon/2\right).$
+[…]
 Let
-$x_{[s]}^{n}=x_{1},x_{1+s},x_{1+2s},\dots,x_{1+\lfloor(n-1)/s\rfloor s}\in\mathcal{X}$
-be some $s$-skipped trajectory, let $t\in\left\{0,\dots,\lfloor(n-1)/s\rfloor\right\}$, $x^{\prime}_{1+ts}\in\mathcal{X}$, and denote
-$x^{n,(t)}=x_{1},x_{1+s},x_{1+2s},\dots,x_{1+(t-1))s},x^{\prime}_{1+ts},x_{1+(t+1)s}\dots,x_{1+\lfloor(n-1)/s\rfloor s}$
-the trajectory where the observation $x_{1+ts}$ at time $1+ts$ was replaced with observation $x^{\prime}_{1+ts}$. We show that the estimator satisfies the bounded differences condition. We decompose,
-$\begin{split}&\left|\left|\widehat{\beta}(s)(X)-\beta(s)\right|-\left|\widehat{\beta}(s)(X^{(t)})-\beta(s)\right|\right|\\ &\leq\left|\widehat{\beta}(s)(X)-\widehat{\beta}(s)(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}(X)-N_{xx^{\prime}}^{(s)}(X^{(t)})\right|\\ &+\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|.\\ \end{split}$
+$\bar{x}=(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n},$
+and for $t\in[n]$ and $x^{\prime}_{t}\in\mathcal{X}$, let us denote
+$\bar{x}^{\left\{t\right\}}\doteq(x_{1},x_{2},\dots,x_{t-1},x^{\prime}_{t},x_{t+1}\dots,x_{n})$
+the trajectory where the observation $x_{t}$ at time $t$ was replaced with observation $x^{\prime}_{t}$.
+We show that the absolute deviation $\left|\widehat{\beta}(s)-\beta(s)\right|\colon\mathcal{X}^{n}\to\mathbb{R}$ satisfies the bounded differences condition.
+We decompose,
+[…]
 We further upper bound the second term as follows,
-$\begin{split}&\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &+\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X^{(t)})-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x^{\prime}\in\mathcal{X}}\left|N_{x^{\prime}}^{(s)}(X)-N_{x^{\prime}}^{(s)}(X^{(t)})\right|+\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x\in\mathcal{X}}\left|N_{x}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})\right|.\\ \end{split}$
+[…]
 We obtain
-$\left|\left|\widehat{\beta}(s)(X)-\beta(s)\right|-\left|\widehat{\beta}(s)(X^{(t)})-\beta(s)\right|\right|\leq\frac{4}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}.$
+$\left|\left|\widehat{\beta}(s)(\bar{x})-\beta(s)\right|-\left|\widehat{\beta}(s)(\bar{x}^{\left\{t\right\}})-\beta(s)\right|\right|\leq\frac{4}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}.$
 From a version of McDiarmid’s inequality for uniformly ergodic Markov chains (Paulin, 2015, Corollary 2.10), and the fact that the mixing time of the $s$-skipped chain $t_{\mathsf{mix}}^{(s)}$ verifies $t_{\mathsf{mix}}^{(s)}\leq\lceil t_{\mathsf{mix}}/s\rceil$,
 $\begin{split}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)&\leq 2\exp\left(-\frac{2(\varepsilon/2)^{2}}{9t_{\mathsf{mix}}^{(s)}\lfloor(n-1)/s\rfloor\left(\frac{8}{\lfloor(n-1)/s\rfloor}\right)^{2}}\right)\\ &\leq 2\exp\left(-\frac{(n-1)\varepsilon^{2}}{9\cdot 2^{8}(t_{\mathsf{mix}}+s)}\right),\end{split}$
-which completes the proof of claim $(ii)$. ∎
-5.6   Proof of Theorem 3.3
-Suppose that
-$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},18\log\left(\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\right)\right\}.$
-Invoking Lemma 5.1 with
-$a\doteq\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon),\qquad b\doteq\frac{2}{\delta_{n-1}},$
-and where $\delta_{n-1}\doteq\frac{\delta}{n-1}$, it holds that
-$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}.$ (14)
-Writing $S=S(n,\varepsilon,\delta,p,\mathcal{J}_{p})$, we thus have
-$\begin{split}n-1&\geq\frac{n-1}{2}+\frac{2^{8}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}\\ &\geq\frac{2^{8}}{\varepsilon^{2}}\left[S+\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\right]\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}\\ &\geq\frac{2^{8}}{\varepsilon^{2}}\max\left\{\left(S+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\mathcal{J}_{p},9(S+t_{\mathsf{mix}})\log\frac{2}{\delta_{n-1}}\right\}.\end{split}$
-Furthermore, observing from (14) that
-$S\geq\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\geq t^{\sharp}_{\mathsf{mix}}(\varepsilon),$
-it follows that $\beta(s)<\varepsilon$ for $s>S$. An application of the union bound and Theorem 3.2$(ii)$ yield that
-$\mathbb{P}_{\pi}\left(\left\|\widehat{\beta}-\beta\right\|_{\infty}>\varepsilon\right)\leq\sum_{s=1}^{S}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)\leq S\delta_{n-1}\leq\delta.$
-The claim then follows from $t_{\mathsf{mix}}(\varepsilon)\leq\lceil\log_{2}\varepsilon^{-1}\rceil t_{\mathsf{mix}}$ for any $\varepsilon\in(0,1/4)$ (Levin et al., 2009, (4.34)). ∎
-5.7   Proof of Theorem 3.4
-We first bound the probability that the estimator overshoots, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)>t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$.
+which completes the proof. ∎
+C.8  Proof of Lemma 3.2
+We first bound the probability that the estimator overshoots the target interval, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)>t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$.
 On this event, for any $s\leq t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, and in particular for $s=t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, it must be that $\widehat{\beta}(s)>\xi$.
-From Theorem 3.2$(ii)$, for
-$n\geq 1+\frac{2^{8}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{4}{\delta}\right\},$
+From our assumption
+$n\geq n_{0}(P,t^{\sharp}_{\mathsf{mix}}(\xi(1-\varepsilon)),\xi\varepsilon,\delta/2),$
 it holds that
@@
 As a second step, we bound the probability that $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ underestimates the true parameter, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)<t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon))$.
-From Theorem 3.2$(ii)$ and an application of the union bound, when for any $s\leq t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))$ the trajectory length satisfies
-$n\geq 1+\frac{2^{8}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{4t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))}{\delta}\right\},$
+From our assumption and an application of the union bound, when for any $s\leq t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))$ the trajectory length satisfies
+$n\geq n_{0}\left(P,s,\xi\varepsilon,\frac{\delta}{2t^{\sharp}_{\mathsf{mix}}(\xi)}\right),$
 it holds that
 $\begin{split}\mathbb{P}_{\pi}\left(\exists s<t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon)),\widehat{\beta}(s)\leq\xi\right)&\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\widehat{\beta}(s)\leq\xi\right)\\ &\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\beta(s)-\widehat{\beta}(s)\geq-\xi+\xi(1+\varepsilon)\right)\\ &\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|\geq\xi\varepsilon\right)\leq\frac{\delta}{2}.\end{split}$
-The theorem follows from $t^{\sharp}_{\mathsf{mix}}(\xi(1-\varepsilon))\leq t_{\mathsf{mix}}(\xi(1-\varepsilon))\leq t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\xi(1-\varepsilon)}\right\rceil\leq t_{\mathsf{mix}}\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil$.
-∎
-6   Example
-In this section, we identify $\mathcal{X}=\mathbb{N}$, and we study the properties of a family of transition operators $(P_{T})_{T\in\mathcal{T}}$.
-Definition 6.1.
-Let $q\in(0,1)$, a sequence $\nu=(\nu_{x})_{x=1,\dots}$ taking values in $[0,1]$, and a probability distribution $\mu=(\mu_{x})_{x=2,\dots}$. We define:
-$P_{T}=\left(\begin{array}[]{c c c c c}1-q&q\mu_{2}&q\mu_{3}&q\mu_{4}&\dots\\ \nu_{2}&1-\nu_{2}&0&0&\dots\\ \nu_{3}&0&1-\nu_{3}&0&\dots\\ \nu_{4}&0&0&1-\nu_{4}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{array}\right).$
-We define $\mathcal{T}$ as the set of triples $T=(q,\mu,\nu)$ such that $q>0$, $\nu_{x}>0$ and $\mu_{x}>0$ for all $x$ and $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<+\infty$.
-A first observation is that when $q>0$ and $\nu_{x}>0$ and $\mu_{x}>0$ for all $x$, then the chain is aperiodic and irreducible.
-Thus it has an invariant positive measure (not necessarily finite).
-Solving $\pi P_{T}=\pi$, we find that the invariant measure satisfies, for $x\geq 2$,
-$\pi(x)=\frac{q\mu_{x}}{\nu_{x}}\pi(1).$
-So, if $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}=+\infty$, the chain is recurrent null.
-But if $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<\infty$ then it is recurrent positive, and the unique invariant probability measure, that will be denoted by $\pi_{T}$, is given by
-$\pi_{T}(1)=\frac{1}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\text{ and }\pi_{T}(x)=\frac{q\mu_{x}}{\nu_{x}\left(1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}\right)}\text{ for }x>1.$ (15)
-Thus, when $T=(q,\mu,\nu)\in\mathcal{T}$ the chain is recurrent positive.
-At first, we will study the ergodic properties of $P_{T}$ and evaluate its $\beta$-mixing coefficients.
-Secondly, we will apply the results of Section 4 to control $\mathcal{J}_{p}^{(s)}$ for $P_{T}$.
-6.1   Ergodic and mixing properties of $P_{T}$
-We remind that the Dobrushin coefficient of $P_{T}$ is defined by
-$\kappa:=\sup_{x,y\in\mathbb{N}}\|P_{T}(x,\cdot)-P_{T}(y,\cdot)\|_{{\rm TV}}.$ (16)
-An application of Theorem 18.2.4 in Douc et al. (2018) gives, for any $t\in\mathbb{N}$,
-$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\leq\kappa^{t}.$ (17)
-Thus, when $\kappa<1$, the chain is uniformly ergodic.
-Note for example that if $q<1$ and $\nu_{x}\geq 1-q$, then for any $x$ we have $P_{T}(x,1)\geq 1-q$.
-Thus, for any $x$ and $y$, $\|P_{T}(x,\cdot)-P_{T}(y,\cdot)\|_{{\rm TV}}\leq q<1$, and thus $\kappa<q<1$.
-So the chain defined by $P_{T}$ is uniformly ergodic.
-In this case, $t_{\mathsf{mix}}(\xi)\leq\lceil\log(\xi)/\log q\rceil$ and the chain is exponentially $\beta$-mixing.
-A first interesting observation is that when $\nu_{x}\rightarrow 0$, the chain cannot be uniformly ergodic.
-Lemma 6.1.
-Assume $T=(q,\mu,\nu)\in\mathcal{T}$. If $\nu_{x}\rightarrow 0$ for $x\rightarrow\infty$, we have
-$d(t)=\sup_{x\geq 1}\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\geq\frac{1}{2}$
-and thus $t_{{\rm mix}}=+\infty$.
-Proof: we have, for a fixed $t$,
-$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\sum_{y=1}^{\infty}|P_{T}^{t}(x,y)-\pi_{T}(y)|$
-$\geq\frac{1}{2}|P_{T}^{t}(x,x)-\pi_{T}(x)|$
-$\geq\frac{1}{2}|(1-\nu_{x})^{t}-\pi_{T}(x)|.$
-Note that $\pi_{T}(x)\rightarrow 0$ for $x\rightarrow\infty$, while $(1-\nu_{x})^{t}\rightarrow 1$. $\square$
-Of course, even though it is not uniformly ergodic, as the chain remains ergodic for $T\in\mathcal{T}$.
-Thus, for $t\rightarrow+\infty$ we have $\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\rightarrow 0$.
-Thus, by the dominated convergence theorem, in the stationary case, we have $\beta(t)=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\rightarrow 0.$ We can actually provide more accurate information on $\beta(t)$.
-This requires a careful analysis that will be simplified by additional technical assumptions on $T$.
-Definition 6.2.
-We define $\mathcal{S}\subset\mathcal{T}$ as the set of triples $T=(q,\mu,\nu)\in\mathcal{T}$ such that $q\in(0,1/4]$, $(\nu_{x})_{x=1,\dots}$ is nonincreasing and takes values in $(0,1/2]$ and $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<+\infty$.
-Lemma 6.2.
-Assume $T=(q,\mu,\nu)\in\mathcal{S}$. For any $K$ large enough to ensure $\sum_{x>K}\pi_{T}(x)\leq 1/2$,
-$\beta(t)\leq 2(1-\nu_{K})^{t}+7\sum_{x>K}\pi_{T}(x)+2tq\sum_{x>K}\mu_{x}.$
-We can simplify a little the above by noting that
-$q\mu_{x}=\pi_{T}(x)\nu_{x}\left(1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}\right)\leq\frac{\pi_{T}(x)\nu_{x}}{\pi_{T}(1)}\leq\frac{\pi_{T}(x)}{2\pi_{T}(1)}$
-and thus
-$\beta(t)\leq 2(1-\nu_{K})^{t}+\left(7+\frac{t}{\pi_{T}(1)}\right)\sum_{x>K}\pi_{T}(x).$
-Before we provide the proof, let us discuss various possible regimes:
-  • if $\nu_{i}x\rightarrow 0$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ for some $C^{\prime},b>0$ then there is a $C>0$ such that for $x>1$, $\pi(x)=C\exp(-bx)$ and thus
-    $\beta(t)$ $\leq 2(1-\nu_{K})^{t}+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}$
-    $\leq 2\exp(-t\nu_{K})+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}.$
-    For example if $\nu_{x}=\min(1/2,1/\log\log(x))$ we can take $K=t$ to obtain:
-    $\beta(t)\leq 2\exp\left(-\frac{t}{\log\log(t)}\right)(1+o(1)).$
-  • if $\nu_{x}=\min(1/2,1/x^{a})$ for some $a>0$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ for some $C^{\prime}>0$ and $b>1$ then there is a $C>0$ such that for $x>1$, $\pi_{T}(x)=C\exp(-bx)$ and thus
-    $\beta(t)\leq 2\exp\left(-\frac{t}{K^{a}}\right)+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}$
-    which, if we choose $K=\lfloor t^{\frac{1}{a+1}}\rfloor$ we obtain
-    $\beta(t)\leq\frac{\frac{Ct}{\pi_{T}(1)}\exp(-bt^{\frac{1}{a+1}})}{1-\exp(-b)}(1+o(1)).$
-  • if $\nu_{x}=\min(1/2,1/x)$ and $\mu_{x}=C^{\prime}\nu_{x}/x^{b}$ for some $C^{\prime}>0$ and $b>1$ then there is a $C>0$ such that for $x>1$, $\pi_{T}(x)=C/x^{b+1}$ and thus there is a $C^{\prime\prime}$ with
-    $\beta(t)\leq 2\exp(-t/K)+C^{\prime\prime}\left(7+\frac{t}{\pi_{T}(1)}\right)\frac{1}{(K+1)^{b}}$
-    and the choice $K=\lfloor t/((b-1)\log t)\rfloor$ gives
-    $\beta(t)$ $\leq\left(\frac{2}{t^{d}}+\frac{C^{\prime\prime}(b-1)^{b}}{\pi_{T}(1)}\frac{\log^{b}(t)}{t^{b-1}}\right)(1+o(1))$
-    $=\frac{C^{\prime\prime}(b-1)^{b}}{\pi(1)}\frac{\log^{b}(t)}{t^{b-1}}(1+o(1)).$
-Of course, these estimates could be seen as pessimistic. However, we also have lower bounds for $\beta(t)$.
-Lemma 6.3.
-Assume $T=(q,\mu,\nu)\in\mathcal{S}$. Assume $\nu_{x}=\min(1/2,1/x^{a})$ for some $a>0$. Then, we have, for $t\geq 2$,
-$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[\frac{1}{4}-\pi_{T}(\lceil t^{1/a}\rceil)\right].$
-That is,
-$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{8}(1+o(1)).$
-This is important with respect to the previous examples:
-  • by taking $a=1$ thus $\nu_{x}=\min(1/2,1/x)$ and $\mu(x)=C^{\prime}\nu_{x}/x^{b}=C^{\prime}/x^{b+1}$ for some $b>1$ we have $\pi_{T}(x)=C/x^{b+1}$ and thus
-    $\frac{C}{8(t+1)^{b+1}}\left[1+o(1)\right]\leq\beta(t)\leq\frac{C^{\prime\prime}(b-1)^{b}}{\pi_{T}(1)}\frac{\log^{b}(t)}{t^{b-1}}(1+o(1)).$
-  • in the case $\nu_{x}=\min(1/2,1/x^{a})$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ and thus $\pi_{T}(x)=C\exp(-bx)$, this gives
-    $\frac{C\exp[-b(t^{\frac{1}{a}}+1)]}{8}\left[1+o(1)\right]\leq\beta(t)\leq\frac{\frac{Ct}{\pi_{T}(1)}\exp(-bt^{\frac{1}{a+1}})}{1-\exp(-b)}(1+o(1)).$
-We conclude this subsection by the proofs of Lemmas 6.2 and 6.3.
-Proof of Lemma 6.2.
-The proof relies on the following approximation $P_{T,K}$ of $P_{T}$:
-$P_{T,K}=\left(\begin{array}[]{c c c c c c}1-q-q\sum_{x>K+1}\mu_{x}&q\mu_{2}&\dots&q\mu_{K}&0&\dots\\ \nu_{2}&1-\nu_{2}&0&0&0&\dots\\ \vdots&\vdots&\ddots&\vdots&\vdots&\dots\\ \nu_{K}&0&&1-\nu_{K}&0&\dots\\ \nu_{K+1}&0&\dots&0&1-\nu_{K+1}&\dots\\ \vdots&\vdots&&\vdots&\vdots&\ddots\end{array}\right).$
-Let $\pi_{T,K}$ denote its invariant distribution:
-$\pi_{T,K}(1)=\frac{1}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\text{ and }\pi_{T,K}(x)=\frac{q\mu_{x}}{\nu_{x}\left(1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}\right)}\text{ for }1<x\leq K$
-and $\pi_{T,K}(x)=0$ if $x>K$. Then, note that
-$\beta(t)$ $=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
-$\leq\sum_{x=1}^{K}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}+\sum_{x>K}\pi_{T}(x).$
-Then, for $x\leq K$, we have:
-$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\\ \leq\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}+\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}+\|\pi_{T,K}(\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}.$ (18)
-Let us control each term in (18) separately. First,
-$2\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}$ $=\left|\frac{1}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}-\frac{1}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\right|$
-$\quad\quad\quad+\sum_{x=2}^{K}\left|\frac{\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}-\frac{\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\right|+\sum_{x>K}\pi_{T}(x)$
-$=\left|1-\frac{1+q\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\right|\sum_{x=1}^{K}\pi_{T}(x)+\sum_{x>K}\pi_{T}(x)$
-$\leq\left|1-\frac{1+q\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\right|+\sum_{x>K}\pi_{T}(x)$
-$=\frac{\sum_{x>K}\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}+\sum_{x>K}\pi_{T}(x)$
-$=\frac{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\frac{\sum_{x>K}\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}+\sum_{x>K}\pi_{T}(x)$
-$=\left(\frac{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}+1\right)\sum_{x>K}\pi_{T}(x)$
-$=\left(\frac{1}{1-\sum_{x>K}\pi_{T}(x)}+1\right)\sum_{x>K}\pi_{T}(x)$
-$\leq 3\sum_{x>K}\pi_{T}(x)$
-as soon as $K$ large enough to ensure $\sum_{x>K}\pi_{T}(x)\leq 1/2$.
-We then tackle the term $\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}$ in (18). In order to do so, we study the Dobrushin coefficient of $P_{T,K}$,
-$\kappa:=\sup_{x,y\leq K}\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$
-and show it satisfies $\kappa<1$. Exactly as we derived (17) from (16) above, an application of Theorem 18.2.4 in Douc et al. (2018) will allow to conclude. When $1<x<y$,
-$\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\left(|\nu_{x}-\nu_{y}|+|1-\nu_{x}|+|1-\nu_{y}|\right)$
-$=1-\nu_{y}.$
-In the case $1=x<y$, and there we need $q\leq 1/4$ and $\nu_{y}\leq 1/2$,
-$\|P_{T,K}(1,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\left(\left|1-q-q\sum_{x>K}\mu_{x}-\nu_{y}\right|+|1-\nu_{y}-q\mu_{y}|+\sum_{x\leq K,x\neq y}q\mu_{x}\right)$
-$=\frac{1}{2}\left[2-2\nu_{z}-q\left(\sum_{x=2}^{\infty}\mu_{x}+\sum_{x>K}\mu_{x}+\mu_{y}-\sum_{x\leq K,x\neq y}\mu_{x}\right)\right]$
-$\leq 1-\nu_{y}.$
-So, $\kappa=\sup_{x,y\leq K}\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}=1-\nu_{K}$ which leads to the bound
-$\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}\leq(1-\nu_{K})^{t}.$
-Finally, we control the term $\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}$ in (18).
-This can be done thanks to the theory of the stability of Markov chains, that studies how fast two Markov chains with very similar transition operators will diverge, see for example Mitrophanov (2005).
-More precisely, (3.17) in the proof of Corollary 3.2 in Mitrophanov (2005) shows that
-$\sup_{1\leq x\leq K}\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}\leq t\sup_{1\leq x\leq K}\|P_{T}(x,\cdot)-P_{T,K}(x,\cdot)\|_{{\rm TV}}.$
-But it is direct to check that
-$\sup_{1\leq x\leq K}\|P_{T}(x,\cdot)-P_{T,K}(x,\cdot)\|_{{\rm TV}}$ $=\|P_{T}(1,\cdot)-P_{T,K}(1,\cdot)\|_{{\rm TV}}$
-$=q\sum_{x>K}\mu_{x}.$
-Thus, we obtain
-$\sup_{1\leq x\leq K}\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}\leq tq\sum_{x>K}\mu_{x}.$
-Putting everything together in (18),
-$\beta(t)\leq\sum_{x=1}^{K}2\pi_{T}(x)\left[(1-\nu_{K})^{t}+3\sum_{x>K}\pi_{T}(x)+qt\sum_{x>K}\mu_{x}\right]+\sum_{x>K}\pi_{T}(x)$
-which ends the proof. ∎
-Proof of Lemma 6.3.
-We have, for a fixed $t$ and $y$,
-$\beta(t)$ $=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
-$\geq\pi_{T}(y)\|P_{T}^{t}(y,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
-$=\frac{\pi_{T}(y)}{2}\sum_{y=1}^{\infty}|P_{T}^{t}(k,y)-\pi_{T}(y)|$
-$\geq\frac{\pi_{T}(y)}{2}|P_{T}^{t}(y,y)-\pi_{T}(y)|$
-$\geq\frac{\pi_{T}(y)}{2}\left[(1-1/(2y^{a}))^{t}-\pi_{T}(y)\right].$
-Then, take $y=\lceil t^{1/a}\rceil$ to get
-$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[(1-1/t)^{t}-\pi_{T}(\lceil t^{1/a}\rceil)\right]\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[\frac{1}{4}-\pi_{T}(\lceil t^{1/a}\rceil)\right]$
-when $t>2$. ∎
-6.2   Control of $\mathcal{J}_{p}^{(s)}$ for $P_{T}$
-Proposition 6.1.
-Assume $T=(q,\mu,\nu)\in\mathcal{T}$ and
-$\sum_{x\geq 2}\frac{\mu_{x}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{x}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
-Then Condition 3 is satisfied for $P_{T}$ with $D(1)=V(1)=1$, $D(k)=V(k)=1/\sqrt{\mu_{k}}$ for $k>1$ and $b=q\sum_{k\geq 2}\sqrt{\mu_{k}}$.
-Moreover, the sumability conditions of Lemma 4.1 are also satisfied, and
-$\mathcal{J}_{p}^{(s)}\leq\left\|\pi_{T}\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi_{T}\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi_{T}\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right)<\infty.$
-This gives a nontrivial example of chain that satisfies Condition 3.
-Proof of Proposition 6.1.
-First, let us evaluate $P_{T}D(x)$. We start with $x>1$,
-$P_{T}D(x)=\sum_{y}P_{T}(x,y)D(y)=\nu_{x}D(1)+(1-\nu_{x})D(x)\leq D(x),$
-and then when $x=1$,
-$P_{T}D(1)=\sum_{y}P_{T}(1,y)D(y)=(1-q)D(1)+q\sum_{y\geq 2}\mu_{y}D(y)\leq D(1)+q\sum_{y\geq 2}\sqrt{\mu_{y}}.$
-Thus $P_{T}D(x)\leq D(x)+b$ with $b=q\sum_{y\geq 2}\sqrt{\mu_{y}}$.
-We can now check the assumption in Lemma 4.3:
-$\sup_{y\in\mathcal{X}}V(y)(P_{T}(x,y)-\pi_{T}(y))\leq D(x).$
-We actually prove the stronger:
-$\forall(x,y)\text{, }P_{T}(x,y)\leq\frac{D(x)}{V(y)}.$
-Indeed, for $x=y=1$,
-$P_{T}(1,1)=1-q\leq 1=\frac{D(1)}{V(1)}.$
-For $x>1$,
-$P_{T}(x,x)=1-\nu_{x}\leq 1=\frac{D(x)}{V(x)}.$
-Still for $x>1$,
-$P_{T}(x,1)=\nu_{x}\leq 1\leq\frac{1}{\sqrt{\mu_{x}}}=\frac{D(x)}{V(1)}.$
-Finally, for $y>1$,
-$P_{T}(1,y)=q\mu_{y}\leq\mu_{y}\leq\sqrt{\mu_{y}}=\frac{D(1)}{V(y)}.$
-Finally, let us check the summability conditions:
-$\sum_{y}\left[\frac{1}{V(y)}\right]^{\frac{1}{2}(1-\frac{1}{p})}=1+\sum_{y>1}\mu_{y}^{\frac{1}{4}(1-\frac{1}{p})}\leq 1+\sum_{y>1}\frac{\mu_{y}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{y}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
-by assumption. Then
-$\sum_{x}\left[\pi_{T}(x)D_{x}\right]^{\frac{1}{2}(1-\frac{1}{p})}\leq 1+\sum_{x>1}\left[\frac{\mu_{x}}{\nu_{x}}\frac{1}{\sqrt{\mu_{x}}}\right]^{\frac{1}{2}(1-\frac{1}{p})}\leq 1+\sum_{x>1}\frac{\mu_{x}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{x}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
-still by assumption. This ends the proof. ∎
 Acknowledgments
 GW would like to thank Luca Zanetti and John Sylvester for posing the original question of estimating the average-mixing time, and Aryeh Kontorovich for the interesting discussions.
+The authors also thank the anonymous referees for their careful reading and insightful remarks.
 Funding